### v1

# Representative Action Selection for Large Action-Space Meta-Bandits

Abstract
We study the problem of selecting a subset from a large action space shared by a family of bandits, with the goal of achieving performance nearly matching that of using the full action space. We assume that similar actions tend to have related payoffs, modeled by a Gaussian process. To exploit this structure, we propose a simple $\epsilon$-net algorithm to select a representative subset. We provide theoretical guarantees for its performance and compare it empirically to Thompson Sampling and Upper Confidence Bound.
1 Introduction
We study a family of bandits that share a common but extremely large action space. We aim to understand whether it is possible—and how—to select a smaller set of representative actions that performs nearly as well as the full action space across all bandit instances. To build intuition, imagine a pharmacy preparing its inventory for the upcoming season. The available drugs (actions) are nearly infinite, and each customer (bandit) has unique characteristics. If two drugs share similar ingredients, their effects on a patient are likely to be similar. Likewise, if two patients have comparable health indices, a drug is likely to have similar effects on both. By modeling the expected outcome of each drug for each patient as a Gaussian process, we can capture these correlations. Now, consider medicine demand: If a drug treats a very rare illness, it is unlikely to be needed frequently, so the store can exclude it to optimize storage space. Conversely, if flu season is approaching, stocking several flu medications is a wise choice.
Different from prior approaches in multi-armed bandits (MAB) (dani2008stochastic, ) that aim for identifying either a single best action or a subset that achieves high cumulative outcomes for a fixed bandit, our objective focuses on selecting a subset that is likely to contain the best action, or one whose best element performs nearly as well for a family of bandits. This problem can be seen as a large-scale combinatorial optimization under uncertainty, with applications where decisions involve a vast number of possibilities but are constrained by computational or time limitations for evaluating all options, e.g., inventory management, online recommendations.
Consider the following MAB setting: In a bandit, if a decision-maker plays an action $a\in\mathcal{A}_{full}\subset\mathbb{R}^{n}$, they observe a random outcome taking values in $\mathbb{R}$. We define the expected outcome of playing action $a$ in this bandit as $\mu_{a}(\theta):=\langle a,\theta\rangle,$ where this bandit instance $\theta\in\mathbb{R}^{n}$ follows a multivariate Gaussian distribution. Thus, the collection of random variables $\{\mu_{a}\}_{a\in\mathcal{A}_{full}}$ forms a Gaussian process (vershynin2018high, ). Our underlying setting aligns with the one considered in contextual bandit (dani2008stochastic, ; bietti2021contextual, ); however, the key difference lies in the objective.
Now consider that the decision-maker has access to a fixed action subset $\mathcal{A}\subset\mathcal{A}_{full}$. For a given bandit instance $\theta$, if the best action over the full space lies within $\mathcal{A}$, the decision-maker benefits from a reduced set of suboptimal actions to explore. Conversely, if the best action lies outside $\mathcal{A}$, regret arises from being unable to select this best action. Thus, we define the regret as the difference in expected outcome between having access to the full action space versus being restricted to the subset:
$\textnormal{Regret}:=\max_{a\in\mathcal{A}_{full}}\mu_{a}-\max_{a^{\prime}\in\mathcal{A}}\mu_{a^{\prime}},$ (1)
which depends on the sampled bandit instance and is therefore a random variable. Our objective is to identify a small subset $\mathcal{A}$ that minimizes the expected regret $\mathbb{E}_{\theta}\left[\operatorname*{Regret}\right]$ over all possible bandit instances, making the underlying optimization both stochastic and combinatorial.
This objective is motivated by practical considerations. MAB regret in finite-horizon settings (auer2002using, ; auer2002nonstochastic, ) typically scales with the number of available actions. However, if a subset $\mathcal{A}$ is carefully chosen, the resulting MAB regret can be lower than when using the full action space, due to a reduced set of suboptimal actions. This is especially beneficial when there are a very large number of actions, as even the initialization phase can be computationally expensive. This improvement arises from a natural connection between the regret in Equation (1) and MAB regret, as discussed in Supplementary A.
The remainder of this section is literature review. The subsequent sections present our main contributions and are organized as follows:
  • Section 2: Subset Selection Framework and Algorithm. We introduce a framework that models expected outcome using a Gaussian process to capture correlations across similar actions and bandit instances. Based on this structure, we construct geometric $\epsilon$-nets and refine them using an importance-based selection mechanism. This mechanism is formalized as a variant of the $\epsilon$-net algorithm (haussler1986epsilon, ), tailored to the bandit setting.
  • Section 3: Regret Analysis. We provide theoretical guarantees, including upper and lower bounds on the expected regret that reflect the geometric and importance structure of the action space.
  • Section 4: Examples and Empirical Validation. We extend our analysis to settings where outcome (payoff) functions are sampled from a Gaussian process with a kernel, and empirically compare the performance to that of Thompson Sampling and Upper Confidence Bound.
The code used for the experiments is available at https://github.com/Quan-Zhou/MAB.
1.1 Related Works
Multi-Armed Bandits The classic MAB problem (lai1985asymptotically, ; auer2002finite, ) is defined by a set of arms, each deliver outcomes that are independently drawn from a fixed and unknown distribution. The decision-maker sequentially selects an arm, observes its outcome, and aims to maximize cumulative outcomes over time. The policy is to choose the next arm based on the sequence of past plays and obtained outcomes. Popular policies include the Upper Confidence Bound (auer2002finite, ), Thompson Sampling (agrawal2012analysis, ), and EXP3 (auer2002nonstochastic, ).
Contextual Bandits Unlike classic MAB, which assumes action outcomes are independent, contextual bandits (auer2002using, ) incorporate observed features (contexts) at each round. The decision-maker selects actions based on context, aiming to learn a policy that maps contexts to high-reward actions. Linear models such as LinUCB (chu2011contextual, ) and linear Thompson sampling (agrawal2013thompson, ) assume linear reward functions, enabling efficient learning. Early work includes LinRel (auer2002using, ), while later efforts extended to nonlinear models (filippi2010parametric, ; jun2017scalable, ). Nonlinear payoff structures have also been explored using Gaussian processes (srinivas2009gaussian, ) and RKHS methods (valko2013finite, ; chowdhury2017kernelized, ). Recent studies examine richer structures: budget-constrained linear contextual bandits (agrawal2016linear, ), reward feedback across contexts (balseiro2019contextual, ), and adversarially corrupted outcomes (he2022nearly, ). Empirical evaluations of common algorithms appear in bietti2021contextual .
Online Combinatorial Optimization Several works have extended stochastic linear optimization (dani2008stochastic, ; rusmevichientong2010linearly, ) to combinatorial settings. This line of work applies the Follow-the-Perturbed-Leader (FPL) algorithm (hannan1957approximation, ; kalai2005efficient, ), which introduces random perturbations to the estimated cumulative loss of each action, resulting in simple and efficient algorithms for various combinatorial structures. This idea has also been extended to online convex optimization. For instance, (hazan2007logarithmic, ) develop an FPL-style algorithm that achieves logarithmic regret for strictly convex payoff functions, particularly in settings where the decision point lies in the simplex. Follow-the-Regularized-Leader (FTRL) (shalev2006convex, ; shalev2012online, ) generalizes Follow-the-Leader by adding a regularization term to stabilize updates. (mcmahan2011follow, ) show that many online learning algorithms, including mirror descent, can be interpreted as instances of FTRL. In the context of online combinatorial optimization, (audibert2011minimax, ; audibert2014regret, ) study minimax regret bounds and propose FTRL-based algorithms alongside other strategies. (ito2021hybrid, ) further develop a hybrid algorithm based on FTRL.
Combinatorial Bandits Combinatorial bandits (cesa2012combinatorial, ) are a subclass of online combinatorial optimization where the decision-maker selects a subset of base arms from $\mathcal{A}_{full}$ in each round, forming a super arm. Early work on online shortest path problems with adversarial edge weights (gyorgy2007line, ; uchiya2010algorithms, ; liu2012adaptive, ) adapts the EXP3 framework (auer2002nonstochastic, ). Subsequent studies establish regret bounds for Thompson Sampling in structured settings (gopalan2014thompson, ) and propose general models with non-linear rewards and semi-bandit feedback (chen2016combinatorial, ; kveton2015tight, ). The matroid bandit model (kveton2014matroid, ) captures independence constraints via matroid optimization (whitney1992abstract, ). Later work reframes the shortest path problem as an adversarial MDP (rosenberg2019online, ) and introduces improved UCB variants with adaptive confidence estimation (zhou2024towards, ).
Epsilon net We identified two definitions of $\epsilon$-nets, with further discussion in Section 2.1. The first net of a given set, referring to (vershynin2018high, ), ensures that closed balls of radius $\epsilon$ centered at net points cover the given set. This relates to the covering number, that is the smallest possible cardinality of an $\epsilon$-net. This concept extends to function classes as a complexity measure, known as metric entropy, with further details in wainwright2019high , and has been used in russo2013eluder on the class of outcome functions. The second definition, as given in (matousek2013lectures, ), relates to importance sampling, requiring that every significant subset of a set intersects with the net. The classic $\epsilon$-net algorithm, introduced by haussler1986epsilon , operates via random sampling, yielding a net of size $\mathcal{O}(\frac{1}{\epsilon}\log\frac{1}{\epsilon})$. As noted in matousek2013lectures , the importance measure extends to general probability measures, inspiring our approach. Further works aim to reduce net size (mustafa2019computing, ; pach2011tight, ; rabani2009explicit, ) and address online settings (bhore2024online, ).
2 Subset Selection Framework
We consider the problem of selecting a small number of representative action from a large (potentially infinite but compact) action space, $\mathcal{A}_{full}\subset\mathbb{R}^{n}$, $n\in\mathbb{N}$. Define the expected outcome of playing action $a$ in bandit instance $\theta$ as
$\mu_{a}(\theta):=\langle a,\theta\rangle,\quad\forall a\in\mathcal{A}_{full},\theta\sim\mathcal{N}(0,I),$ (2)
where the bandit instance $\theta$ follows a multivariate normal distribution with zero mean and identity covariance matrix $I$. Thus, the collection of random variables $\{\mu_{a}\}_{a\in\mathcal{A}_{full}}$ forms a canonical Gaussian process (vershynin2018high, ). In fact, any constant-mean Gaussian process can be represented as the canonical process; see Supplementary B for details.
Let $\operatorname*{Conv}(\mathcal{A}_{full})$ denote the convex hull that consists of all convex combinations of finitely many actions in $\mathcal{A}_{full}$. The extreme points of the convex hull are those points $x\in\operatorname*{Conv}(\mathcal{A}_{full})$ for which there do not exist distinct points $a,a^{\prime}\in\operatorname*{Conv}(\mathcal{A}_{full})$ and $\lambda\in(0,1)$ such that $x=\lambda a+(1-\lambda)a^{\prime}$.
By the extreme point theorem, i.e., the optimum of the linear function on a convex set is attained at an extreme point—it follows that the optimal action for any bandit instance is always an extreme point of the full action space. Therefore, if we use all the extreme points as representatives of the full action space, the regret (as defined in Equation (1)) is zero. See Supplementary C for details.
This example highlights how a geometric approach can be used to solve the stochastic combination problem. However, even if one only needs the extreme points, the set of extreme points may still be large or even infinite, e.g, the extreme points of a Euclidean ball is infinite. To address this, we will later introduce the notions of $\epsilon$-nets. Without loss of generality, we assume $\mathcal{A}_{full}$ consists only of the extreme points of $\operatorname*{Conv}(\mathcal{A}_{full})$, as they are the only points of interest.
2.1 Epsilon Nets
There are at least two definitions of $\epsilon$-net: one from a geometric perspective (vershynin2018high, ) and another from a measure-theoretic perspective (matousek2013lectures, ). Let $\|\cdot\|_{2}$ denote the Euclidean norm.
  • A subset $\mathcal{A}\subseteq\mathcal{A}_{full}$ is called a Geometric $\epsilon$-net if, for all $a\in\mathcal{A}_{full}$, there exists $a^{\prime}\in\mathcal{A}$ such that
    $\|a-a^{\prime}\|_{2}<\epsilon.$
  • Let $\mathcal{R}$ be a finite partition of the extreme points into disjoint clusters such that $\cup_{r\in\mathcal{R}}\;r=\mathcal{A}_{full}$. Given a measure $q$ assigning a value to each cluster $r\in\mathcal{R}$. A subset $\mathcal{A}\subseteq\mathcal{A}_{full}$ is called a Measure-Theoretic $\epsilon$-net with respect to measure $q$ if, for any cluster $r\in\mathcal{R}$, we have:
    $r\cap\mathcal{A}\neq\emptyset\quad\text{whenever}\quad q(r)>\epsilon.$
Intuitively, the geometric definition ensures point-wise closeness, while the measure-theoretic definition ensures representative sampling. In well-behaved spaces, the two coincide. For example, let $\mathcal{A}_{full}=[0,1]$ and $q$ be the uniform distribution over this interval. A grid with spacing $\epsilon$ forms a geometric $\epsilon$-net and also ensures any subinterval of measure at least $\epsilon$ contains at least a grid point.
Both definitions are useful. A geometric $\epsilon$-net ensures small regret, as it guarantees that an action close to the optimal one always exists in $\mathcal{A}$. However, it suffers from the curse of dimensionality: for $[0,1]^{n}$, the number of points needed to form a geometric $\epsilon$-net grows as $(1/\epsilon)^{n}$, exponentially in $n$. To address this, we incorporate the measure-theoretic $\epsilon$-net to restrict the geometric construction to only the important clusters—as determined by the $q$-measure.
The partition $\mathcal{R}$ is the central concept that bridges the two definitions of $\epsilon$-nets. If two actions $a,a^{\prime}$ are in the same cluster, they are close in the Euclidean sense, and the deviation between $\mu_{a}$ and $\mu_{a^{\prime}}$ is small in the $L^{2}$-sense (i.e., in expected squared difference). This holds because, for the canonical Gaussian process, the increments of the process correspond to the Euclidean distance:
$\|\mu_{a}-\mu_{a^{\prime}}\|_{L^{2}}=\left(\mathbb{E}(\mu_{a}-\mu_{a^{\prime}})^{2}\right)^{1/2}=\left(\mathbb{E}(a-a^{\prime})^{T}\theta\theta^{T}(a-a^{\prime})\right)^{1/2}=\|a-a^{\prime}\|_{2},\;\forall a,a^{\prime}\in\mathcal{A}_{full},$
where $\theta^{T}$ denotes the transpose of $\theta$. The middle & right equalities use Equation (2) and $\mathbb{E}\theta\theta^{T}=I$. Then, selecting one point from each cluster $r\in\mathcal{R}$ forms a geometric $\epsilon$-net (which will later be formally defined as a reference subset), where $\epsilon$ is the maximum diameter of all the clusters. This construction yields bounded expected regret, as each selected point serves as a representative for its cluster, and all other points within the same cluster have bounded deviation from the representative. Furthermore, clusters in $\mathcal{A}_{full}$ that are deemed unimportant for expected regret can be ignored: even if no representative points are selected from these clusters, the resulting expected regret will not increase significantly compared to that of the reference subset.
2.2 Epsilon Net Algorithm
We propose Algorithm 1, a variant of the $\epsilon$-net algorithm originally introduced by haussler1986epsilon . It selects $K$ i.i.d. random actions, aligned with the distribution of bandit instances. Since repetitions are allowed, the resulting subset $\mathcal{A}$ may have fewer than $K$ distinct actions.
Algorithm 1 Epsilon Net Algorithm
1:Input: Action space $\mathcal{A}_{full}$, Sample size $K$, Distribution of bandits.
2:Output: A subset of actions $\mathcal{A}$.
3:$\mathcal{A}\leftarrow\emptyset$
4:for $1,\dots,K$ do
5:     Sample a bandit instance $\theta$.
6:     Find the optimal action $a^{*}(\theta):=\operatorname*{argmax}_{a\in\mathcal{A}_{full}}\langle a,\theta\rangle$.
7:     $\mathcal{A}\leftarrow\mathcal{A}\cup\{a^{*}\}$triangleright Repetitions of actions are allowed
8:end for
We define the optimal action in a bandit instance $\theta$ as $a^{*}(\theta):=\arg\max_{a\in\mathcal{A}_{full}}\mu_{a}(\theta)$.
Assumption 1 (Unique optimal action).
The optimal action $a^{*}(\theta)$ is unique with probability $1$ for all bandit instances.
This assumption does not exclude the case where the action space is a polytope, and there may be specific bandit instances for which the optimal actions lie on the edges, resulting in an infinite number of optimal actions. Since the bandit instances follow a continuous distribution, the probability of encountering these specific bandits is zero.
Then, define the measure $q$:
[…]
Assumption 2 (Compact support).
the support of measure $q$ is compact.
The compactness assumption ensures that $\mathbb{E}_{\theta}[\max_{a\in\mathcal{A}_{full}}\mu_{a}]$ is finite and that the unique optimal action exists. Without loss of generality, let $\mathcal{A}_{full}$ be the support of measure $q$.
By definition, Algorithm 1 samples $K$ i.i.d. extreme points from clusters in $\mathcal{R}$ according to the probability measure $q$. If a cluster has a higher measure under $q$, its elements are more likely to be included in the output of Algorithm 1. In fact, with high probability, Algorithm 1 outputs a measure-theoretic $\epsilon$-net of $\mathcal{A}_{full}$ with respect to measure $q$. (This is a simplified version of the $\epsilon$-net theorem; see Theorem 10.2.4 of (matousek2013lectures, ).)
Lemma 3 (Measure-theoretic $\epsilon$-net).
Given a partition $\mathcal{R}$ of the full action space, and the importance measure $q$ assigning a value to each cluster $r\in\mathcal{R}$. Let $\mathcal{A}$ be the output of Algorithm 1 after $K$ samples. Then, with probability at least $1-\frac{1}{\epsilon}\exp(-K\epsilon)$, it holds that for any cluster $r\in\mathcal{R}$,
$r\cap\mathcal{A}\neq\emptyset\quad\text{whenever}\quad q(r)>\epsilon.$
3 Regret Analysis
In this section, we begin by analyzing a special class of geometric $\epsilon$-nets, constructed by partitioning the action space into clusters and selecting a single representative action from each cluster. We then extend the analysis to obtain algorithm-dependent bounds for the output of Algorithm 1.
For any cluster $r\in\mathcal{R}$, we define its diameter as $\operatorname{diam}(r):=\max_{a,b\in r}\|a-b\|_{2}$. Throughout the following, we set $\epsilon$ to be the maximum diameter over all clusters, i.e., $\epsilon:=\max_{r\in\mathcal{R}}\operatorname{diam}(r)$.
Definition 1 (Reference subsets).
Consider a partition :=$\mathcal{R}:=\{r_{\ell}\}_{\ell\leq m}$ of the full action space, with $\epsilon:=\max_{r\in\mathcal{R}}\operatorname{diam}(r)$. A reference subset is a set $\mathcal{A}:=\{a_{1},\dots,a_{m}\}$, where each representative $a_{\ell}\in\mathcal{A}$ corresponds to a cluster $r_{\ell}$, and each cluster $r_{\ell}$ is contained within a closed Euclidean ball of radius $\epsilon$ centered at $a_{\ell}$, i.e., $r_{\ell}\subset B(a_{\ell},\epsilon)$.
We impose the following assumption on the partition, used only for lower bounds. This means that when the optimal action falls within an arbitrary cluster $r$, any action in this cluster is better than any action outside of it. It ensures that the clusters are relatively well-separated.
Assumption 4 (Effective partition).
For any $r\in\mathcal{R}$, whenever the optimal action lies in cluster $r$, i.e., $a^{*}(\theta)\in r$, then $\mu_{a}\geq\max_{a^{\prime}\in\mathcal{A}_{full}\setminus r}\mu_{a^{\prime}},\;\forall a\in r$.
Theorem 5 (Regret bounds of reference subsets).
Consider a partition :=$\mathcal{R}:=\{r_{\ell}\}_{\ell\leq m}$ of the full action space, with $\epsilon:=\max_{r\in\mathcal{R}}\operatorname{diam}(r)$, and an arbitrary reference subset $\mathcal{A}$. Then, there is an absolute constant $C>0$, such that
$\mathbb{E}_{\theta}[\operatorname*{Regret}]\leq\max_{r\in\mathcal{R}}\mathbb{E}_{\theta}\left[\max_{a\in r}\mu_{a}\right]+C\epsilon\sqrt{\log|\mathcal{R}|}.$
If the partition $\mathcal{R}$ satisfies Assumption 4, then
$\mathbb{E}_{\theta}[\operatorname*{Regret}]\geq\min_{r\in\mathcal{R}}\mathbb{E}_{\theta}\left[\max_{a\in r}\mu_{a}\right]-C\epsilon\sqrt{\log|\mathcal{R}|}.$
Proof sketch. For each cluster $\ell\leq m$, define a simple Gaussian process $\left\{Z_{a}\right\}_{a\in r_{\ell}}$ where $Z_{a}:=\mu_{a}-\mu_{a_{\ell}}$. Define a non-negative random variable $Y_{\ell}:=\sup_{a\in r_{\ell}}Z_{a}$. When $a^{*}(\theta)\in r_{\ell}$, regret is upper bounded by $Y_{\ell}$ (or equal to it under Assumption 4). Thus, for any bandit $\theta$, the regret is bounded between $\min_{\ell\leq m}Y_{\ell}$ and $\max_{\ell\leq m}Y_{\ell}$. Finally, the expectations $\mathbb{E}\left[\min_{\ell\leq m}Y_{\ell}\right]$ and $\mathbb{E}\left[\max_{\ell\leq m}Y_{\ell}\right]$ can be bounded using concentration property of Gaussian process. $\square$
Theorem 5 reveals that the expected regret grows linearly with the cluster diameter but logarithmically with the number of clusters. If the action space has a large overall diameter but exhibits a clustered structure—that is, actions are naturally grouped into a few tight clusters—then the reference subset can achieve significantly lower regret. In contrast, if the action space has a scattered or uniform structure, using a moderate number of clusters would lead to large cluster diameters, which in turn increases the regret of the reference subset.
As shown in the next section, Theorem 5 paves the way for analyzing Algorithm 1 by bounding its expected regret in terms of a reference subset, plus an additional term—since the proof relies on the possibility that its output matches a reference subset. Thus, the structure-regret relationship applies to Algorithm 1’s output.
In Figure 1, we study the effect of the cluster diameters (controlled by a spread parameter) on regret. We structure the clustered action space: Five center points are fixed on the unit sphere in $\mathbb{R}^{3}$, and around each center, 200 points are sampled to form five clusters. Each point is obtained by adding Gaussian noise (mean zero, standard deviation equal to the spread parameter) to the center direction, followed by projection back onto the unit sphere. Bandits are sampled from a 3-dimensional standard Gaussian distribution, i.e., $\theta\sim\mathcal{N}(0,I)$. The left subplot shows the expected regret of Algorithm 1 with $K=10$, computed using $10^{4}$ additional bandits, as the spread varies from 0.01 to 0.5. The curves and error shade represent the mean $\pm$ one standard deviation of expected regret over 30 repetitions. The middle and right subplots display example action spaces for spread values of 0.01 and 0.5, with representative actions (purple stars) selected by Algorithm 1 with $K=10$.
Figure 1: Illustration of clustered action spaces on unit sphere in $\mathbb{R}^{3}$ and the effect of cluster diameters on regret. Five clusters are formed by generating 5 fixed center points, with 200 points sampled around each using Gaussian noise (spread controls the variance). Bandits are drawn from $\mathcal{N}(0,I)$. The left subplot shows the mean $\pm$ standard deviation of the expected regret (over 30 trials) as the spread varies from 0.01 to 0.5, using $10^{4}$ additional bandits. The middle and right subplots show example action spaces (blue dots) for spread values 0.01 and 0.5, with representative actions (purple stars) selected by Algorithm 1 with $K=10$.
3.1 Regret Bounds of Algorithm
The expected regret in previous results is taken over bandit instances $\theta$. In contrast, since the output of Algorithm 1 is random, the expected regret analyzed in this section is taken with respect to both the algorithm’s randomness (i.e., the sampled $\mathcal{A}$) and the distribution over $\theta$.
The algorithm’s regret bound is established by comparing it to that of a reference subset, for which we already have known expected regret bounds. The key difference is that the algorithm may omit some clusters, whereas the reference subset includes a representative from each cluster. However, although it may select a smaller subset, the algorithm achieves regret comparable to that of the reference subset, as it tends to omit clusters that contribute minimally to the expected regret.
Theorem 6 (Upper bound).
Consider a partition :=$\mathcal{R}:=\{r_{\ell}\}_{\ell\leq m}$ of the action space, with $\epsilon:=\max_{r\in\mathcal{R}}\operatorname{diam}(r)$. Let $\mathcal{A}$ be output of Algorithm 1. For the same constant $C>0$ in Theorem 5,
$\begin{split}\mathbb{E}_{\theta,\mathcal{A}}[\textnormal{Regret}]\leq\max_{r\in\mathcal{R}}\mathbb{E}_{\theta}\left[\max_{a\in r}\mu_{a}\right]+C\epsilon\sqrt{\log|\mathcal{R}|}+\left(\mathbb{E}_{q}\left[(1-q(r))^{2K}\right]\cdot\mathbb{E}_{\theta}\left[\max_{a\in\mathcal{A}_{full}}\mu^{2}_{a}\right]\right)^{1/2}.\end{split}$
Proof sketch. If $r_{\ell}\cap\mathcal{A}\neq\emptyset$, choose $a_{\ell}\in r_{\ell}\cap\mathcal{A}$ as the representative point. If $r_{\ell}\cap\mathcal{A}=\emptyset$, choose an arbitrary point $a_{\ell}\in r_{\ell}$. The new set $\mathcal{A}^{\prime}:=\{a_{\ell}\}_{\ell\leq m}$ forms a reference subset. Then, for each $\ell\leq m$, define a simple Gaussian process $\{Z_{a}\}_{a\in r_{\ell}}$, where $Z_{a}:=\mu_{a}-\mu_{a_{\ell}}$, and let $Y_{\ell}:=\sup_{a\in r_{\ell}}Z_{a}$. When $a^{*}(\theta)\in r_{\ell}$, we consider two cases:
  • If $r_{\ell}\cap\mathcal{A}\neq\emptyset$, the regret is upper bounded by $Y_{\ell}$, and hence by $\max_{\ell\leq m}Y_{\ell}$.
  • If $r_{\ell}\cap\mathcal{A}=\emptyset$, the regret is still bounded by $\max_{a\in\mathcal{A}_{full}}\mu_{a}$. This is justified by the assumption of $0\in\mathrm{conv}(\mathcal{A})$, since $\mathbb{E}_{\theta}[\max_{a\in\mathcal{A}}\mu_{a}]$ is invariant under affine unitary transformations of $\mathcal{A}$. $\square$
Comparing Theorem 6 with Theorem 5, we observe that the algorithm-dependent upper bound exceeds that of the reference subset by an additional term:
$\left(\mathbb{E}_{q}\left[(1-q(r))^{2K}\right]\cdot\mathbb{E}_{\theta}\left[\max_{a\in\mathcal{A}_{full}}\mu^{2}_{a}\right]\right)^{1/2}.$
Note that the term $\mathbb{E}_{q}\left[(1-q(r))^{2K}\right]$ behaves similarly to entropy. To illustrate, each cluster $r$ contributes $q(r)\cdot(1-q(r))^{2K}$. This expression is small when $q(r)$ is small, and decays rapidly when $q(r)$ is large, making the overall term negligible when $q$ is concentrated on few clusters.
In general, the partition $\mathcal{R}$ may be unknown, and while the corresponding $q$ measure exists, it remains unspecified. We therefore provide a worst-case bound using the covering number $N(\mathcal{A}_{full},\epsilon)$, which is the smallest number of points needed to form a geometric $\epsilon$-net of $\mathcal{A}_{full}$:
$N(\mathcal{A}_{full},\epsilon)=\min\left\{m\in\mathbb{N}:\exists\{a_{1},\dots,a_{m}\}\subseteq\mathcal{A}_{full},\ \forall a\in\mathcal{A}_{full},\ \exists\ell,\ \|a-a_{\ell}\|_{2}\leq\epsilon\right\}.$
Theorem 7 (Partition-independent upper bound).
Under Assumption 2, there exists a point $a_{0}$ and a constant $M>0$ such that $\mathcal{A}_{full}\subset B(a_{0},M)$, a closed Euclidean ball of radius $M$ centered at $a_{0}$. Let the action space have dimension $n$, and fix a constant $0<\epsilon<M$. Let $\mathcal{A}$ be the output of Algorithm 1. For the same constant $C>0$ in Theorem 5, and another absolute constant $c>0$,
$\mathbb{E}_{\theta,\mathcal{A}}[\operatorname*{Regret}]\leq 2\epsilon\sqrt{n}+C\epsilon\sqrt{\log N(\mathcal{A}_{full},\epsilon)},\quad\text{where }K\geq c\cdot(M/\epsilon)^{2}\cdot N(\mathcal{A}_{full},\epsilon).$
As $\epsilon\to 0^{+}$, we have $\mathbb{E}_{\theta,\mathcal{A}}[\operatorname*{Regret}]\to 0$ as $K\to\infty$.
The partition-independent upper bound in Theorem 7 depends on a positive constant $\epsilon$, which serves as a tolerance parameter up to the diameter of the action space. Given a $\epsilon$, the required number of samples $K$ scales with the square of the diameter-to-$\epsilon$ ratio, multiplied by the covering number—a geometric property of the action space. The resulting expected regret is then bounded in terms of $\epsilon$, the dimensionality, and the logarithm of the covering number. As $\epsilon$ decreases, more samples are needed, but the regret bound becomes tighter.
Further, we give a corresponding lower bound of Algorithm 1:
Theorem 8 (Lower bound).
Consider a partition :=$\mathcal{R}:=\{r_{\ell}\}_{\ell\leq m}$ of the full action space, where each cluster contains more than one action and satisfies Assumption 4. Let $\epsilon:=\max_{r\in\mathcal{R}}\operatorname{diam}(r)$. Let $\mathcal{A}$ be the output of Algorithm 1. For the same constant $C$ in Theorem 5 and another constant $c>0$,
$\begin{split}\mathbb{E}_{\theta,\mathcal{A}}[\operatorname*{Regret}]\geq&c\cdot\left(\mathbb{E}_{q}\left[(1-q(r))^{2K}\right]\right)^{1/2}\cdot\left(\min_{r\in\mathcal{R}}\mathbb{E}_{\theta}\left[\max_{a\in r}\mu_{a}\right]-C\epsilon\sqrt{\log|\mathcal{R}|}\right).\end{split}$
4 Examples
In general cases, the decision-maker may not have explicit access to the full structure of the action space, especially in high-dimensional settings. Instead, they are given a list of actions and can observe the expected outcomes of these actions through sampling. We therefore adopt a dimension-free view by treating the expected outcomes $\{\mu_{a}\}_{a\in\mathcal{A}_{full}}$ as a family of random variables indexed by an abstract set $\mathcal{A}_{full}$. In a single bandit, these expected outcomes define an outcome function $f$ over the action space $\mathcal{A}_{full}$:
$f(a):=\mu_{a},\quad\forall a\in\mathcal{A}_{full}.$ (4)
[…]
$k(a,a^{\prime})=\sum\nolimits_{i\leq\infty}\lambda_{i}\phi_{i}(a)\phi_{i}(a^{\prime}),$
where $(\phi_{i})_{i\leq\infty}$ are orthonormal eigenfunctions under $\mathbb{P}$, and $(\lambda_{i})_{i\leq\infty}$ are non-negative eigenvalues. Let outcome $f$ be of the form $f(\cdot)=\sum_{i=1}^{N}\alpha_{i}k(\cdot,a_{i})$ for some integer $N\geq 1$, and a set of points $\{a_{i}\}_{i=1}^{N}\subset\mathcal{A}_{full}$ and a weight vector $\alpha\in\mathbb{R}^{N}$. As shown in (williams2006gaussian, ; wainwright2019high, ), the function $f$ can also be written as
$\mu_{a}=f(a)=\langle\mathbf{f},\Phi(a)\rangle,$
where $\mathbf{f}$ is a vector of coefficients, and $\Phi(a)$ usually refereed as a feature map, has entries $\Phi_{i}(a)=\sqrt{\lambda_{i}}\phi_{i}(a)$. If the coefficients $\{\mathbf{f}_{i}\}_{i=1}^{\infty}$ are i.i.d. Gaussian, this formulation can be represented as the canonical Gaussian process model in Equation (2). Note that in this case, the actual action space is the one formed by the feature vectors $\Phi(a)$ for $a\in\mathcal{A}_{full}$.
Limitations: Gaussian assumptions may limit applicability in practice. We plan to extend this work to general random process with sub-Gaussian property.
4.1 Varying-dependence Actions
We study the effect of varying action dependence using a Gaussian process with a kernel. To control the degree of dependence, we use stationary RBF kernel and non-stationary Gibbs kernel (williams2006gaussian, ).
[…]
where $l$ is a length-scale parameter and $l(a):=0.1+0.9\cdot\exp(-\|a\|^{2})$ is a location-dependent length-scale function. Both of them control the dependence between actions. But, unlike the stationary RBF kernels, the Gibbs kernel allows some regions of the action space to be highly correlated over large distances, and others to have short-range correlations. That is, the correlation depends not only on the distance between actions, but also on where $a,a^{\prime}$ are located. When $l(a)=l$ is a constant, the Gibbs kernel reduces to the RBF kernel.
Sampling Outcome Functions from a Kernel: We first construct the kernel matrix $\mathbf{K}$, where each entry is given by $\mathbf{K}_{a,a^{\prime}}=k(a,a^{\prime})$, for $a,a^{\prime}\in\mathcal{A}_{full}$, depending on the choice of kernel. We then sample a Gaussian vector (a Gaussian process function evaluated at finite input) $f\sim\mathcal{N}(0,\mathbf{K})$. Under either kernels defined in Equation (5), the variance of the function value $f(a)$ is one for all $a\in\mathcal{A}_{full}$.
Figure 2: Experiments with outcome functions sampled from RBF/Gibbs kernels in Equation (5). Left: Sampled outcome functions from Gibbs kernel over fixed 1000 grid points in $[0,2]$ (blue curves, right y-axis). The histogram (purple bars, left y-axis) shows action selection frequencies by Algorithm 1 with $K=5000$, favoring regions with rougher functions and edge points. Right: Performance on solving stochastic combinatorial optimization in Equation (6), selecting $K=5$ distinct actions from 15 grid points in [0,2], with outcome functions sampled from RBF kernel at varying length-scales. Solid curves and shaded regions represent the mean and $\pm$ standard deviation of the expected regret over 50 repetitions. Results are presented for our method (green), Thompson Sampling (yellow), and UCB (orange), with TS/UCB run for 3000 rounds per repetition.
To study the effect of varying dependence on the output of Algorithm 1, we consider a fixed action space consisting of 1000 grid points in the interval $[0,2]$, using Gibbs kernel in Equation (5) to sample outcome functions. To simplify computations, we marginalize a Gaussian process defined by the kernel over the grid. The left subplot of Figure 2 provides examples of sampled outcome functions (blue curves, with the y-axis on the right-hand side), which become smoother as the actions approach the left end of the interval—indicating stronger correlations among function values in that region. We run Algorithm 1 on this action space with $K=5000$ to select actions and record the frequency of each action being selected. The resulting histogram (purple bars, with the y-axis on the left) reflects the importance measure $q$, highlighting that Algorithm 1 tends to select more actions from regions where the outcome functions are rougher—i.e., where action outcomes are less correlated and their features $\Phi(a)$ are are farther apart. Another interesting aspect of this subplot is the two high bars at the edges. Recall that the actual action space consists of feature vectors $\Phi(a)$ for $a\in\mathcal{A}_{ful}$. For actions indexed closer to 0, their feature vectors become more densely packed compared to those indexed closer to 2, resulting in more correlated outcomes. The two actions at the edges, indexed by 0 and 2, correspond to the two farthest points in the actual feature space.
As an extreme case, suppose $\mathcal{A}_{full}$ is a set of independent actions. This corresponds to the canonical process with the action space given by the orthogonal basis in $\mathbb{R}^{n}$ where $n=|\mathcal{A}_{full}|$. To see this, we associate each action $a$ with a unit vector $e_{a}$, which is $1$ at the $a^{th}$ entry but is zero otherwise, and define the expected outcome as $\mu_{a}(\theta):=\langle e_{a},\theta\rangle$. With this construction, the collection $\{\mu_{a}\}_{a\in\mathcal{A}_{full}}$ consists of mutually independent random variables. Then, as discussed in Supplementary I, for any subset $\mathcal{A}\subset\mathcal{A}_{full}$, including the output of Algorithm 1, the regret suffers from the lower bound
$\mathbb{E}_{\theta}[\textnormal{Regret}]\geq\frac{\sqrt{\log|\mathcal{A}_{full}|}}{\sqrt{\pi\log 2}}-\sqrt{2\log|\mathcal{A}|}.$
As a result, any algorithm, including Algorithm 1, would perform poorly unless the subset size $|\mathcal{A}|$ is sufficiently large.
4.2 Combinatorial Optimization under Uncertainty
Returning to the objective stated in the Introduction, we aim to find a subset $\mathcal{A}$ of cardinality $K$ that minimizes the expected regret defined in Equation (1). Since the term $\mathbb{E}[\max_{a\in\mathcal{A}_{full}}\mu_{a}]$ is a constant independent of $\mathcal{A}$, this is equivalent to the following optimization problem:
$\max_{\mathcal{A}\subseteq\mathcal{A}_{full}}\mathbb{E}\left[\max_{a\in\mathcal{A}}\mu_{a}\right]\quad\text{subject to }|\mathcal{A}|=K.$ (6)
In general, this problem is non-trivial because the objective function involves an expectation over a collection of random variables whose distribution is neither specified nor directly accessible. Moreover, the expression inside the expectation is the maximum over a subset of potentially correlated random variables. As a result, small changes in $\mathcal{A}$ can lead to non-smooth variations in the maximum.
Our method is computationally efficient, avoiding the combinatorial complexity of the optimization problem in Equation (6). To assess the effectiveness in identifying near-optimal solutions, we compare it against Thompson Sampling (TS) and Upper Confidence Bound (UCB) on a simple example: We consider a fixed action space consisting of 15 grid points in $[0,2]$, using an RBF kernel in Equation (5) to sample outcome functions while varying the length-scale $l\in\{0.5,1,\dots,4\}$. In our method, we run Algorithm 1 until $K=5$ distinct actions are selected, where in each round, the optimal action is founded via an exhaustive search over the action space. In TS/UCB methods, we treat each $K$-tuple of actions as a super arm and adopt a bandit feedback setting: at each round, the decision-maker selects a super arm $\mathcal{A}$, observes the payoff $\max_{a\in\mathcal{A}}\mu_{a}$, and updates its policy accordingly, repeating this process for 3000 rounds. Thus, they are tailored to maximize the same objective. Since the payoff $\max_{a\in\mathcal{A}}\mu_{a}$ is unbounded, we assume Gaussian payoffs for both TS and UCB, with the algorithms summarized in Supplementary J, along with examples of selected subsets and sampled outcome functions. We evaluate expected regret of all methods over $10^{5}$ randomly sampled outcome functions.
The right subplot of Figure 2 reports the expected regret of our method (green), TS (yellow) and UCB (orange), where solid curves and error shades indicate the mean $\pm$ one standard deviation of expected regret over 50 repetitions. As shown in this subplot, the expected regret decreases as the length-scale increases. This is because the length-scale $l$ of the RBF kernel controls the number of approximately independent actions in the process $\{\mu_{a}\}_{a\in\mathcal{A}_{full}}$: When actions $a$ and $a^{\prime}$ are far apart relative to the length-scale, the kernel value becomes very small, so the covariance between the function values $\mu_{a}$ and $\mu_{a^{\prime}}$ is also close to zero.
Appendix A Connection to Bayesian Regret of Multi-Armed Bandits
Given a subset $\mathcal{A}\subseteq\mathcal{A}_{full}$, let $\mu^{*}(\mathcal{A},\theta):=\max_{a\in\mathcal{A}}\mu_{a}(\theta)$ be the highest expected outcome in a bandit instance $\theta$. We can equivalently rewrite the regret in Equation (1):
$\textnormal{Regret}(\theta)=\mu^{*}(\mathcal{A}_{full},\theta)-\mu^{*}(\mathcal{A},\theta).$
Bandit Regret: In the classic multi-armed bandit setting [7, 32], outcome of an arm $a\in\mathcal{A}_{full}$ in a bandit instance $\theta$ is generated i.i.d. according to an unknown but fixed law with expectation $\mu_{a}(\theta)$. A decision-maker would follow a policy $\pi$ that chooses the next arm based on the sequence of rounds and obtained outcomes. The bandit regret of policy $\pi$ in this bandit environment $\theta$ over $N$ rounds is defined as the expected difference in cumulative outcomes over $N$ rounds between always selecting the optimal action and following a policy $\pi$ to choose actions. Bayesian (bandit) regret [2, 32] is the expectation of this bandit regret, taken over the randomness in the distribution of bandit instances, and the action chosen by a policy:
$\begin{split}\textnormal{BanditRegret}(\theta):=&\mu^{*}(\mathcal{A}_{full},\theta)-\sum_{t=1}^{N}\mu_{A_{t}}(\theta),\\ \textnormal{BayesianRegret}:=&\mathbb{E}\left[\mu^{*}(\mathcal{A}_{full},\theta)-\sum_{t=1}^{N}\mu_{A_{t}}(\theta)\right],\end{split}$
where $A_{t}$ is the action chosen by policy $\pi$ in round $t$. Since we consider settings where only a subset of the full action space is accessible, the action chosen by the policy can only from the subset $\mathcal{A}\subseteq\mathcal{A}_{full}$. This quantifies the expected difference in cumulative outcomes over $N$ rounds between always selecting the optimal action from the full action set $\mathcal{A}_{full}$ and following policy $\pi$ to choose actions from the restricted subset $\mathcal{A}$.
Next, we establish the connection between the regret defined in this paper and the Bayesian (bandit) regret.
A.1 Optimal policy
If $\pi$ is the optimal policy, it selects the best action in $\mathcal{A}$ in every round. Therefore, the regret of only having access to $\mathcal{A}$ instead of $\mathcal{A}_{full}$ for $N$ rounds, is determined by the optimal actions in $\mathcal{A}$ and $\mathcal{A}_{full}$:
[…]
A.2 Thompson sampling policy
If $\pi$ is the Thompson sampling policy, the Bayesian regret is bounded by
[…]
$=$ $\mathbb{E}\left[\sum_{t=1}^{N}\left(\mu^{*}(\mathcal{A})-\mu_{A_{t}}+\mu^{*}(\mathcal{A}_{full})-\mu^{*}(\mathcal{A})\right)\right]$
$\leq$ $C\sqrt{|\mathcal{A}|\cdot N\log N}+N\cdot\mathbb{E}_{\theta}[\textnormal{Regret}],$
where $|\mathcal{A}|$ denotes the cardinality, and $C>0$ is a constant. The inequality uses the Bayesian (bandit) regret of Thompson sampling (Theorem 36.1 in [32]). Note that with access to the full action space, the Bayesian regret is bounded by $C\sqrt{|\mathcal{A}_{full}|\cdot N\log N}$. However, this can be worse than restricting to a subset $\mathcal{A}$ when $\mathbb{E}_{\theta}[\textnormal{Regret}]$ is sufficiently small and the time horizon $N$ is finite. Intuitively speaking, using a well-chosen subset of representative actions instead of the full action space can reduce Bayesian regret in multi-armed bandits.
Appendix B Canonical Representation of a Gaussian Process
The expected outcome depends on both the chosen action $a\in\mathbb{R}^{n}$ and an observed context $g\in\mathbb{R}^{n}$, such that:
$\mu_{a}:=\langle a,g\rangle+c,\quad\;\forall\;a\in\mathcal{A}_{full},\;g\sim\mathcal{N}(0,\Sigma).$ (7)
We show how to transform it into the canonical Gaussian process. Note that
$g\equiv\Sigma^{1/2}\theta,\;\theta\sim\mathcal{N}(0,I),$
where $\equiv$ denotes equivalence in distribution. Let $\sigma_{j}$ denote the $j$-row of the matrix $\Sigma^{1/2}$. Thus, $g=(\langle\sigma_{j},\theta\rangle)_{j\leq n}$, and
$\mu_{a}=\langle a,g\rangle=\sum_{j\leq n}a_{j}g_{j}=\sum_{j\leq n}a_{j}\langle\sigma_{j},\theta\rangle=\left\langle\sum_{j\leq n}a_{j}\sigma_{j},\theta\right\rangle=\langle\Sigma^{1/2}a,\theta\rangle.$
Therefore, the setting in Equation (7) is equivalent to
$\mu_{a}=\langle a,\theta\rangle+c,\;\forall\;a\in\Sigma^{1/2}\mathcal{A}_{full},\;\theta\sim\mathcal{N}(0,I),$
where $\Sigma^{1/2}\mathcal{A}_{full}$ is the projected set of $\mathcal{A}_{full}$ by the matrix $\Sigma^{1/2}$.
Appendix C Convex hull dominated by extreme points
Lemma 9.
Given a subset $\mathcal{A}\subset\mathcal{A}_{full}$. If $\mathcal{A}$ includes all extreme points of $\operatorname*{Conv}(\mathcal{A}_{full})$, i.e., $\mathcal{A}_{full}\subseteq\operatorname*{Conv}(\mathcal{A})$, then it always holds
$\max_{a\in\mathcal{A}_{full}}\mu_{a}-\max_{a^{\prime}\in\mathcal{A}}\mu_{a^{\prime}}=0\;\;\forall\theta\in\mathbb{R}^{n}.$
Proof.
Let $\mathcal{A}=\{a_{1},a_{2},\dots,a_{K}\}$. Since $\mathcal{A}_{full}\subseteq\operatorname*{Conv}(\mathcal{A})$, for any $a\in\mathcal{A}_{full}$, there exist non-negative numbers $\lambda_{1},\dots,\lambda_{K}$ that sum up to $1$ and such that $a=\lambda_{1}a_{1}+\dots+\lambda_{K}a_{K}$. Thus, by Equation (2), for any $\theta\in\mathbb{R}^{n}$
$\mu_{a}=\langle a,\theta\rangle=\left\langle\sum_{i\in 1:K}\lambda_{i}a_{i},\theta\right\rangle=\sum_{i\in 1:K}\lambda_{i}\langle a_{i},\theta\rangle\leq\sum_{i\in 1:K}\lambda_{i}\max_{a^{\prime}\in\mathcal{A}}\langle a^{\prime},\theta\rangle=\max_{a^{\prime}\in\mathcal{A}}\mu_{a^{\prime}}.$
It yields
$\max_{a^{\prime}\in\mathcal{A}}\mu_{a^{\prime}}\leq\max_{a\in\mathcal{A}_{full}}\mu_{a}\leq\max_{a^{\prime}\in\mathcal{A}}\mu_{a^{\prime}},$
where the first inequality uses $\mathcal{A}\subset\mathcal{A}_{full}$. Thus, the two quantities are equal. ∎
Appendix D Proof of Lemma 3
Statement: Given a partition $\mathcal{R}$ of the full action space, and the importance measure $q$ assigning a value to each cluster $r\in\mathcal{R}$. Let $\mathcal{A}$ be the output of Algorithm 1 after $K$ samples. Then, with probability at least $1-\frac{1}{\epsilon}\exp(-K\epsilon)$, it holds that for any cluster $r\in\mathcal{R}$,
$r\cap\mathcal{A}\neq\emptyset\quad\text{whenever}\quad q(r)>\epsilon.$
Proof.
By definition, the algorithm draws $K$ i.i.d. samples from clusters in $\mathcal{R}$ according to the distribution $q$. Define a set of typical clusters $R_{\epsilon}:=\{r\in\mathcal{R}:q(r)>\epsilon\}$. Then,
$\Pr[r\cap\mathcal{A}\!=\!\emptyset]=(1-q(r))^{K}<(1-\epsilon)^{K}\leq\exp(-K\epsilon),\;\forall r\in R_{\epsilon},$ (8)
where the equality is the probability of missing the cluster $r$ for $K$ times, the first inequality uses the definition of $R_{\epsilon}$. The last inequality uses $\log(1-\epsilon)^{K}=K\log(1-\epsilon)$ and the inequality $\log(1-\epsilon)\leq-\epsilon$ for $\epsilon<1$ [46]. Therefore,
$\Pr[r\cap\mathcal{A}\neq\emptyset\quad\text{whenever}\quad q(r)>\epsilon]=$ $1-\Pr[r\cap\mathcal{A}\neq\emptyset\quad\exists r\in R_{\epsilon}]$
$\geq$ $1-\sum_{r\in R_{\epsilon}}\Pr[r\cap\mathcal{A}\neq\emptyset]$
$>$ $1-\sum_{r\in R_{\epsilon}}\exp(-K\epsilon)$
$=$ $1-|R_{\epsilon}|\exp(-K\epsilon)\geq 1-1/\epsilon\cdot\exp(-K\epsilon),$
where the first inequality uses union bound, the second inequality uses Equation (8). The last inequality uses the fact that there could be at most $\lfloor 1/\epsilon\rfloor$ clusters in $R_{\epsilon}$. ∎
Appendix E Proof of Theorem 5
Lemma 10 (Regret decomposition of reference subsets).
Consider a partition :=$\mathcal{R}:=\{r_{\ell}\}_{\ell\leq m}$ of the full action space and an arbitrary reference subset $\mathcal{A}$. For each $\ell\leq m$, define a Gaussian process $\left\{Z_{a}\right\}_{a\in r_{\ell}}$ where $Z_{a}:=\mu_{a}-\mu_{a_{\ell}}$. Define a non-negative random variable
$Y_{\ell}:=\sup_{a\in r_{\ell}}Z_{a}=\sup_{a\in r_{\ell}}\mu_{a}-\mu_{a_{\ell}}.$
Under Assumption 4, $\operatorname*{Regret}=Y_{\ell}$ where the cluster $r_{\ell}$ contains the optimal action. Hence,
$\mathbb{E}_{\theta}\left[\operatorname*{Regret}\right]=\sum_{\ell\leq m}q(r_{\ell})\cdot\mathbb{E}_{\theta}\left[Y_{\ell}\Big{|}a^{*}(\theta)\in r_{\ell}\right].$
Proof.
Let the cluster $r_{\ell}$ contains the optimal action, i.e., $a^{*}(\theta)\in r_{\ell}$, then
$\operatorname*{Regret}=\max_{a\in\mathcal{A}_{full}}\mu_{a}-\max_{a^{\prime}\in\mathcal{A}}\mu_{a^{\prime}}=\max_{a\in r_{\ell}}\mu_{a}-\mu_{a_{\ell}}=Y_{\ell},$
where the middle equality uses that when $a^{*}(\theta)\in r_{\ell}$, $\max_{a\in\mathcal{A}_{full}}\mu_{a}=\max_{a\in r_{\ell}}\mu_{a}$ and further with Assumption 4, $\max_{a^{\prime}\in\mathcal{A}}\mu_{a^{\prime}}\geq\mu_{a_{\ell}}$ when $a^{*}(\theta)\in r_{\ell}$. Then,
$\mathbb{E}\left[\operatorname*{Regret}\right]=\sum_{\ell\leq m}\Pr[a^{*}(\theta)\in r_{\ell}]\cdot\mathbb{E}\left[Y_{\ell}\Big{|}a^{*}(\theta)\in r_{\ell}\right]=\sum_{\ell\leq m}q(r_{\ell})\cdot\mathbb{E}\left[Y_{\ell}\Big{|}a^{*}(\theta)\in r_{\ell}\right],$
where the left equality uses tower rule, the right equality uses the definition of $q$ measure. ∎
To bound the expected regret of reference subsets, we need a few more lemmas:
Lemma 11 (Expectation integral identity).
Given a non-negative random variables $X$. If $\Pr[X\geq u]\leq c\exp{\left(-\frac{u^{2}}{\epsilon^{2}}\right)}$ for any $u>0$, then $\mathbb{E}X\leq C\epsilon\sqrt{\log c}$, where $\epsilon,c,C$ are positive constants.
Proof.
Let $u_{0}:=\epsilon\sqrt{\log c}$, then we have
[…]
$=$ $u_{0}+\exp\left(-\frac{u_{0}^{2}}{\epsilon^{2}}\right)\cdot\frac{c\epsilon^{2}}{2u_{0}}$
$=$ $\epsilon\sqrt{\log c}+\frac{\epsilon}{2\sqrt{\log c}}\leq C\epsilon\sqrt{\log c}.$
Lemma 12 (Borell-TIS inequality; Lemma 2.4.7 [45]).
Given a set $\mathcal{S}$, and a zero-mean Gaussian process $\left(X_{a}\right)_{a\in\mathcal{S}}$. Let $\epsilon:=\sup_{a\in\mathcal{S}}\left(\mathbb{E}X_{a}^{2}\right)^{\frac{1}{2}}$. Then for $u>0$, we have
$\Pr\left[\left|\sup_{a\in\mathcal{S}}X_{a}-\mathbb{E}\sup_{a\in\mathcal{S}}X_{a}\right|\geq u\right]\leq 2\exp{\left(-\frac{u^{2}}{2\epsilon^{2}}\right)}.$
It means that the size of the fluctuations of $\mathbb{E}\sup_{a\in\mathcal{S}}X_{a}$ is governed by the size of the individual random variables $X_{a}$.
Statement of Theorem 5:
Consider a partition :=$\mathcal{R}:=\{r_{\ell}\}_{\ell\leq m}$ of the full action space, with $\epsilon:=\max_{r\in\mathcal{R}}\operatorname{diam}(r)$, and an arbitrary reference subset $\mathcal{A}$. Then, for some constant $C>0$,
$\mathbb{E}_{\theta}[\operatorname*{Regret}]\leq\max_{r\in\mathcal{R}}\mathbb{E}_{\theta}\left[\max_{a\in r}\mu_{a}\right]+C\epsilon\sqrt{\log|\mathcal{R}|}.$
If the partition $\mathcal{R}$ satisfies Assumption 4, then
$\mathbb{E}_{\theta}[\operatorname*{Regret}]\geq\min_{r\in\mathcal{R}}\mathbb{E}_{\theta}\left[\max_{a\in r}\mu_{a}\right]-C\epsilon\sqrt{\log|\mathcal{R}|}.$
Proof.
Fix $\ell$, define a Gaussian process $\left\{Z_{a}\right\}_{a\in r_{\ell}}$ where $Z_{a}:=\mu_{a}-\mu_{a_{\ell}}$. Define a non-negative random variable
$Y_{\ell}:=\sup_{a\in r_{\ell}}Z_{a}=\sup_{a\in r_{\ell}}\mu_{a}-\mu_{a_{\ell}}.$
Since $\mathbb{E}\mu_{a}=0$ for all $a\in\mathcal{A}_{full}$, it holds that $\mathbb{E}Y_{\ell}=\mathbb{E}\max_{a\in r_{\ell}}\mu_{a}$.
When the cluster $r_{\ell}$ contains the optimal action, i.e., $a^{*}(\theta)\in r_{\ell}$, we have
$\operatorname*{Regret}=\max_{a\in\mathcal{A}_{full}}\mu_{a}-\max_{a^{\prime}\in\mathcal{A}}\mu_{a^{\prime}}=\max_{a\in r_{\ell}}\mu_{a}-\max_{a^{\prime}\in\mathcal{A}}\mu_{a^{\prime}}\leq\max_{a\in r_{\ell}}\mu_{a}-\mu_{a_{\ell}}=Y_{\ell}\leq\max_{\ell^{\prime}\leq m}Y_{\ell^{\prime}},$
where the first equality follows from the definition of regret in Equation (1), and the second equality follows from the assumption that $a^{*}(\theta)\in r_{\ell}$. The inequality holds because $a_{\ell}\in\mathcal{A}$, and hence $\max_{a\in\mathcal{A}}\mu_{a}\geq\mu_{a_{\ell}}$. The final equality follows from the definition of $Y_{\ell}$. On the other hand, according to Lemma 10, under Assumption 4, $\operatorname*{Regret}=Y_{\ell}$, and thus $\operatorname*{Regret}\geq\min_{\ell^{\prime}\leq m}Y_{\ell^{\prime}}$, where the cluster $r_{\ell}$ contains the optimal action.
Therefore, we reach the important conclusion that
$\min_{\ell\leq m}Y_{\ell}\leq\operatorname*{Regret}\leq\max_{\ell\leq m}Y_{\ell},$ (9)
where the left inequality holds under Assumption 4. Further, by definition $r_{\ell}\subset B(a_{\ell},\epsilon)$, such that $\mathbb{E}Z_{a}^{2}=\mathbb{E}(\mu_{a}-\mu_{a_{\ell}})^{2}=\|a-a_{\ell}\|^{2}_{2}\leq\epsilon^{2}$. Using Lemma 12 on the process $\left\{Z_{a}\right\}_{a\in r_{\ell}}$, we have
$\Pr\left[\left|Y_{\ell}-\mathbb{E}Y_{\ell}\right|\geq u\right]\leq 2\exp{\left(-\frac{u^{2}}{2\epsilon^{2}}\right)}.$
By union bound, we have
$\Pr\left[\max_{\ell\leq m}\left|Y_{\ell}-\mathbb{E}Y_{\ell}\right|\geq u\right]\leq 2m\exp{\left(-\frac{u^{2}}{2\epsilon^{2}}\right)}.$
Using Lemma 11, we have for some absolute constant $C>0$
$\mathbb{E}\max_{\ell\leq m}\left|Y_{\ell}-\mathbb{E}Y_{\ell}\right|\leq C\epsilon\sqrt{\log m}.$ (10)
Upper bound:
$\mathbb{E}\operatorname*{Regret}\leq\mathbb{E}\max_{\ell\leq m}Y_{\ell}\leq\max_{\ell\leq m}\mathbb{E}Y_{\ell}+\mathbb{E}\max_{\ell\leq m}\left|Y_{\ell}-\mathbb{E}Y_{\ell}\right|\leq\max_{\ell\leq m}\mathbb{E}\max_{a\in r_{\ell}}\mu_{a}+C\epsilon\sqrt{\log m},$ (11)
where the first inequality uses Equation (9). The second inequality uses $\max_{\ell\leq m}Y_{\ell}\leq\max_{\ell\leq m}\mathbb{E}Y_{\ell}+\max_{\ell\leq m}\left|Y_{\ell}-\mathbb{E}Y_{\ell}\right|$, since $Y_{\ell}\leq\mathbb{E}Y_{\ell}+\left|Y_{\ell}-\mathbb{E}Y_{\ell}\right|$. The last inequality uses Equation (10) and the identity $\mathbb{E}[Y_{\ell}]=\mathbb{E}[\max_{a\in r_{\ell}}\mu_{a}]$.
Lower bound:
$\mathbb{E}\operatorname*{Regret}\geq\mathbb{E}\min_{\ell\leq m}Y_{\ell}\geq\min_{\ell\leq m}\mathbb{E}Y_{\ell}-\mathbb{E}\max_{\ell\leq m}\left|Y_{\ell}-\mathbb{E}Y_{\ell}\right|\geq\min_{\ell\leq m}\mathbb{E}\max_{a\in r_{\ell}}\mu_{a}-C\epsilon\sqrt{\log m},$ (12)
where the first inequality uses Equation (9) under Assumption 4. The second inequality uses $\min_{\ell\leq m}Y_{\ell}\geq\min_{\ell\leq m}\mathbb{E}Y_{\ell}-\max_{\ell\leq m}\left|Y_{\ell}-\mathbb{E}Y_{\ell}\right|$, since $Y_{\ell}\geq\mathbb{E}Y_{\ell}-\left|Y_{\ell}-\mathbb{E}Y_{\ell}\right|$. The last inequality uses Equation (10) and the identity $\mathbb{E}[Y_{\ell}]=\mathbb{E}[\max_{a\in r_{\ell}}\mu_{a}]$. ∎
Appendix F Proof of Theorem 6
Lemma 13 (Transformation invariance).
Given any vector $x\in\mathbb{R}^{n}$, let $S+x:=\{s+x:s\in\mathcal{S}\}$. If $\mathbb{E}[\theta]=0$, then $\mathbb{E}\left[\max_{a\in\mathcal{S}+x}\langle a,\theta\rangle\right]=\mathbb{E}\left[\max_{a^{\prime}\in\mathcal{S}}\langle a^{\prime},\theta\rangle\right]$.
Proof.
$\mathbb{E}\left[\max_{a\in\mathcal{S}+x}\langle a,\theta\rangle\right]=\mathbb{E}\left[\max_{a^{\prime}\in\mathcal{S}}\langle a^{\prime}+x,\theta\rangle\right]=\mathbb{E}\left[\max_{a^{\prime}\in\mathcal{S}}\langle a^{\prime},\theta\rangle\right]+\mathbb{E}\left[\langle x,\theta\rangle\right]=\mathbb{E}\left[\max_{a^{\prime}\in\mathcal{S}}\langle a^{\prime},\theta\rangle\right],$
where the last equality uses $\mathbb{E}\left[\langle x,\theta\rangle\right]=\left\langle x,\mathbb{E}[\theta]\right\rangle=0$. ∎
Lemma 14.
Consider a partition $\mathcal{R}$ of the full action space. Let $\mathcal{A}$ be the output of Algorithm 1. Then, the event that the optimal action falls in a cluster $r$, i.e., $\{a^{*}(\theta)\in r\}$, is independent of whether the subset $\mathcal{A}$ intersects with the cluster. It holds that
$\Pr[a^{*}(\theta)\in r,\,r\cap\mathcal{A}\!=\!\emptyset]=$ $q(r)(1-q(r))^{K},$
$\Pr[a^{*}(\theta)\in r,r\cap\mathcal{A}\!\neq\!\emptyset]\leq$ $q(r).$
Proof.
Since $\mathcal{A}$ is the output of Algorithm 1, the event $\{a^{*}(\theta)\in r\}$ is independent from $\{r\cap\mathcal{A}\!\neq\!\emptyset\}$ or $\{r\cap\mathcal{A}\!=\!\emptyset\}$. We have
$\Pr[a^{*}(\theta)\in r,r\cap\mathcal{A}\!=\!\emptyset]=\Pr[a^{*}(\theta)\in r]\Pr[r\cap\mathcal{A}\!=\!\emptyset]=q(r)\Pr[r\cap\mathcal{A}\!=\!\emptyset]=q(r)(1-q(r))^{K},$
where the first equality uses independence between $\{a^{*}(\theta)\in r\}$ and $\{r\cap\mathcal{A}\!\neq\!\emptyset\}$, the second equality uses the definition of measure $q$, the third equality use the probability of missing cluster $r$ in $K$ i.i.d. samplings. Similarly,
$\Pr[a^{*}(\theta)\in r,r\cap\mathcal{A}\!\neq\!\emptyset]=q(r)\cdot\Pr[r\cap\mathcal{A}\!\neq\!\emptyset].$
Using $\Pr[r\cap\mathcal{A}\!\neq\!\emptyset]\leq 1$, we complete the proof. ∎
Statement of Theorem 6:
Let $\mathcal{A}$ be output of Algorithm 1. Consider a partition :=$\mathcal{R}:=\{r_{\ell}\}_{\ell\leq m}$ of the full action space, with $\epsilon:=\max_{r\in\mathcal{R}}\operatorname{diam}(r)$. Then, for the same constant $C>0$ in Theorem 5,
$\begin{split}\mathbb{E}_{\theta,\mathcal{A}}[\textnormal{Regret}]\leq\max_{r\in\mathcal{R}}\mathbb{E}_{\theta}\left[\max_{a\in r}\mu_{a}\right]+C\epsilon\sqrt{\log|\mathcal{R}|}+\left(\mathbb{E}_{q}\left[(1-q(r))^{2K}\right]\cdot\mathbb{E}_{\theta}\left[\max_{a\in\mathcal{A}_{full}}\mu^{2}_{a}\right]\right)^{1/2}.\end{split}$
Proof.
If $r_{\ell}\cap\mathcal{A}\neq\emptyset$, define $a_{\ell}\in r_{\ell}\cap\mathcal{A}$. If $r_{\ell}\cap\mathcal{A}\!=\!\emptyset$, choose an arbitrary point $a_{\ell}\in r_{\ell}$ as the representative. The set $\mathcal{A}^{\prime}:=\{a_{\ell}\}_{\ell\leq m}$ forms a reference subset of Definition 1. The cluster $r_{\ell}$ is contained in a closed Euclidean ball of radius $\epsilon$ centered at $a_{\ell}$, i.e., $r_{\ell}\subset B(a_{\ell},\epsilon)$. Define a Gaussian process $\left\{Z_{a}\right\}_{a\in r_{\ell}}$ where $Z_{a}:=\mu_{a}-\mu_{a_{\ell}}$. Define a random variable
$Y_{\ell}:=\sup_{a\in r_{\ell}}Z_{a}=\sup_{a\in r_{\ell}}\mu_{a}-\mu_{a_{\ell}}.$
Since $\mathbb{E}\mu_{a}=0$ for all $a\in\mathcal{A}_{full}$, we have $\mathbb{E}Y_{\ell}=\mathbb{E}\sup_{a\in r_{\ell}}\mu_{a}$.
Consider the case that $a^{*}(\theta)\in r_{\ell}$. We have
$\begin{split}\mathbb{E}\left[\operatorname*{Regret}\Big{|}r_{\ell}\cap\mathcal{A}\!\neq\!\emptyset,a^{*}(\theta)\in r_{\ell}\right]\leq&\mathbb{E}\left[Y_{\ell}\Big{|}r_{\ell}\cap\mathcal{A}\!\neq\!\emptyset,a^{*}(\theta)\in r_{\ell}\right]\\ =&\mathbb{E}\left[Y_{\ell}\Big{|}a^{*}(\theta)\in r_{\ell}\right],\\ \leq&\mathbb{E}\left[\max_{\ell\leq m}Y_{\ell}\Big{|}a^{*}(\theta)\in r_{\ell}\right],\end{split}$ (13)
where the first inequality uses $\operatorname*{Regret}\leq Y_{\ell}$ if $r_{\ell}\cap\mathcal{A}\!\neq\!\emptyset$, the equality uses that $Y_{\ell}$ is independent of $r_{\ell}\cap\mathcal{A}\!\neq\!\emptyset$. On the other hand, we assume that $0\in\operatorname*{Conv}(\mathcal{A})$ because even if it doesn’t hold, we can always find a vector $x\in\mathbb{R}^{n}$ such that $0\in\operatorname*{Conv}(\mathcal{A}+x)$, without changing the value of $\mathbb{E}[\max_{a\in\mathcal{A}}\mu_{a}]$ (c.f. Lemma 13). Therefore,
$\begin{split}\mathbb{E}\left[\operatorname*{Regret}\Big{|}r_{\ell}\cap\mathcal{A}\!=\!\emptyset,a^{*}(\theta)\in r_{\ell}\right]\leq\mathbb{E}\left[\max_{a\in\mathcal{A}_{full}}\mu_{a}\Big{|}a^{*}(\theta)\in r_{\ell}\right],\end{split}$ (14)
where the inequality uses $\operatorname*{Regret}\leq\max_{a\in\mathcal{A}_{full}}\mu_{a}$, as a consequence of $0\in\operatorname*{Conv}(\mathcal{A})$, and that $\max_{a\in\mathcal{A}_{full}}\mu_{a}$ is independent of $r_{\ell}\cap\mathcal{A}\!=\!\emptyset$. Further,
$\mathbb{E}[\operatorname*{Regret}]=$ $\sum_{r\in\mathcal{R}}\Pr[r\cap\mathcal{A}\neq\emptyset,a^{*}(\theta)\in r]\cdot\mathbb{E}\left[\operatorname*{Regret}\Big{|}r\cap\mathcal{A}\neq\emptyset,a^{*}(\theta)\in r\right]$
$+\sum_{r\in\mathcal{R}}\Pr[r\cap\mathcal{A}=\emptyset,a^{*}(\theta)\in r]\cdot\mathbb{E}\left[\operatorname*{Regret}\Big{|}r\cap\mathcal{A}=\emptyset,a^{*}(\theta)\in r\right]$
$\leq$ $\sum_{\ell\leq m}q(r_{\ell})\cdot\left(\mathbb{E}\left[\max_{\ell\leq m}Y_{\ell}\Big{|}a^{*}(\theta)\in r_{\ell}\right]+(1-q(r_{\ell}))^{K}\cdot\mathbb{E}\left[\max_{a\in\mathcal{A}_{full}}\mu_{a}\Big{|}a^{*}(\theta)\in r_{\ell}\right]\right)$
$=$ $\mathbb{E}\left[\max_{\ell\leq m}Y_{\ell}\right]+\sum_{r\in\mathcal{R}}q(r)\cdot(1-q(r))^{K}\cdot\mathbb{E}\left[\max_{a\in\mathcal{A}_{full}}\mu_{a}\Big{|}a^{*}(\theta)\in r\right]$
$\leq$ $\max_{r\in\mathcal{R}}\mathbb{E}\max_{a\in r}\mu_{a}+C\epsilon\sqrt{\log|\mathcal{R}|}+\left(\mathbb{E}_{q}\left[(1-q(r))^{2K}\right]\cdot\mathbb{E}_{\theta}\left[\max_{a\in\mathcal{A}_{full}}\mu^{2}_{a}\right]\right)^{1/2},$
where the first equality uses tower rule. The first inequality uses Equations (13-14), and Lemma 14. The last equality uses tower rule again. The last inequality uses Equation (11) in the proof of Theorem 5 on the summand $\mathbb{E}\left[\max_{\ell\leq m}Y_{\ell}\right]$. The other summand is bounded via Cauchy-Schwarz inequality, which states that for any two random variables $X$ and $Y$, we have $|\mathbb{E}[XY]|\leq\sqrt{\mathbb{E}[X^{2}]\mathbb{E}[Y^{2}]}$. ∎
Appendix G Proof of Theorem 7
Lemma 15.
Let $q$ denote a discrete probability distribution over a finite support $\mathcal{R}$. Define
$M:=\max_{q}\mathbb{E}_{q}\left[(1-q(r))^{K}\right]\quad\text{s.t. }\sum_{r\in\mathcal{R}}q(r)=1,\;q(r)\in[0,1)\;\forall r\in\mathcal{R}.$ (15)
When $|\mathcal{R}|\geq K+1$, the maximum is $M=\left(1-\frac{1}{|\mathcal{R}|}\right)^{K}$, and it is attained when $q$ is uniform.
When $|\mathcal{R}|<K+1$, the maximum is upper bounded by $M\leq\frac{|\mathcal{R}|}{K+1}\left(\frac{K}{K+1}\right)^{K}$, and there exists a feasible distribution $q^{\prime}$ such that $\mathbb{E}_{q^{\prime}}\left[(1-q^{\prime}(r))^{K}\right]\geq\frac{|\mathcal{R}|-1}{K+1}\left(\frac{K}{K+1}\right)^{K}$.
Proof.
Let =$\mathcal{R}=\{r_{1},\dots,r_{m}\}$ be the finite support of the measure $q$, where $m:=|\mathcal{R}|$. Let $q_{i}:=q(r_{i})$ denote the probability mass at each support point. Define the function $f(q_{\ell}):=q_{\ell}\cdot(1-q_{\ell})^{K}$.
Case I $m<K+1$: The first derivative of $f$ is
$f^{\prime}(q_{\ell})=(1-q_{\ell})^{K-1}\big{(}1-(K+1)q_{\ell}\big{)},$
which is positive on the interval $\left[0,\frac{1}{K+1}\right)$ and negative on $\left(\frac{1}{K+1},1\right)$. Therefore, $f(q_{\ell})$ attains its maximum over $[0,1)$ at
$q^{*}_{\ell}=\frac{1}{K+1},$
with the corresponding maximum value
$f(q^{*}_{\ell})\leq\frac{1}{K+1}\left(\frac{K}{K+1}\right)^{K}.$
Since there are $m$ support points, this yields the upper bound $M\leq\sum_{\ell=1}^{m}f(q^{*}_{\ell})\leq\frac{m}{K+1}\left(\frac{K}{K+1}\right)^{K}$. Also, since $\frac{1}{K+1}<\frac{1}{m}$, the solution $q_{1}=\dots=q_{m-1}=\frac{1}{K+1}$ and $q_{m}=\frac{K-m+2}{K+1}$ is feasible. Thus,
$M\geq\sum_{\ell=1}^{m-1}f\left(\frac{1}{K+1}\right)+f\left(\frac{K-m+2}{K+1}\right)\geq\frac{m-1}{K+1}\left(\frac{K}{K+1}\right)^{K},$
where the right inequality uses that $f(q_{\ell})\geq 0$ for $q_{\ell}\in[0,1]$.
Case II $m\geq K+1$: Consider the relaxed maximization problem of Equation (15):
$\max_{q_{1},\dots,q_{m}\in[0,1)}\sum_{\ell=1}^{m}f(q_{\ell})\quad\text{subjectto }\sum_{\ell=1}^{m}q_{\ell}\leq 1.$ (16)
Let $\lambda\geq 0$ be the Lagrange multiplier associated with the constraint. Define the Lagrangian:
$\mathcal{L}(q_{1},\dots,q_{m},\lambda)=\sum_{\ell=1}^{m}q_{\ell}(1-q_{\ell})^{K}-\lambda\left(\sum_{\ell=1}^{m}q_{\ell}-1\right).$
For each $\ell=1,\dots,m$, compute the partial derivative of $\mathcal{L}$ with respect to $q_{\ell}$:
$\frac{\partial}{\partial q_{\ell}}\left[q_{\ell}(1-q_{\ell})^{K}\right]=(1-q_{\ell})^{K}-Kq_{\ell}(1-q_{\ell})^{K-1}.$
Setting this derivative equal to zero yields the stationary condition:
$(1-q_{\ell})^{K-1}(1-(K+1)q_{\ell})=\lambda,\quad\text{with }\lambda\geq 0.$
To find critical points of Equation (16), we solve the system:
$(1-q_{\ell})^{K-1}(1-(K+1)q_{\ell})=\lambda\geq 0\quad\forall\ell\leq m,\quad\text{and}\quad\sum_{\ell=1}^{m}q_{\ell}\leq 1.$
Define the function $g(q_{\ell}):=(1-q_{\ell})^{K-1}(1-(K+1)q_{\ell})$. For $g(q_{\ell})\geq 0$, it must hold that $1-(K+1)q_{\ell}\geq 0$, i.e., $q_{\ell}\leq\frac{1}{K+1}$. Therefore, any feasible solution to this system must satisfy $q_{\ell}\in\left[0,\frac{1}{K+1}\right]$ for all $\ell\leq m$.
Then, over the interval $q_{\ell}\in\left[0,\frac{1}{K+1}\right]$, both factors $(1-q_{\ell})^{K-1}$ and $1-(K+1)q_{\ell}$ are positive and decreasing. Hence, $g(q_{\ell})$ is positive and strictly decreasing, so the equation $g(q_{\ell})=\lambda\geq 0$ has at most one solution. Therefore, all $q_{\ell}$’s must be equal at a critical point. Let $q_{\ell}=c$ for all $\ell\leq m$. Due to the assumption of $m\geq K+1$, we have $\frac{1}{m}\leq\frac{1}{K+1}$, so any choice of $c\in\left[0,\frac{1}{m}\right]$ satisfies the constraint $\sum_{\ell=1}^{m}q_{\ell}\leq 1$ and is feasible for the system.
Therefore, $q_{\ell}=c$ for all $\ell\leq m$ is a feasible critical point of Equation (16). The corresponding objective value is:
$\sum_{\ell=1}^{m}f(c)=m\cdot c\cdot(1-c)^{K},$
which is increasing in $c$ over the interval $\left[0,\frac{1}{m}\right]$. Hence, the maximum is attained at $c=\frac{1}{m}$, and the optimal value is:
$\sum_{\ell=1}^{m}f\left(\frac{1}{m}\right)=\left(1-\frac{1}{m}\right)^{K}=\left(1-\frac{1}{|\mathcal{R}|}\right)^{K},$
achieved when $q_{\ell}=\frac{1}{m}$ for all $\ell$.
To confirm that this critical point is indeed a maximum, observe that the Hessian of the objective function $f(q_{\ell})$ is diagonal (since the function is separable), and the diagonal entries are:
$\frac{\partial^{2}}{\partial q_{\ell}^{2}}\left[q_{\ell}(1-q_{\ell})^{K}\right]=-(1-q_{\ell})^{K-2}\left(2K-K(K+1)q_{\ell}\right),$
which is negative for $q_{\ell}\leq\frac{1}{K+1}$, because then $02-(K+1)q_{\ell}>$ 0. Hence, the Hessian is negative definite, and the critical point is a local (and thus global) maximum.
Finally, note that Equation (16) is a relaxation of Equation (15). While the optimum of the original problem is upper bounded by that of the relaxed problem, the optimal solution to the relaxed problem also lies within the feasible region of the original problem. Therefore, the maximum of Equation (15) is also attained when $q$ is uniform, with the maximum value being $\left(1-\frac{1}{|\mathcal{R}|}\right)^{K}$. ∎
Statement of Theorem 7
Under Assumption 2, there exists a point $a_{0}$ and a constant $M>0$ such that $\mathcal{A}_{full}\subset B(a_{0},M)$, a closed Euclidean ball of radius $M$ centered at $a_{0}$. Let the action space have dimension $n\in\mathbb{N}$, and fix a constant $0<\epsilon<M$. Let $\mathcal{A}$ be the output of Algorithm 1. For the same constant $C>0$ in Theorem 5, we have:
$\mathbb{E}_{\theta,\mathcal{A}}[\operatorname*{Regret}]\leq 2\epsilon\sqrt{n}+C\epsilon\sqrt{\log N(\mathcal{A}_{full},\epsilon)},\quad\text{where }K\geq\frac{1}{2}\left(\frac{M^{2}N(\mathcal{A}_{full},\epsilon)}{\epsilon^{2}e}-1\right).$
As $\epsilon\to 0^{+}$, we have $\mathbb{E}_{\theta,\mathcal{A}}[\operatorname*{Regret}]\to 0$ as $K\to\infty$.
Proof.
By Lemma 13, we can shift the action space to be centered at the origin. So, without loss of generality, we assume that $\mathcal{A}_{full}\subset M\cdot B_{2}^{n}$, the scaled unit Euclidean ball in $\mathbb{R}^{n}$. Then,
$\mathbb{E}_{\theta}\left[\max_{a\in\mathcal{A}_{full}}\mu_{a}^{2}\right]\leq M^{2}\cdot\mathbb{E}\|\theta\|_{2}^{2}=M^{2}n,$ (17)
where the inequality follows from $\mu_{a}=\langle\theta,a\rangle\leq M\|\theta\|_{2}$ for all $a\in MB_{2}^{n}$, and the equality uses that $\theta\sim\mathcal{N}(0,I)$, so each coordinate has variance 1 and $\|\theta\|_{2}^{2}=\sum_{i=1}^{n}\theta_{i}^{2}$ has expectation $n$.
Let $\{a_{1},\dots,a_{m}\}\subseteq\mathcal{A}_{full}$ be a minimal geometric $\epsilon$-net under the Euclidean norm, so that $m=N(\mathcal{A}_{full},\epsilon)$. Define $\pi(a)$ as the closest point in the $\epsilon$-net to $a$, and let the partition =$\mathcal{R}=\{r_{1},\dots,r_{m}\}$ be given by
$r_{\ell}=\left\{a\in\mathcal{A}_{full}:\pi(a)=a_{\ell}\right\}.$
Then,
$\max_{\ell\leq m}\mathbb{E}_{\theta}\left[\max_{a\in r_{\ell}}\mu_{a}\right]\leq\mathbb{E}_{\theta}\left[\max_{a\in B(a_{\ell},\epsilon)}\mu_{a}\right]=\mathbb{E}_{\theta}\left[\max_{a\in\epsilon B_{2}^{n}}\mu_{a}\right]\leq\epsilon\sqrt{n},$ (18)
where the first inequality follows from $r_{\ell}\subset B(a_{\ell},\epsilon)$ by construction; the equality uses Lemma 13 to shift; and the final bound uses Claim 3 from Supplementary K. Then,
$\mathbb{E}_{\theta,\mathcal{A}}[\textnormal{Regret}]\leq$ $\max_{r\in\mathcal{R}}\mathbb{E}_{\theta}\left[\max_{a\in r}\mu_{a}\right]+C\epsilon\sqrt{\log m}+\left(\mathbb{E}_{q}\left[(1-q(r))^{2K}\right]\cdot\mathbb{E}_{\theta}\left[\max_{a\in\mathcal{A}_{full}}\mu^{2}_{a}\right]\right)^{1/2},$
$\leq$ $\epsilon\sqrt{n}+C\epsilon\sqrt{\log m}+\sqrt{\frac{m}{2K+1}}\left(\frac{2K}{2K+1}\right)^{K}\cdot M\sqrt{n}$
$\leq$ $\epsilon\sqrt{n}+C\epsilon\sqrt{\log m}+\epsilon\sqrt{n},$
where the first inequality follows from the algorithm-dependent upper bound in Theorem 6. The second inequality follows from Equations (17) and (18), and the term $\mathbb{E}_{q}\left[(1-q(r))^{2K}\right]$ is bounded by $\frac{m}{2K+1}\left(\frac{2K}{2K+1}\right)^{2K}$ (see Lemma 15). The last inequality comes from
$\log\left(\sqrt{\frac{m}{2K+1}}\left(\frac{2K}{2K+1}\right)^{K}M\right)=$ $\tfrac{1}{2}\log\left(\tfrac{m}{2K+1}\right)+K\log\left(\tfrac{2K}{2K+1}\right)+\log M$
$\leq$ $\tfrac{1}{2}\log m-\tfrac{1}{2}\log(2K+1)-\tfrac{1}{2}+\log M$
$\leq$ $\tfrac{1}{2}\log m-\tfrac{1}{2}\log\left(\frac{mM^{2}}{\epsilon^{2}e}\right)-\tfrac{1}{2}+\log M\leq\log\epsilon,$
where the first inequality uses $\tfrac{1}{2}\log\left(\tfrac{m}{2K+1}\right)=\tfrac{1}{2}(\log m-\log(2K+1))$. Also, $\log\left(\tfrac{2K}{2K+1}\right)=\log\left(1-\tfrac{1}{2K+1}\right)\leq-\tfrac{1}{2K+1}$ due to the inequality $\log(1-X)\leq-X$ for $X<1$, and for a large $K$, we approximate $-\tfrac{K}{2K+1}\approx-\frac{1}{2}$. The last inequality uses the assumption $K\geq\frac{1}{2}\left(\frac{mM^{2}}{\epsilon^{2}e}-1\right)$.
Finally, since $\mathcal{A}_{full}\subset MB_{2}^{n}$, the covering number satisfies $N(\mathcal{A}_{full},\epsilon)\leq C^{\prime}(M/\epsilon)^{n}$ for some constant $C^{\prime}>0$; see Proposition 4.2.12 of [49]. Since $\epsilon\sqrt{\log(M/\epsilon)}\to 0$ as $\epsilon\to 0^{+}$, it follows that $\epsilon\sqrt{\log N(\mathcal{A}_{full},\epsilon)}\to 0$ as $\epsilon\to 0^{+}$. ∎
Appendix H Proof of Theorem 8
Consider a partition :=$\mathcal{R}:=\{r_{\ell}\}_{\ell\leq m}$ of the full action space, where each cluster contains more than one action and satisfies Assumption 4. Let $\epsilon:=\max_{r\in\mathcal{R}}\operatorname{diam}(r)$. Let $\mathcal{A}$ be the output of Algorithm 1. Then, for the same constant $C$ in Theorem 5 and another constant $c>0$, for any reference subset $\mathcal{A}^{\prime}$:
$\begin{split}\mathbb{E}_{\theta,\mathcal{A}}[\operatorname*{Regret}]\geq&c\cdot\left(\mathbb{E}_{q}(1-q(r))^{2K}\right)^{1/2}\cdot\mathbb{E}_{\theta}\left[\operatorname*{Regret}\textnormal{of }\mathcal{A}^{\prime}\right]\\ \geq&c\cdot\left(\mathbb{E}_{q}(1-q(r))^{2K}\right)^{1/2}\cdot\left(\min_{r\in\mathcal{R}}\mathbb{E}_{\theta}\left[\max_{a\in r}\mu_{a}\right]-C\epsilon\sqrt{\log|\mathcal{R}|}\right).\end{split}$
Proof.
Choose an arbitrary point $a_{\ell}\in r_{\ell}$ as the representative, such that the subset $\mathcal{A}^{\prime}=\{a_{\ell}\}_{\ell\leq m}$ forms a reference subset of Definition 1. Fix $\ell$. As in Lemma 10, define a Gaussian process $(Z_{a})_{a\in r_{\ell}}$ by setting $Z_{a}:=\mu_{a}-\mu_{a_{\ell}}$, and define random variable $Y_{\ell}:=\sup_{a\in r_{\ell}}Z_{a}$. The key idea is that $\operatorname*{Regret}\geq Y_{\ell}$ whenever $r_{\ell}\cap\mathcal{A}=\emptyset$ and $a^{*}(\theta)\in r_{\ell}$, such that
$\begin{split}\mathbb{E}\left[\operatorname*{Regret}\Big{|}r_{\ell}\cap\mathcal{A}\!=\!\emptyset,a^{*}(\theta)\in r_{\ell}\right]=&\mathbb{E}\left[\max_{a\in r_{\ell}}\mu_{a}-\max_{a^{\prime}\in\mathcal{A}}\mu_{a^{\prime}}\Big{|}r_{\ell}\cap\mathcal{A}\!=\!\emptyset,a^{*}(\theta)\in r_{\ell}\right]\\ \geq&\mathbb{E}\left[\max_{a\in r_{\ell}}\mu_{a}-\mu_{a_{\ell}}\Big{|}r_{\ell}\cap\mathcal{A}\!=\!\emptyset,a^{*}(\theta)\in r_{\ell}\right]\\ =&\mathbb{E}\left[Y_{\ell}\Big{|}a^{*}(\theta)\in r_{\ell}\right],\end{split}$ (19)
where the inequality follows from Assumption 4. The last equality follows from $Y_{\ell}$ is independent from $r_{\ell}\cap\mathcal{A}\!=\!\emptyset$. Further,
$\begin{split}\mathbb{E}\left[\operatorname*{Regret}\right]=&\sum_{r\in\mathcal{R}}\Pr[r\cap\mathcal{A}\neq\emptyset,a^{*}(\theta)\in r]\cdot\mathbb{E}\left[\operatorname*{Regret}\Big{|}r\cap\mathcal{A}\neq\emptyset,a^{*}(\theta)\in r\right]\\ &+\sum_{r\in\mathcal{R}}\Pr[r\cap\mathcal{A}=\emptyset,a^{*}(\theta)\in r]\cdot\mathbb{E}\left[\operatorname*{Regret}\Big{|}r\cap\mathcal{A}=\emptyset,a^{*}(\theta)\in r\right]\\ \geq&\sum_{r\in\mathcal{R}}q(r)\cdot(1-q(r))^{K}\cdot\mathbb{E}\left[\operatorname*{Regret}\Big{|}r\cap\mathcal{A}=\emptyset,a^{*}(\theta)\in r\right],\\ \geq&\sum_{\ell\leq m}q(r_{\ell})\cdot(1-q(r_{\ell}))^{K}\cdot\mathbb{E}\left[Y_{\ell}\Big{|}a^{*}(\theta)\in r_{\ell}\right]\\ \geq&c\cdot\left(\mathbb{E}_{q}(1-q(r))^{2K}\right)^{1/2}\cdot\left(\sum_{\ell\leq m}q(r_{\ell})\cdot\mathbb{E}^{2}\left[Y_{\ell}\Big{|}a^{*}(\theta)\in r_{\ell}\right]\right)^{1/2}\\ \geq&c\cdot\left(\mathbb{E}_{q}(1-q(r))^{2K}\right)^{1/2}\cdot\mathbb{E}_{\theta}\left[\operatorname*{Regret}\textnormal{of }\mathcal{A}^{\prime}\right]\\ \geq&c\cdot\left(\mathbb{E}_{q}(1-q(r))^{2K}\right)^{1/2}\cdot\left(\min_{r\in\mathcal{R}}\mathbb{E}_{\theta}\left[\max_{a\in r}\mu_{a}\right]-C\epsilon\sqrt{\log|\mathcal{R}|}\right)\end{split}$
where the equality uses the tower rule. The first inequality holds because the regret is lower bounded by zero when $r\cap\mathcal{A}\neq\emptyset$ and $a^{*}(\theta)\in r$, and Lemma 14. The second inequality follows from Equation (19). The third inequality uses the Pólya-Szegő inequality [17], corresponding to a constant $c>0$:
$c:=\frac{1}{2}\left(\sqrt{\frac{M_{1}M_{2}}{m_{1}m_{2}}}+\sqrt{\frac{m_{1}m_{2}}{M_{1}M_{2}}}\right),$
where $m_{1},m_{2},M_{1},M_{2}$ are constants such that
[…]
Note that $m_{2}$ is positive due to the assumption that each cluster contains more than one action. The fourth inequality follows from Jensen’s inequality that for a random variable $X$, $\mathbb{E}[X^{2}]\geq\mathbb{E}[X]^{2}$, and Lemma 10. Since the reference subset $\mathcal{A}^{\prime}$ is arbitrary, the lower bound $c\cdot\left(\mathbb{E}_{q}(1-q(r))^{2K}\right)^{1/2}\cdot\mathbb{E}_{\theta}\left[\operatorname*{Regret}\textnormal{of }\mathcal{A}^{\prime}\right]$ holds for any reference subset. The last inequality uses Theorem 5. ∎
Appendix I Example of orthogonal basis
In general, even when the distribution of bandit instances $\theta$ is well-behaved, e.g., the standard Gaussian distribution in $\mathbb{R}^{n}$, the importance measure $q$ is usually non-trivial. In this section, we consider an example of the action space being orthogonal basis.
Let the set $\mathcal{A}_{full}=\{e_{i}:i=1,\dots,n\}$ consists of the $n$ unit vectors aligned with the coordinate axes in $\mathbb{R}^{n}$. Hence, $\operatorname{diam}(\mathcal{A}_{full})=\sqrt{2}$. In this case, $\max_{a\in\mathcal{A}_{full}}\langle a,\theta\rangle$ is equivalent to the maximum among the $n$ entries of $\theta$, where by definition, each entry follows i.i.d. from the standard normal distribution $\mathcal{N}(0,1)$, such that each entry has the same probability of having the maximum value. Let each cluster contain only one unit vector, thus the importance measure of clusters $q$ is a uniform distribution.
Expected maximum in $\mathcal{A}_{full}$: Let $X_{i}$, $i=1,\dots,n$ be i.i.d. samples from $\mathcal{N}(0,1)$, then
$\mathbb{E}\max_{a\in\mathcal{A}_{full}}\mu_{a}=\mathbb{E}_{\theta}\max_{i=1,\dots,n}\theta_{i}=\mathbb{E}\max_{i=1,\dots,n}X_{i},$
where the equivalence comes from that each entry $\theta_{i}$ is i.i.d. sample from $\mathcal{N}(0,1)$. Then, according to the bounds of expected maximum of Gaussian in Supplementary L, we have
$\frac{\sqrt{\log n}}{\sqrt{\pi\log 2}}\leq\;\mathbb{E}\max_{a\in\mathcal{A}_{full}}\mu_{a}\leq\sqrt{2\log n},\quad\frac{\sqrt{\log|\mathcal{A}|}}{\sqrt{\pi\log 2}}\leq\;\mathbb{E}\max_{a\in\mathcal{A}}\mu_{a}\leq\sqrt{2\log|\mathcal{A}|}.$
Bounds of expected regret of arbitrary $\mathcal{A}$: By definition of regret in Equation (1),
$\frac{\sqrt{\log n}}{\sqrt{\pi\log 2}}-\sqrt{2\log|\mathcal{A}|}\leq\mathbb{E}_{\theta}[\textnormal{Regret}]\leq\sqrt{2\log n}-\frac{\sqrt{\log|\mathcal{A}|}}{\sqrt{\pi\log 2}}.$
Upper bound of expected regret of Algorithm 1: Since the importance measure of clusters $q$ is a uniform distribution, we have
$\mathbb{E}_{\theta,\mathcal{A}}[\textnormal{Regret}]\leq\left(1-\frac{1}{n}\right)^{K}\cdot\mathbb{E}_{\theta}\left[\max_{a\in\mathcal{A}_{full}}\mu_{a}\right]\leq\left(1-\frac{1}{n}\right)^{K}\sqrt{2\log n},$
where the left inequality uses Corollary 16 and that if each cluster contains one action, $\mathbb{E}_{\theta}\left[\max_{a\in r}\mu_{a}\right]=0$ for each cluster and $\epsilon:=\max_{r\in\mathcal{R}}\operatorname{diam}(r)=0$.
Corollary 16.
Let $\mathcal{A}$ be output of Algorithm 1. Consider a partition :=$\mathcal{R}:=\{r_{\ell}\}_{\ell\leq m}$ of the full action space, where the $q$ measure is uniform over all clusters. Let $\epsilon:=\max_{r\in\mathcal{R}}\operatorname{diam}(r)$. Then, for the same constant $C>0$ in Theorem 5, it holds that
$\mathbb{E}_{\theta,\mathcal{A}}[\textnormal{Regret}]\leq\max_{r\in\mathcal{R}}\mathbb{E}_{\theta}\left[\max_{a\in r}\mu_{a}\right]+C\epsilon\sqrt{\log|\mathcal{R}|}+\left(1-\frac{1}{|\mathcal{R}|}\right)^{K}\mathbb{E}_{\theta}\left[\max_{a\in\mathcal{A}_{full}}\mu_{a}\right].$
Under Assumption 4,
$\mathbb{E}_{\theta,\mathcal{A}}[\operatorname*{Regret}]\geq\left(1-\frac{1}{|\mathcal{R}|}\right)^{K}\cdot\left(\min_{r\in\mathcal{R}}\mathbb{E}_{\theta}\left[\max_{a\in r}\mu_{a}\right]-C\epsilon\sqrt{\log|\mathcal{R}|}\right).$
Proof.
Consider the variables $Y_{\ell}$ defined in the proof of Theorem 6,
$\mathbb{E}[\operatorname*{Regret}]\leq$ $\mathbb{E}\left[\max_{\ell\leq m}Y_{\ell}\right]+\sum_{r\in\mathcal{R}}q(r)\cdot\left(1-\frac{1}{|\mathcal{R}|}\right)^{K}\cdot\mathbb{E}\left[\max_{a\in\mathcal{A}_{full}}\mu_{a}\Big{|}a^{*}(\theta)\in r\right]$
$=$ $\mathbb{E}\left[\max_{\ell\leq m}Y_{\ell}\right]+\left(1-\frac{1}{|\mathcal{R}|}\right)^{K}\cdot\mathbb{E}_{\theta}\left[\max_{a\in\mathcal{A}_{full}}\mu_{a}\right],$
where the first inequality comes from the proof of Theorem 6, the equality uses $q(r)=1/|\mathcal{R}|$ for all $r\in\mathcal{R}$. Then, using Equation (11) in the proof of Theorem 5, we derive the upper bound.
Consider the variables $Y_{\ell}$ defined in the proof of Theorem 8,
$\mathbb{E}[\operatorname*{Regret}]\geq$ $\sum_{\ell\leq m}q(r_{\ell})\cdot\left(1-\frac{1}{|\mathcal{R}|}\right)^{K}\cdot\mathbb{E}\left[Y_{\ell}\Big{|}a^{*}(\theta)\in r_{\ell}\right]$
$=$ $\left(1-\frac{1}{|\mathcal{R}|}\right)^{K}\cdot\sum_{\ell\leq m}q(r_{\ell})\cdot\mathbb{E}\left[Y_{\ell}\Big{|}a^{*}(\theta)\in r_{\ell}\right]\geq\mathbb{E}\left[\min_{\ell\leq m}Y_{\ell}\right],$
where the first inequality comes from the proof of Theorem 8, the equality uses $q(r)=1/|\mathcal{R}|$ for all $r\in\mathcal{R}$. The last inequality uses the definition of $q$ measure and $Y_{\ell}\geq\min_{\ell\leq m}Y_{\ell}$. Then, using Equation (12) in the proof of Theorem 5, we derive the lower bound. ∎
Appendix J Comparisons with UCB & TS
To understand why Algorithm 1 outperforms UCB and TS, we still consider the fixed action space of 15 grid points over the interval $[0,2]$ and indexed by $0,1,\ldots,14$. Outcome functions are generated using an RBF kernel with a fixed length-scale $l=1$. The middle upper subplot of Figure 4 displays ten example outcome functions sampled from this setup. In each trial, Algorithms 1, 2 (UCB), and 3 (TS) are run to select $K=5$ distinct representative actions. UCB and TS are each executed for 3000 rounds (excluding initialization), with each round corresponding to a bandit instance. The number of rounds is chosen to approximately match the number of possible super arms, given by $\binom{15}{5}=3003$.
The upper part of Figure 3 visualizes the representative actions selected by each algorithm in three individual trials. Each row corresponds to one trial. The lower part shows the frequency with which each of the 15 actions is selected across 50 independent trials. The frequencies for actions indexed 1 through 13 are relatively uniform across algorithms, but a notable distinction is that Algorithm 1 (green) exhibits a strong preference toward the boundary actions—those indexed 0 and 14.
To understand this preference of Algorithm 1, recall that its output frequencies directly reflect the underlying measure $q$, since the algorithm samples actions according to $q$ by design. In Figure 4, we vary the RBF kernel length-scale across values 0.1, 1, and 4. For each setting, we sample ten example outcome functions (shown in the upper part) and run Algorithm 1 for 50 trials, recording the frequency of each selected action as an approximation of the $q$-measure (shown in the lower part).
In Figure 4, when the length-scale is 4, the outcome functions are very smooth—almost linear—so the maximum is typically attained at the boundary actions (indices 0 and 14), causing the $q$-measure to concentrate heavily on them. In contrast, with a length-scale of 0.1, the functions are very rough, with outcomes nearly independent across actions, resulting in an approximately uniform $q$-measure over the action space. Further, since the outcome functions are generated using a stationary RBF kernel, the corresponding feature representations $\Phi(a)$, for $a\in\mathcal{A}_{full}$, are uniformly distributed. When action outcomes are nearly independent (e.g., with a very small length-scale), the feature vectors $\Phi(a)$ form an orthogonal basis, with each pair $(\Phi(a),\Phi(a^{\prime}))$ equidistant. As discussed in Supplementary I, this leads to a uniform $q$-measure when each cluster contains a single action. Indeed, the lower-left part of Figure 4 confirms this: for length-scale $l=0.1$, the action selection frequencies are nearly uniform. As the length-scale increases, action outcomes become more correlated, and the pairwise distances between $\Phi(a)$ vectors shrink. The boundary actions (indices 0 and 14), remain the farthest apart and thus are more likely to be optimal in a given bandit instance.
This explains why Algorithm 1 performs better: it closely aligned with the $q$-measure, whereas UCB and TS select actions more uniformly. Further, in any given round $\theta$ of TS/UCB, the reward $\max_{a\in\mathcal{A}}\mu_{a}$ for optimal and suboptimal super arms can be indistinguishable if both include the optimal base action $a^{*}(\theta)$. This ambiguity slows down the convergence of the learning algorithms because they cannot effectively differentiate between good and bad super arms—a problem exacerbated by the combinatorial number of possible super arms.
Figure 3: Comparison of representative actions selected by three algorithms in a synthetic setting. Outcome functions are generated using an RBF kernel (length-scale $l=1$) over 15 grid points in $[0,2]$. In each trial, $K=5$ distinct actions are selected per algorithm (Algorithm 1 in green, TS in yellow, and UCB in red); UCB and TS are run for 3000 rounds. Upper: Representative actions selected in three individual trials; each row shows one trial. Lower: Action selection frequencies of Algorithm 1, TS, and UCB over 50 trials. Algorithm 1 consistently selects boundary actions (indices 0 and 14), while UCB and TS select more uniformly.
Figure 4: Effect of kernel length-scale on importance measure $q$. Upper: Example outcome functions generated using RBF kernels with varying length-scales $l\in\{0.1,1,4\}$, over 15 grid points in $[0,2]$, with actions indexed from 0 to 14. Lower: Action selection frequencies of Algorithm 1 over 50 trials for each length-scale setting. These frequencies approximate the importance measure $q$. As the length-scale increases, the outcome functions become smoother and more linear, causing $q$-measure to concentrate at the boundary actions (indices 0 and 14).
Algorithm 2 Upper Confidence Bound with Gaussian Rewards
1:Input: Base actions $\mathcal{A}_{full}$, subset cardinality $K$, kernel matrix $\mathbf{K}$, time horizon $T=3000$
2:Initialize: Super arms $\{\mathcal{A}_{i}\}_{i\leq M}$ with $M=\binom{|\mathcal{A}_{full}|}{K}$, reward variance $\sigma^{2}=1$
3:Initialize: $N_{i}\leftarrow 0$, $\hat{\mu}_{i}\leftarrow 0$ for all $i=1,\dots,K$
4:for $t=1$ to $T+M$ do $\triangleright$ Each round corresponds to a bandit
5:     Sample an outcome function $f\sim\mathcal{N}(0,\mathbf{K})$
6:     if $t\leq M$ then
7:         Select arm $i_{t}=t$
8:     else
9:         Select arm $i_{t}=\arg\max_{i}\left(\hat{\mu}_{i}+\sqrt{\frac{2\log t}{N_{i}+\epsilon}}\right)$
10:     end if
11:     Pull arm $i_{t}$, observe reward $r_{t}=\max_{a\in\mathcal{A}_{i_{t}}}f(a)$
12:     Update:
$N_{i_{t}}\leftarrow N_{i_{t}}+1$
$\hat{\mu}_{i_{t}}\leftarrow\hat{\mu}_{i_{t}}+\frac{r_{t}-\hat{\mu}_{i_{t}}}{N_{i_{t}}}$
13:end for
Algorithm 3 Thompson Sampling with Gaussian Priors
1:Input: Base actions $\mathcal{A}_{full}$, subset cardinality $K$, kernel matrix $\mathbf{K}$, time horizon $T=3000$
2:Initialize: Super arms $\{\mathcal{A}_{i}\}_{i\leq M}$ with $M=\binom{|\mathcal{A}_{full}|}{K}$, reward variance $\sigma^{2}=1$
3:Initialize: prior mean $\mu_{i}\leftarrow 0$, prior precision $\tau_{i}\leftarrow\epsilon$ for all $i=1,\dots,K$
4:for $t=1$ to $T$ do $\triangleright$ Each round corresponds to a bandit
5:     Sample an outcome function $f\sim\mathcal{N}(0,\mathbf{K})$
6:     for $i=1$ to $M$ do
7:         Sample $\tilde{\mu}_{i}\sim\mathcal{N}(\mu_{i},\sigma^{2}/\tau_{i})$
8:     end for
9:     Select arm $i_{t}=\arg\max_{i}\tilde{\mu}_{i}$
10:     Pull arm $i_{t}$, observe reward $r_{t}=\max_{a\in\mathcal{A}_{i_{t}}}f(a)$
11:     Update:
$\tau_{i_{t}}\leftarrow\tau_{i_{t}}+1$
$\mu_{i_{t}}\leftarrow\mu_{i_{t}}+\frac{r_{t}-\mu_{i_{t}}}{\tau_{i_{t}}}$
12:end for
Appendix K Properties of Gaussian width
Given a set $\mathcal{S}\in\mathbb{R}^{n}$, the term $\mathbb{E}[\max_{a\in\mathcal{S}}\mu_{a}]$ where $\theta\sim\mathcal{N}(0,I)$ is called Gaussian (mean) width.
Claim 1: $\mathbb{E}\left[\max_{a\in\mathcal{S}}\langle a,\theta\rangle\right]=\mathbb{E}\left[\max_{a^{\prime}\in-\mathcal{S}}\langle a^{\prime},\theta\rangle\right]$.
$\mathbb{E}\left[\max_{a\in\mathcal{S}}\langle a,\theta\rangle\right]=\mathbb{E}\left[\max_{a\in\mathcal{S}}\langle a,-\theta\rangle\right]=\mathbb{E}\left[\max_{a\in\mathcal{S}}\langle-a,\theta\rangle\right]=\mathbb{E}\left[\max_{a^{\prime}\in-\mathcal{S}}\langle a^{\prime},\theta\rangle\right],$
where the first equality uses $-\theta$ and $\theta$ are identically distributed. The third equality uses for any $a\in\mathcal{S}$, it holds that $-a\in-\mathcal{S}$.
Claim 2: $\mathbb{E}\left[\max_{a\in\mathcal{S}}\langle a,\theta\rangle\right]\leq\frac{1}{2}\mathbb{E}\left[\max_{a,a^{\prime}\in\mathcal{S}}\langle a-a^{\prime},\theta\rangle\right]$.
Let $a^{*}(-\mathcal{S},\theta)$ denote the optimal action in $\mathcal{S}$ for bandit instance $\theta$.
$2\cdot\mathbb{E}\left[\max_{a\in\mathcal{S}}\langle a,\theta\rangle\right]=$ $\mathbb{E}\left[\max_{a\in\mathcal{S}}\langle a,\theta\rangle\right]+\mathbb{E}\left[\max_{a^{\prime}\in-\mathcal{S}}\langle a^{\prime},\theta\rangle\right]$
$=$ $\mathbb{E}\left[\langle a^{*}(\mathcal{S},\theta),\;\theta\rangle\right]+\mathbb{E}\left[\langle a^{*}(-\mathcal{S},\theta),\;\theta\rangle\right]$
$=$ $\mathbb{E}\left[\langle a^{*}(\mathcal{S},\theta)+a^{*}(-\mathcal{S},\theta),\;\theta\rangle\right]$
$\leq$ $\mathbb{E}\left[\max_{a,a^{\prime}\in\mathcal{S}}\langle a-a^{\prime},\theta\rangle\right],$
where the first equality uses Claim 1, the second equality uses the definition of $a^{*}(-\mathcal{S},\theta)$. The third equality uses linearity of expectation. The inequality uses that the vector $a^{*}(\mathcal{S},\theta)+a^{*}(-\mathcal{S},\theta)$ belongs to the set of vectors $\{a-a^{\prime}:a,a^{\prime}\in\mathcal{S}\}$. In fact, one can prove the equality that $\mathbb{E}\left[\max_{a\in\mathcal{S}}\langle a,\theta\rangle\right]=\frac{1}{2}\mathbb{E}\left[\max_{a,a^{\prime}\in\mathcal{S}}\langle a-a^{\prime},\theta\rangle\right]$, but we only need this inequality to prove the next claim.
Claim 3: $\mathbb{E}\left[\max_{a\in\mathcal{S}}\mu_{a}\right]\leq\frac{\operatorname{diam}(\mathcal{S})}{2}\cdot\sqrt{n}$.
$\mathbb{E}\left[\max_{a\in\mathcal{S}}\mu_{a}\right]=$ $\mathbb{E}\left[\max_{a\in\mathcal{S}}\langle a,\theta\rangle\right]$
$\leq$ $\frac{1}{2}\mathbb{E}\left[\max_{a,a^{\prime}\in\mathcal{S}}\langle a-a^{\prime},\theta\rangle\right]$
$\leq$ $\frac{1}{2}\mathbb{E}\max_{a,a^{\prime}\in\mathcal{S}}\|\theta\|_{2}\|a-a^{\prime}\|_{2}$
$\leq$ $\frac{1}{2}\mathbb{E}\operatorname{diam}(\mathcal{S})\|\theta\|_{2}\leq\frac{\operatorname{diam}(\mathcal{S})}{2}\cdot\sqrt{n},$
where the first equality uses the definition of $\mu_{a}$. The first inequality uses Claim 2. The second inequality uses Cauchy-Schwarz inequality. The third inequality uses the definition of $\operatorname{diam}(\cdot)$. The last inequality uses $\mathbb{E}\|\theta\|_{2}\leq\sqrt{n}$.
Appendix L Bounds of expected maximum of Gaussian
Let $X_{1},\dots,X_{N}$ be $N$ random Gaussian variables (no necessarily independent) with zero mean and variance of marginals smaller than $\sigma^{2}$, then
$\mathbb{E}\left[\max_{i=1,\dots,N}X_{i}\right]\leq\sigma\sqrt{2\log N}.$
Proof.
for any $\delta>0$,
$\mathbb{E}\left[\max_{i=1,\dots,N}X_{i}\right]=$ $\frac{1}{\delta}\mathbb{E}\left[\log\exp(\delta\max_{i=1,\dots,N}X_{i})\right]\leq\frac{1}{\delta}\log\mathbb{E}\left[\exp(\delta\max_{i=1,\dots,N}X_{i})\right]$
$=$ $\frac{1}{\delta}\log\mathbb{E}\left[\max_{i=1,\dots,N}\exp(\delta X_{i})\right]\leq\frac{1}{\delta}\log\sum_{i=1}^{N}\mathbb{E}\left[\exp(\delta X_{i})\right]$
$\leq$ $\frac{1}{\delta}\log\sum_{i=1}^{N}\exp(\sigma^{2}\delta^{2}/2)=\frac{\log N}{\delta}+\frac{\sigma^{2}\delta}{2},$
where the first inequality uses Jensen’s inequality. Taking $\delta:=\sqrt{2(\log N)/\sigma^{2}}$ yields the results. ∎
Let $X_{1},\dots,X_{N}$ be i.i.d. $\mathcal{N}(0,\sigma^{2})$ random variables, then according to [28]:
$\mathbb{E}_{\theta}\left[\max_{i=1,\dots,N}X_{i}\right]\geq\frac{\sigma\sqrt{\log N}}{\sqrt{\pi\log 2}}.$