### v2

# Representative Action Selection for Large Action Space Meta-Bandits

Abstract
We study the problem of selecting a subset from a large action space shared by a family of bandits, with the goal of achieving performance nearly matching that of using the full action space. We assume that similar actions tend to have related payoffs, modeled by a Gaussian process. To exploit this structure, we propose a simple $\epsilon$-net algorithm to select a representative subset. We provide theoretical guarantees for its performance and compare it empirically to Thompson Sampling and Upper Confidence Bound.
1 Introduction
We study a family of bandits that share a common but extremely large action space. We aim to understand whether it is possible—and how—to select a smaller set of representative actions that performs nearly as well as the full action space across all bandit instances. To build intuition, imagine a pharmacy preparing its inventory for the upcoming season. The available drugs (actions) are nearly infinite, and each customer (bandit) has unique characteristics. If two drugs share similar ingredients, their effects on a patient are likely to be similar. Likewise, if two patients have comparable health indices, a drug is likely to have similar effects on both. By modeling the expected outcome of each drug for each patient as a Gaussian process, we can capture these correlations. Now, consider medicine demand: If a drug treats a very rare illness, it is unlikely to be needed frequently, so the store can exclude it to optimize storage space. Conversely, if flu season is approaching, stocking several flu medications is a wise choice.
Different from prior approaches in multi-armed bandits (MAB) (Dani et al., 2008) that aim for identifying either a single best action or a subset that achieves high cumulative outcomes for a fixed bandit, our objective focuses on selecting a subset that is likely to contain the best action, or one whose best element performs nearly as well for a family of bandits. This problem can be seen as a large-scale combinatorial optimization under uncertainty, with applications where decisions involve a vast number of possibilities but are constrained by computational or time limitations for evaluating all options, e.g., inventory management, online recommendations.
Consider the following MAB setting: In a bandit, if a decision-maker plays an action with a fixed but unknown feature vector $a\in\mathcal{A}_{\operatorname*{full}}\subset\mathbb{R}^{n}$, they observe a random outcome taking values in $\mathbb{R}$. We define the expected outcome of playing action $a$ in this bandit as $\mu_{a}(\theta):=\langle a,\theta\rangle$ where this bandit instance $\theta\in\mathbb{R}^{n}$ is drawn from an unknown multivariate Gaussian distribution. Thus, the collection of random variables $\{\mu_{a}\}_{a\in\mathcal{A}_{\operatorname*{full}}}$ forms a Gaussian process (GP) (Vershynin, 2018, Chapter 7), and we also provide results under general sub-Gaussian assumption. Our underlying setting aligns with the one considered in contextual bandit (Dani et al., 2008); however, the key difference lies in the objective.
Now consider that the decision-maker has access to a fixed action subset $\mathcal{A}\subset\mathcal{A}_{\operatorname*{full}}$. For a given bandit instance $\theta$, if the best action over the full space lies within $\mathcal{A}$, the decision-maker benefits from reduced suboptimal actions to explore. Conversely, if the best action lies outside $\mathcal{A}$, regret arises from being unable to select this best action. Thus, we define the regret as the difference in expected outcome between having access to the full action space versus being restricted to the subset:
$\boxed{\textnormal{Regret}:=\max_{a\in\mathcal{A}_{\operatorname*{full}}}\mu_{a}-\max_{a^{\prime}\in\mathcal{A}}\;\mu_{a^{\prime}},}$ (1)
which depends on the sampled bandit instance and is therefore a random variable. Our objective is to identify a small subset $\mathcal{A}$ that minimizes the expected regret $\mathbb{E}_{\theta}\left[\operatorname*{Regret}\right]$ over all possible bandit instances, making the underlying optimization both stochastic and combinatorial.
This objective is motivated by practical considerations. The classic Bayesian regret in the bandit literature (Agrawal and Goyal, 2012) typically scales with the number of available actions. However, if a subset $\mathcal{A}$ is carefully chosen, the resulting Bayesian bandit regret can be significantly lower due to the reduced number of suboptimal actions. This is especially beneficial when the action space is large, as even the initialization phase can be computationally expensive. To see this, we can decompose the bandit regret as follows:
[…]
$=\mathbb{E}\sum_{t=1}^{N}\left[\max_{a\in\mathcal{A}}\mu_{a}(\theta)\!-\!\mu_{A_{t}}+\max_{a\in\mathcal{A}_{\operatorname*{full}}}\mu_{a}(\theta)\!-\!\max_{a\in\mathcal{A}}\mu_{a}(\theta)\right]$
$\leq C\sqrt{|\mathcal{A}|\cdot N\log N}+N\cdot\mathbb{E}_{\theta}[\textnormal{Regret}].$
In the first equality, $A_{t}\in\mathcal{A}$ denotes the action chosen by a policy, e.g., Thompson Sampling (Agrawal and Goyal, 2012), in round $t$, and $N$ is the number of rounds. The expectation is taken over the randomness in the distribution of bandit instances and in actions selected by the policy. The inequality follows from the well-known regret bounds for Thompson Sampling (Lattimore and Szepesvári, 2020). $|\mathcal{A}|$ denotes the cardinality of the action set and $C>0$ is a constant. Note that if the policy has access to the full action space, the Bayesian bandit regret is instead bounded by $C\sqrt{|\mathcal{A}_{\operatorname*{full}}|\cdot N\log N}$.
Our main contributions are organized as follows:
  • Meta-Bandits Framework. We propose a meta-bandits framework that specifically tackles combinatorial action selection by leveraging correlations across similar actions and bandit instances. To the best of our knowledge, this is the first such framework.
  • $\epsilon$-Net Algorithm. We introduce a simple algorithm within this framework. It starts with the intuitive idea of placing a grid over the action space, then refines it using an importance-based selection mechanism.
  • Regret Analysis. We provide theoretical guarantees for both the grid and the algorithm’s output, including upper and lower bounds on expected regret, along with results under general sub-Gaussian processes. We also discuss the cost of not using a grid, which depends on the importance-structure of the action space.
  • Generalization and Empirical Validation. We extend the analysis to settings where outcome functions are sampled from a reproducing kernel Hilbert space (RKHS), and empirically compare our algorithm to Thompson Sampling (TS) and Upper Confidence Bound (UCB).
1.1 Related Works
Multi-Armed Bandits (Lai and Robbins, 1985; Auer et al., 2002a) is defined by a set of actions (arms), each deliver outcomes that are independently drawn from a fixed and unknown distribution. The decision-maker sequentially selects an action, observes its outcome, and aims to maximize cumulative outcomes over time. Popular methods include the UCB (Auer et al., 2002a), TS (Agrawal and Goyal, 2012), and EXP3 (Auer et al., 2002b) for adversarial settings.
Optimal Action Identification focuses on identifying the action with the highest expected outcome in a MAB setting using as few independent samples as possible (Jamieson and Nowak, 2014; Kaufmann et al., 2016). Popular methods in fixed confidence setting include Action Elimination Even-Dar et al. (2006); Karnin et al. (2013), UCB, and LUCB, all of which achieve sample complexity within a $\log(|\mathcal{A}_{\operatorname*{full}}|)$ factor of the optimum. In fixed budget setting, there is Successive halving Karnin et al. (2013), successive reject Audibert and Bubeck (2010). Our algorithm includes a step for identifying the optimal action. While this line of work assumes (non-centered) sub-Gaussian outcome distributions within a bandit instance, we model each instance with a linear structure, allowing for further simplification.
Stochastic Linear Optimization assumes that the expected outcome of each action depends through the inner product between a context $\theta$ and an action $a\in\mathcal{A}_{\operatorname*{full}}$ Auer (2002); Dani et al. (2008); Rusmevichientong and Tsitsiklis (2010), and our model of each bandit instance align with this setting. This is a sequential method for maximizing a linear function—i.e., $f(a):=\mu_{a}(\theta)$ with fixed $\theta$—based on noisy observations of the function values. Soare et al. (2014) study the sample complexity of optimal action identification in this setting. However, this line of work assumes the action feature vectors are known.
GP Optimization can deal with the case that the feature vectors of actions are unknown. Srinivas et al. (2009) models the outcome function as a sample from a GP prior with a kernel function (Williams and Rasmussen, 2006). Their Bayesian algorithm, GP-UCB, uses this GP prior to first select points $x_{t}$ that enable global estimation of the function, and then plays the point with the highest posterior mean. The widespread adoption of this method in bandit settings (Valko et al., 2013; Li and Scarlett, 2022), as well as in continuous action spaces (Chowdhury and Gopalan, 2017), highlights the practicality of assuming that action outcomes are correlated. This also motivates our extension, where the outcome function of each bandit instance is modeled as a sample from a RKHS.
Before discussing two seemingly relevant lines of work that select a subset from the action space, we first highlight a advantage of our approach. These methods assume a fixed $|\mathcal{A}|=K$, which is often unclear in practice and requires restarting the algorithm when changed. In contrast, our algorithm can adapt the subset size on the fly. For example, if 100 runs all return the same action, choosing this one may suffice ($|\mathcal{A}|=1$); if each run yields a different action, we may need more iterations to ensure a small regret ($|\mathcal{A}|>100$).
[…]
Example 1.
Consider three actions: $a_{1}=[1,0]$, $a_{2}=[0.9,0.1]$, and $a_{3}=[-0.1,1]$, and suppose bandits are sampled uniformly from $\theta_{1}=[1,0]$ and $\theta_{2}=[0,1]$. Then,
$\mathbb{E}\langle a_{1},\theta\rangle=\mathbb{E}\langle a_{2},\theta\rangle=0.5,\quad\mathbb{E}\langle a_{3},\theta\rangle=0.45.$
So under the Best-222-Action perspective, $a_{1}$ and $a_{2}$ would be selected. However, this is suboptimal in our framework, since $\mu_{a_{1}}$ and $\mu_{a_{2}}$ are positively correlated:
$\mathbb{E}\max_{a\in\{a_{1},a_{2}\}}\langle a,\theta\rangle=0.55,\quad\mathbb{E}\max_{a\in\{a_{1},a_{3}\}}\langle a,\theta\rangle=1.$
In contrast, our algorithm, if run until it selects two distinct actions, would output $a_{1}$ and $a_{3}$, which is the true optimal.
Combinatorial Bandits considers that the decision maker selects $K$ of base arms from $\mathcal{A}_{\operatorname*{full}}$ in each round, forming a super arm $\mathcal{A}$, with $|\mathcal{A}|=K$. Popular methods include CUCB Chen et al. (2016), CTS Wang and Chen (2018). We argue that this line of work is not applicable to our framework: (1) It assumes that the expected outcome of a super arm depends only on the expected outcomes of its individual base arms, or imposes a stricter monotonicity condition. In our case, even though $\mathbb{E}[\mu_{a}]=c$ for all $a\in\mathcal{A}_{\operatorname*{full}}$, super arm expected outcomes $\mathbb{E}[\max_{a\in\mathcal{A}}\mu_{a}]$ can differ significantly due to correlations among actions. (2) It assumes independence across base arms, whereas we explicitly model correlations. Ignoring these correlations misses the core challenge—an issue illustrated in Example 1.
[…]
Expected supremum of Gaussian process for a given set $\mathcal{S}$ refers to the term $\mathbb{E}\left[\max_{a\in\mathcal{S}}\mu_{a}\right]$. It is an important topic in high-dimensional probability (Vershynin, 2018). The sharpest known bounds are due to Talagrand (2014).
2 Subset Selection Framework
We consider the problem of selecting a small number of representative actions from a large action space $\mathcal{A}_{\operatorname*{full}}\subset\mathbb{R}^{n}$, where $n\in\mathbb{N}$. (This framework applies to the case $n=+\infty$, with the additional assumption $\sum_{i\geq 1}a^{2}_{i}\!<\!\infty$.) The expected outcome depends on both the chosen action $a\in\mathbb{R}^{n}$ and an observed context $g\in\mathbb{R}^{n}$, and includes a constant $c\in\mathbb{R}$:
$\mu_{a}:=\langle a,g\rangle+c,\quad\forall,a\in\mathcal{A}_{\operatorname*{full}},;g\sim\mathcal{N}(0,\Sigma),$ (2)
where $\Sigma$ is a positive semi-definite matrix. Let $\theta\in\mathbb{R}^{n}$ follows a multivariate normal distribution with zero mean and identity covariance matrix. The distribution of $g$ is in fact equivalent to $\Sigma^{1/2}\theta$. Now, let $\sigma_{j}$ denote the $j$-row of the matrix $\Sigma^{1/2}$. We have $g=(\langle\sigma_{j},\theta\rangle)_{j\leq n}$ and
$\langle a,g\rangle=\sum_{j\leq n}a_{j}\langle\sigma_{j},\theta\rangle=\left\langle\sum_{j\leq n}a_{j}\sigma_{j},\theta\right\rangle=\langle\Sigma^{1/2}a,\theta\rangle.$
Therefore, the setting in Equation (2) is equivalent to
$\mu_{a}:=\langle a,\theta\rangle+c,\;\forall\;a\in\Sigma^{1/2}\mathcal{A}_{\operatorname*{full}},\;\theta\sim\mathcal{N}(0,I),$
where $\Sigma^{1/2}\mathcal{A}_{\operatorname*{full}}$ denotes the image of $\mathcal{A}_{\operatorname*{full}}$ under the linear transformation $\Sigma^{1/2}$. Since the constant $c$ does not affect the regret (as defined in Equation (1)), we can, without loss of generality, focus on this canonical Gaussian process (Vershynin, 2018, Chapter 7) in the remainder:
$\boxed{\mu_{a}(\theta):=\langle a,\theta\rangle,\quad\forall a\in\mathcal{A}_{\operatorname*{full}},\theta\sim\mathcal{N}(0,I).}$ (3)
We define the extreme points as those $x\in\mathcal{A}_{\operatorname*{full}}$ for which there do not exist distinct $a,a^{\prime}\in\mathcal{A}_{\operatorname*{full}}$ and $\lambda\in(0,1)$ such that $x=\lambda a+(1-\lambda)a^{\prime}$. By the extreme point theorem, if we select all extreme points—denoted $\mathcal{A}=\{a_{1},\dots,a_{K}\}$—as representatives of the full action space, the regret is zero. This is because any $a\in\mathcal{A}_{\operatorname*{full}}$ can be expressed as a convex combination of the extreme points: $a=\lambda_{1}a_{1}+\dots+\lambda_{K}a_{K}$, where $\lambda_{i}\geq 0$ and $\sum_{i=1}^{K}\lambda_{i}=1$. Thus, for any $\theta\in\mathbb{R}^{n}$:
$\begin{split}\langle a,\theta\rangle=\sum_{i=1}^{K}\lambda_{i}\langle a_{i},\theta\rangle\leq\sum_{i=1}^{K}\lambda_{i}\max_{a^{\prime}\in\mathcal{A}}\langle a^{\prime},\theta\rangle=\max_{a^{\prime}\in\mathcal{A}}\mu_{a^{\prime}}.\end{split}$
By Equation (3), it yields
$\max_{a^{\prime}\in\mathcal{A}}\mu_{a^{\prime}}\leq\max_{a\in\mathcal{A}_{\operatorname*{full}}}\mu_{a}\leq\max_{a^{\prime}\in\mathcal{A}}\mu_{a^{\prime}},$
where the left inequality uses $\mathcal{A}\subseteq\mathcal{A}_{\operatorname*{full}}$. Thus, the two quantities $\max_{a\in\mathcal{A}}\mu_{a}$ and $\max_{a\in\mathcal{A}_{\operatorname*{full}}}\mu_{a}$ are equal.
This example highlights how a geometric approach can be used to solve the stochastic combination problem. However, even if one only needs the extreme points, the set of extreme points may still be large, e.g, the extreme points of a Euclidean ball is infinite. To address this, we will later introduce the notions of $\epsilon$-nets. Without loss of generality, we assume $\mathcal{A}_{\operatorname*{full}}$ consists only of the extreme points of $\mathcal{A}_{\operatorname*{full}}$, as they are the only points of interest.
2.1 Epsilon Nets
If the set of extreme points $\mathcal{A}_{\operatorname*{full}}$ is still large, a natural approach is to construct a grid over the action space, where the grid points serve as representative actions. This ensures that for every action in the full space, there exists a representative that is close to it. This idea is formally captured by the notion of a (geometric) $\epsilon$-net.
To proceed, we clarify what we mean by an $\epsilon$-net, as there are at least two definitions: one from a geometric perspective (Vershynin, 2018, Chapter 4) and another from a measure-theoretic perspective (Matousek, 2013, Chapter 10). Let $\|\cdot\|_{2}$ denote the Euclidean norm. Define the diameter of a compact set $r\in\mathbb{R}^{n}$ as $\mathrm{diam}(r):=\max_{a,b\in r}\|a-b\|_{2}$.
  • A subset $\mathcal{A}\subseteq\mathcal{A}_{\operatorname*{full}}$ is called a Geometric $\epsilon$-net if, for all $a\in\mathcal{A}_{\operatorname*{full}}$, there exists $a^{\prime}\in\mathcal{A}$ such that
    $\|a-a^{\prime}\|_{2}<\epsilon.$
  • Let $\mathcal{R}$ be a finite partition of the extreme points into disjoint clusters such that $\cup_{r\in\mathcal{R}}\;r=\mathcal{A}_{\operatorname*{full}}$. Given a measure $q$ assigning a value to each cluster $r\in\mathcal{R}$. A subset $\mathcal{A}\subseteq\mathcal{A}_{\operatorname*{full}}$ is called a Measure-Theoretic $\epsilon$-net with respect to measure $q$ if, for any cluster $r\in\mathcal{R}$, we have:
    $r\cap\mathcal{A}\neq\emptyset\quad\text{whenever}\quad q(r)>\epsilon.$
A geometric $\epsilon$-net ensures small regret because if two actions $a,a^{\prime}\in\mathcal{A}_{\operatorname*{full}}$ are close in the Euclidean sense, then the deviation between $\mu_{a}$ and $\mu_{a^{\prime}}$ is small in the $L^{2}$-sense (i.e., their expected squared difference is small):
$\|\mu_{a}\!-\!\mu_{a^{\prime}}\|_{L^{2}}=\!\left(\mathbb{E}(a\!-\!a^{\prime})^{\top}\theta\theta^{\top}(a\!-\!a^{\prime})\right)^{1/2}\!=\|a\!-\!a^{\prime}\|_{2},$
where $\theta^{\top}$ denotes the transpose of $\theta$. The equalities use Equation (3) and $\mathbb{E}\theta\theta^{\top}=I$. Therefore, by definition, a geometric $\epsilon$-net guarantees the existence of an action $a\in\mathcal{A}$ whose expected outcome $\mu_{a}$ is close to that of the optimal action for any given bandit instance. However, this net suffers from the curse of dimension: e.g., for $[0,1]^{n}$, the number of points needed to form a geometric $\epsilon$-net grows as $(1/\epsilon)^{n}$.
The measure-theoretic $\epsilon$-net addresses this issue. Put simply, the measure-theoretic $\epsilon$-net restricts the grid construction to only the most important clusters $r\in\mathcal{R}$, as determined by the $q$-measure.
2.2 Epsilon Net Algorithm
We propose Algorithm 1, a variant of the $\epsilon$-net algorithm originally introduced by Haussler and Welzl (1986). It selects $K$ i.i.d. random actions, aligned with the distribution of bandit instances. Since repetitions are allowed, the resulting subset $\mathcal{A}$ may have fewer than $K$ distinct actions.
Algorithm 1 Epsilon Net Algorithm
1:Input: Action space $\mathcal{A}_{\operatorname*{full}}$, Sample size $K$.
2:Output: A subset of actions $\mathcal{A}$.
[…]
4:for $1,\dots,K$ do
5:  Sample a bandit instance $\theta$.
6:  Find optimal action $a^{*}(\theta)\!:=\!\operatorname*{argmax}_{a\in\mathcal{A}_{\operatorname*{full}}}\langle a,\theta\rangle$.
7:  $\mathcal{A}\leftarrow\mathcal{A}\cup\{a^{*}\}$triangleright Repetition of actions is allowed
8:end for
We define the optimal action in a bandit instance $\theta$ as
$a^{*}(\theta):=\operatorname*{argmax}_{a\in\mathcal{A}_{\operatorname*{full}}}\mu_{a}(\theta).$
Assumption 2 (Unique optimal action).
The optimal action $a^{*}(\theta)$ is unique with probability 111 for all bandit instances.
Define the Importance Measure $q$ over a partition $\mathcal{R}$:
$\boxed{q(r):=\Pr[a^{*}(\theta)\in r]=\!\int\!\mathbb{1}\{a^{*}(\theta)\in r\}\;p(\theta)d\theta,}$ (4)
[…]
Assumption 3.
The support of measure $q$ is compact.
The compactness assumption ensures that the term $\mathbb{E}_{\theta}\left[\max_{a\in\mathcal{A}_{\operatorname*{full}}}\mu_{a}\right]$ is finite and guarantees the attainment of a unique optimal action. Without loss of generality, we assume that $\mathcal{A}_{\operatorname*{full}}$ is the support of the measure $q$.
By definition, Algorithm 1 samples $K$ i.i.d. extreme points from clusters in $\mathcal{R}$ according to measure $q$. If a cluster $r\in\mathcal{R}$ has a higher measure $q(r)$, its elements are more likely to be included in the output. In fact, with high probability, this algorithm outputs a measure-theoretic $\epsilon$-net of $\mathcal{A}_{\operatorname*{full}}$ with respect to measure $q$. (This is a simplified version of Theorem 10.2.4 of Matousek (2013).)
Lemma 4.
Given a partition $\mathcal{R}$ of the full action space, and the importance measure $q$ assigning a value to each cluster $r\in\mathcal{R}$. Let $\mathcal{A}$ be the output of Algorithm 1 after $K$ samples. Then, with probability at least $1-\frac{1}{\epsilon}\exp(-K\epsilon)$, it holds that for any cluster $r\in\mathcal{R}$,
$r\cap\mathcal{A}\neq\emptyset\quad\text{whenever}\quad q(r)>\epsilon.$
The partition $\mathcal{R}$ bridges the two definitions of $\epsilon$-net: choosing one point from each cluster gives an $\epsilon$-net in both senses, though with different values of $\epsilon$. Geometrically, $\epsilon$ is the largest cluster diameter $\epsilon:=\max_{r\in\mathcal{R}}\mathrm{diam}(r)$; measure-theoretically, $\epsilon$ is the smallest cluster measure $\epsilon:=\min_{r\in\mathcal{R}}q(r)$. Hence, we have the following corollary:
Corollary 5.
Under the same conditions as Lemma 4, let $\epsilon:=\min_{r\in\mathcal{R}}q(r)$. Then, with probability at least $1-\frac{1}{\epsilon}\exp(-K\epsilon)$, the following holds:
$\forall a\in\mathcal{A}_{\operatorname*{full}},\ \exists a^{\prime}\in\mathcal{A}\text{ such that }\|a-a^{\prime}\|_{2}<\max_{r\in\mathcal{R}}\operatorname{diam}(r).$
3 Regret Analysis
In this section, we begin by analyzing a special class of geometric $\epsilon$-nets, constructed by partitioning the action space into clusters and selecting a single representative action from each cluster. We then extend the analysis to obtain algorithm-dependent bounds for the output of Algorithm 1.
Definition 1 (Reference subsets).
Consider a partition :=$\mathcal{R}:=\{r_{\ell}\}_{\ell\leq m}$ of the full action space, with $\epsilon:=\max_{r\in\mathcal{R}}\operatorname{diam}(r)$. A reference subset is a set $\mathcal{A}:=\{a_{1},\dots,a_{m}\}$, where each representative $a_{\ell}\in\mathcal{A}$ corresponds to a cluster $r_{\ell}$, and each cluster $r_{\ell}$ is contained within a closed Euclidean ball of radius $\epsilon$ centered at $a_{\ell}$, i.e., $r_{\ell}\subset B(a_{\ell},\epsilon)$.
We impose the following assumption on the partition, used only for lower bounds: if the optimal action lies in cluster $r$, then all actions in $r$ outperform those outside it. This ensures clusters are well-separated.
Assumption 6 (Effective partition).
For any $r\in\mathcal{R}$, whenever the optimal action lies in cluster $r$, i.e., $a^{*}(\theta)\in r$, then $\mu_{a}\geq\max_{a^{\prime}\in\mathcal{A}_{\operatorname*{full}}\setminus r}\mu_{a^{\prime}},\;\forall a\in r$.
Theorem 7 (Regret bounds of reference subsets).
Consider a partition :=$\mathcal{R}:=\{r_{\ell}\}_{\ell\leq m}$ of the full action space, with $\epsilon:=\max_{r\in\mathcal{R}}\operatorname{diam}(r)$, and an arbitrary reference subset $\mathcal{A}$. Then, there is an absolute constant $C>0$, such that
$\mathbb{E}_{\theta}[\operatorname*{Regret}]\leq\max_{r\in\mathcal{R}}\mathbb{E}_{\theta}\left[\max_{a\in r}\mu_{a}\right]+C\epsilon\sqrt{\log|\mathcal{R}|}.$
If the partition $\mathcal{R}$ satisfies Assumption 6, then
$\mathbb{E}_{\theta}[\operatorname*{Regret}]\geq\min_{r\in\mathcal{R}}\mathbb{E}_{\theta}\left[\max_{a\in r}\mu_{a}\right]-C\epsilon\sqrt{\log|\mathcal{R}|}.$
Proof sketch. For each cluster $\ell\leq m$, define a simple Gaussian process $\left\{Z_{a}\right\}_{a\in r_{\ell}}$ where $Z_{a}:=\mu_{a}-\mu_{a_{\ell}}$. Define a non-negative random variable $Y_{\ell}:=\sup_{a\in r_{\ell}}Z_{a}$. When $a^{*}(\theta)\in r_{\ell}$, regret is upper bounded by $Y_{\ell}$ (or equal to it under Assumption 6). Thus, for any bandit $\theta$, the regret is bounded between $\min_{\ell\leq m}Y_{\ell}$ and $\max_{\ell\leq m}Y_{\ell}$. Finally, the expectations $\mathbb{E}\left[\min_{\ell\leq m}Y_{\ell}\right]$ and $\mathbb{E}\left[\max_{\ell\leq m}Y_{\ell}\right]$ can be bounded via concentration property of Gaussian process. $\square$
3.1 Regret Bounds of Algorithm
The algorithm’s regret bound is established by comparing it to that of a reference subset, for which we already have known expected regret bounds. The key difference is that whereas the reference subset includes a representative from each cluster, the algorithm may miss some clusters. However, although it may select a smaller subset, the algorithm achieves regret comparable to that of the reference subset, as it tends to miss clusters that contribute minimally to the expected regret.
The expected regret in previous results is taken over bandit instances $\theta$. In contrast, since the output of Algorithm 1 is random, the expected regret analyzed in this section is taken with respect to both the algorithm’s randomness (i.e., the sampled $\mathcal{A}$) and the distribution over $\theta$.
Upper Bound.
Theorem 8.
Consider a partition :=$\mathcal{R}:=\{r_{\ell}\}_{\ell\leq m}$ of the action space, with $\epsilon:=\max_{r\in\mathcal{R}}\operatorname{diam}(r)$. Let $\mathcal{A}$ be output of Algorithm 1. For the same constant $C>0$ in Theorem 7,
$\begin{split}\mathbb{E}_{\theta,\mathcal{A}}[\textnormal{Regret}]\leq&\max_{r\in\mathcal{R}}\mathbb{E}_{\theta}\left[\max_{a\in r}\mu_{a}\right]+C\epsilon\sqrt{\log|\mathcal{R}|}\\ &+\left(\mathbb{E}_{q}\left[(1-q(r))^{2K}\right]\cdot\mathbb{E}_{\theta}\left[\max_{a\in\mathcal{A}_{\operatorname*{full}}}\mu^{2}_{a}\right]\right)^{1/2}.\end{split}$
Proof sketch. If $r_{\ell}\cap\mathcal{A}\neq\emptyset$, choose $a_{\ell}\in r_{\ell}\cap\mathcal{A}$ as the representative point. If $r_{\ell}\cap\mathcal{A}=\emptyset$, choose an arbitrary point $a_{\ell}\in r_{\ell}$. The new set $\mathcal{A}^{\prime}:=\{a_{\ell}\}_{\ell\leq m}$ forms a reference subset. Then, for each $\ell\leq m$, define a simple Gaussian process $\{Z_{a}\}_{a\in r_{\ell}}$, where $Z_{a}:=\mu_{a}-\mu_{a_{\ell}}$, and let $Y_{\ell}:=\sup_{a\in r_{\ell}}Z_{a}$. When $a^{*}(\theta)\in r_{\ell}$, we consider two cases:
  • $r_{\ell}\cap\mathcal{A}\neq\emptyset$: The regret is upper bounded by $Y_{\ell}$, and hence by $\max_{\ell\leq m}Y_{\ell}$.
  • $r_{\ell}\cap\mathcal{A}=\emptyset$: The regret is bounded by $\max_{a\in\mathcal{A}_{\operatorname*{full}}}\!\mu_{a}$. $\square$
The term $\mathbb{E}_{q}\left[(1-q(r))^{2K}\right]$ in Theorem 8 behaves similarly to entropy: it is maximized when $q$ is uniform, provided $|\mathcal{R}|\geq 2K+1$ (since we can choose any partition); see Lemma 18 in Supplementary. Intuitively, each cluster $r$ contributes $q(r)\cdot(1-q(r))^{2K}$. This expression is small when $q(r)$ is small, and decays faster when $q(r)$ gets larger, making the overall term negligible if $q$ is highly concentrated.
The connection between Theorem 7 and 8 is insightful:
Remark 9.
Consider placing a grid over the action space and defining a partition $\mathcal{R}$ by assigning each point in the space to its nearest grid point. In this way, the grid acts as a reference subset with respect to the partition $\mathcal{R}$. The regret upper bound in Theorem 7 applies to this grid, as it holds for any reference subset. Meanwhile, Theorem 8 applies to any partition, including $\mathcal{R}$. As a result, we obtain a regret bound for Algorithm 1 that exceeds the grid’s bound by only one additional term:
$\left(\mathbb{E}_{q}\left[(1-q(r))^{2K}\right]\cdot\mathbb{E}_{\theta}\left[\max_{a\in\mathcal{A}_{\operatorname*{full}}}\mu^{2}_{a}\right]\right)^{1/2},$
where the part $\mathbb{E}_{\theta}\left[\max_{a\in\mathcal{A}_{\operatorname*{full}}}\mu^{2}_{a}\right]$ is simply a constant, as the action space $\mathcal{A}_{\operatorname*{full}}$ is fixed.
If the distribution $q$ is highly concentrated, then the expectation $\mathbb{E}_{q}[(1-q(r))^{2K}]$ is small, making the extra term negligible. In this case, the benefit of constructing an explicit grid—often a non-trivial task when the action feature vectors are unknown or high-dimensional—is limited.
On the other hand, if the measure $q$ is close to uniform, then $\min_{r\in\mathcal{R}}q(r)$ is large. In this case, with relatively high probability, Algorithm 1 outputs a geometric $\max_{r\in\mathcal{R}}\mathrm{diam}(r)$-net (see Corollary 5). Moreover, since the partition is constructed using a grid, the cluster diameters are similar, so $\max_{r\in\mathcal{R}}\mathrm{diam}(r)$ closely matches the spacing of the original grid—making the output resemble a slightly coarser grid.
Partition-Independent Upper Bound.
Generally speaking, the partition $\mathcal{R}$ may be unknown, and while the corresponding $q$ measure exists, it remains unspecified. We therefore provide a worst-case bound using the covering number $N(\mathcal{A}_{\operatorname*{full}},\epsilon)$, which is the smallest number of points needed to form a geometric $\epsilon$-net of $\mathcal{A}_{\operatorname*{full}}$:
$N(\mathcal{A}_{\operatorname*{full}},\!\epsilon)=\min\Big{\{}m\!\in\!\mathbb{N}:$$\;\exists\{a_{\ell}\}_{\ell\leq m}\subseteq\mathcal{A}_{\operatorname*{full}},$
$\forall a\in\mathcal{A}_{\operatorname*{full}},\exists\ell,\|a\!-\!a_{\ell}\|_{2}\leq\epsilon\Big{\}}.$
Theorem 10.
Under Assumption 3, there exists a point $a_{0}$ and a constant $M>0$ such that $\mathcal{A}_{\operatorname*{full}}\subset B(a_{0},M)$, a closed Euclidean ball of radius $M$ centered at $a_{0}$. Let the action space have dimension $n$, and fix a constant $0<\epsilon<M$. Let $\mathcal{A}$ be the output of Algorithm 1. For the same constant $C>0$ in Theorem 7, and another absolute constant $c>0$,
$\mathbb{E}_{\theta,\mathcal{A}}[\operatorname*{Regret}]\leq\;$ $2\epsilon\sqrt{n}+C\epsilon\sqrt{\log N(\mathcal{A}_{\operatorname*{full}},\epsilon)},$
$\text{where }K\geq$ $c\cdot(M/\epsilon)^{2}\cdot N(\mathcal{A}_{\operatorname*{full}},\epsilon).$
As $\epsilon\to 0^{+}$, we have $\mathbb{E}_{\theta,\mathcal{A}}[\operatorname*{Regret}]\to 0$ as $K\to\infty$.
[…]
Lower Bound.
Theorem 11.
Under the same condition of Theorem 8. Let each cluster contains more than one action and satisfies Assumption 6. For the same constant $C$ in Theorem 7 and another absolute constant $c>0$,
$\begin{split}\mathbb{E}_{\theta,\mathcal{A}}[\operatorname*{Regret}]\geq&\left(\min_{r\in\mathcal{R}}\mathbb{E}_{\theta}\left[\max_{a\in r}\mu_{a}\right]-C\epsilon\sqrt{\log|\mathcal{R}|}\right)\\ &\times c\cdot\left(\mathbb{E}_{q}\left[(1-q(r))^{2K}\right]\right)^{1/2}.\end{split}$
3.2 Regret Bound for Sub-Gaussian Process
In fact, the Gaussian process assumption can be relaxed to a sub-Gaussian one, where the random process $\{\mu_{a}\}_{a\in\mathcal{S}}$ satisfies the following increment condition:
$\forall u>0,\quad\Pr\left[|\mu_{a}-\mu_{a^{\prime}}|\geq u\right]\leq 2\exp\left(-\frac{u^{2}}{2\|a-a^{\prime}\|_{2}^{2}}\right).$
Let $\gamma_{2}(\mathcal{S})$ denote Talagrand’s chaining functional (Talagrand, 2014) for a set $\mathcal{S}\subset\mathbb{R}^{n}$, $n\in\mathbb{N}$, equipped with the Euclidean norm. It is the sharpest known bound that for a sub-Gaussian process indexed by $\mathcal{S}$, there exists a constant $c>0$ such that $\mathbb{E}\left[\max_{a\in\mathcal{S}}\mu_{a}\right]\leq c\gamma_{2}(\mathcal{S})$. Using this functional, we establish a regret bound for our algorithm:
Theorem 12.
Let $\{\mu_{a}\}_{a\in\mathcal{A}_{\operatorname*{full}}}$ be a mean-zero sub-Gaussian process. Consider a partition :=$\mathcal{R}:=\{r_{\ell}\}_{\ell\leq m}$ of the full action space. Let $\mathcal{A}$ be output of Algorithm 1. Then, for a constant $C>0$,
$\begin{split}\mathbb{E}_{\theta,\mathcal{A}}[\textnormal{Regret}]\leq&C\sqrt{\log m}\cdot\max_{\ell\leq m}\gamma_{2}(r_{\ell})\\ &+\left(\mathbb{E}_{q}\left[(1-q(r))^{2K}\right]\cdot\mathbb{E}_{\theta}\left[\max_{a\in\mathcal{A}_{\operatorname*{full}}}\mu^{2}_{a}\right]\right)^{1/2}.\end{split}$
4 Generalization and Empirical Validation
In general cases, the decision-maker may not have explicit access to the full structure of the action space, especially in high-dimensional settings. Instead, they are given a list of actions and can observe the expected outcomes of these actions through sampling. We therefore adopt a dimension-free view by treating the expected outcomes $\{\mu_{a}\}_{a\in\mathcal{A}_{\operatorname*{full}}}$ as a family of random variables indexed by an abstract set $\mathcal{A}_{\operatorname*{full}}$. In a single bandit, these expected outcomes define an outcome function $f$ over the action space $\mathcal{A}_{\operatorname*{full}}$:
$f(a):=\mu_{a},\quad\forall a\in\mathcal{A}_{\operatorname*{full}}.$
[…]
$k(a,a^{\prime})=\sum\nolimits_{i\leq\infty}\lambda_{i}\phi_{i}(a)\phi_{i}(a^{\prime}),$
where $(\phi_{i})_{i\leq\infty}$ are orthonormal eigenfunctions under $\mathbb{P}$, and $(\lambda_{i})_{i\leq\infty}$ are non-negative eigenvalues. Let outcome $f$ be of the form $f(\cdot)=\sum_{i=1}^{N}\alpha_{i}k(\cdot,a_{i})$ for some integer $N\geq 1$, and a set of points $\{a_{i}\}_{i=1}^{N}\subset\mathcal{A}_{\operatorname*{full}}$ and a weight vector $\alpha\in\mathbb{R}^{N}$. The function $f$ can be rewritten as
$\mu_{a}=f(a)=\langle\mathbf{f},\Phi(a)\rangle,$
where $\mathbf{f}$ is a vector of coefficients, and $\Phi(a)$ usually refereed as a feature map, has entries $\Phi_{i}(a)=\sqrt{\lambda_{i}}\phi_{i}(a)$. If the coefficients $\{\mathbf{f}_{i}\}_{i=1}^{\infty}$ are i.i.d. Gaussian, this formulation can be represented as the canonical Gaussian process model in Equation (3). Note that in this case, the actual action space is the one formed by the feature vectors $\Phi(a)$ for $a\in\mathcal{A}_{\operatorname*{full}}$.
4.1 Varying-dependence Actions
Assuming the outcome function is generated from a kernel inherently implies that the expected outcomes of actions are correlated. To study the effect of varying correlation levels among actions, we conduct the following experiments:
Sampling Outcome Functions from a Kernel.
We use the stationary RBF kernel (Williams and Rasmussen, 2006), and provide additional illustrations using the non-stationary Gibbs kernel in Supplementary:
[…]
where $l$ is a length-scale parameter that control the dependence between actions. By (Kanagawa et al., 2018, Theorem 4.12), we sample functions from a RKHS by first constructing the kernel matrix $\mathbf{K}$ with entries $\mathbf{K}_{a,a^{\prime}}=k(a,a^{\prime})$ for $a,a^{\prime}\in\mathcal{A}_{\operatorname*{full}}$, and then drawing $f\sim\mathcal{N}(0,\mathbf{K})$.
Figure 1: Performance comparison on stochastic combinatorial optimization in Equation (6), selecting $K=5$ distinct actions from 15 grid points, with outcome functions $f(a):=\mu_{a}$ sampled from an RBF kernel at varying length-scales. Left: Expected regret for our method (green), Thompson Sampling (yellow), UCB (orange), and Successive Halving (SH, blue), averaged over 50 repetitions at each length-scale. TS/UCB are run for 3000 rounds; SH is given a budget of 37,000 pulls. Right: SH’s expected regret (blue) and pull counts (yellow bars) per round. Our method’s performance (green) is shown for comparison. SH requires nearly 10,000 pulls—about 3× the number of super arms—to match our method, making it infeasible for large $\mathcal{A}_{\operatorname*{full}}$.
Returning to the objective stated in the Introduction, we aim to find a subset $\mathcal{A}$ of cardinality $K$ that minimizes the expected regret defined in Equation (1). Since the term $\mathbb{E}[\max_{a\in\mathcal{A}_{\operatorname*{full}}}\mu_{a}]$ is a constant independent of $\mathcal{A}$, this is equivalent to the following optimization problem:
$\max_{\mathcal{A}\subseteq\mathcal{A}_{\operatorname*{full}}}\mathbb{E}\left[\max_{a\in\mathcal{A}}\mu_{a}\right]\quad\text{subject to }|\mathcal{A}|=K.$ (6)
In general, this problem is non-trivial because the objective function involves an expectation over a collection of random variables whose distribution is neither specified nor directly accessible. Moreover, the expression inside the expectation is the maximum over a subset of potentially correlated random variables. As a result, small changes in $\mathcal{A}$ can lead to non-smooth variations in the maximum.
Our method is computationally efficient, avoiding the combinatorial complexity of the optimization problem in Equation (6). To assess effectiveness in identifying near-optimal solutions, we compare our method against Thompson Sampling (TS) and Upper Confidence Bound (UCB) in a simple example, using Successive Halving (SH) as a reference for approximating the optimal subset—though SH becomes impractical when the full action space is large.
We consider a fixed action space consisting of 15 grid points in [0,2][0,2][ 0 , 2 ], using an RBF kernel in Equation (5) to sample outcome functions while varying the length-scale $l\in\{0.5,1,\dots,4\}$. In our method, we run Algorithm 1 until $K=5$ distinct actions are selected, where the payoff of each action in a round is given by $\mu_{a}$. We use exhaustive search over the action space to find the optimal action $\arg\max_{a\in\mathcal{A}_{\operatorname*{full}}}\mu_{a}$. For other methods, we treat each $K$-tuple of actions as a super arm, and the payoff of each super arm in a round $\mathcal{A}$ is given by $\max_{a\in\mathcal{A}}\mu_{a}$. Thus, these methods are tailored to find the best super arm that maximizes $\mathbb{E}\left[\max_{a\in\mathcal{A}}\mu_{a}\right]$, aligning with the same objective. In TS/UCB methods, we adopt a bandit feedback setting: at each round, the decision-maker selects a super arm $\mathcal{A}$, observes the payoff $\max_{a\in\mathcal{A}}\mu_{a}$, and updates its policy accordingly, repeating this process for 3000 rounds, chosen to roughly match the number of super arms ($N=3003$). Since the payoff $\max_{a\in\mathcal{A}}\mu_{a}$ is unbounded, we assume Gaussian payoffs for both TS and UCB. The algorithms are summarized in Supplementary, with the prior of TS set to $\mathcal{N}(0,1)$. In the SH method, all super arms are evaluated using a fixed budget of arm pulls (we use the minimum budget required in Karnin et al. (2013), $N\log_{2}N\approx 37{,}000$) over a few rounds. In each round, the number of remaining super arms is halved. This process continues until only the best-performing super arm remains or the budget is exhausted. We evaluate expected regret of all methods over $10^{5}$ randomly sampled outcome functions.
The left subplot of Figure 1 reports the expected regret of our method (green), TS (yellow), UCB (orange) and SH (blue), where solid curves and error shades indicate the mean $\pm$ one standard deviation of expected regret over 50 repetitions. As shown in this subplot, the expected regret decreases as the length-scale increases. This is because the length-scale $l$ of the RBF kernel controls the number of approximately independent actions in the process $\{\mu_{a}\}_{a\in\mathcal{A}_{\operatorname*{full}}}$: When actions $a$ and $a^{\prime}$ are far apart relative to the length-scale, the kernel value becomes very small, so the covariance between the function values $\mu_{a}$ and $\mu_{a^{\prime}}$ is close to zero.
Regarding computational complexity: in a given bandit instance $\theta$, identifying the optimal action $a^{*}(\theta):=\arg\max_{a\in\mathcal{A}_{\operatorname*{full}}}\mu_{a}(\theta)$ is a well-studied problem in both best-arm identification and GP optimization. However, if computing $a^{*}(\theta)$ is not computationally efficient, then identifying the optimal super arm, as in SH—whose search space grows combinatorially—is infeasible. Next, in the right subplot of Figure 1, we examine sample complexity. We run our method (green) and Successive Halving (SH, blue) for 50 additional repetitions, using a fixed length-scale $l=1$. Since SH proceeds in rounds, we track the current best super arm and the total number of super arm pulls so far after each round. The expected regret of the current best super arm, is evaluated over another $10^{5}$ randomly sampled outcome functions. The blue solid curves and shades for SH represent the mean $\pm$ one standard deviation of expected regret over 50 repetitions over each SH round. The yellow bars (corresponding to the right y-axis) show the number of super arm pulls by SH at the end of each round. For comparison, the green curve and shaded region represent the mean $\pm$ one standard deviation of expected regret for our method over the same 50 repetitions. As shown, SH’s performance roughly matches ours after three rounds, requiring 9005 super arm pulls—approximately three times the number of super arms. If the action space $|\mathcal{A}_{\operatorname*{full}}|$ is already large, running SH is intractable.
4.2 Conclusion and Future Work
We proposed a framework for selecting a subset of correlated actions, modeling payoff correlations with a Gaussian process and extending to the sub-Gaussian case. A simple algorithm was introduced and shown to effectively identify near-optimal subsets. A key direction for future work is developing a stopping criterion. When the subset size is flexible but sampling new bandit instances is costly, ideas from species discovery Roswell et al. (2021) may be useful—treating the discovery of a new action as analogous to discovering a new species. One could use Turing’s formula to estimate the probability of finding a new action, or frame the process as a one-armed bandit that stops when the discovery probability becomes sufficiently low.
Appendix A Connection to Bayesian Regret of Multi-Armed Bandits
Given a subset $\mathcal{A}\subseteq\mathcal{A}_{\operatorname*{full}}$, let $\mu^{*}(\mathcal{A},\theta):=\max_{a\in\mathcal{A}}\mu_{a}(\theta)$ be the highest expected outcome in a bandit instance $\theta$. We can equivalently rewrite the regret in Equation (1):
$\textnormal{Regret}(\theta)=\mu^{*}(\mathcal{A}_{\operatorname*{full}},\theta)-\mu^{*}(\mathcal{A},\theta).$
Bandit Regret: In the classic multi-armed bandit setting [Auer et al., 2002a, Lattimore and Szepesvári, 2020], outcome of an arm $a\in\mathcal{A}_{\operatorname*{full}}$ in a bandit instance $\theta$ is generated i.i.d. according to an unknown but fixed law with expectation $\mu_{a}(\theta)$. A decision-maker would follow a policy $\pi$ that chooses the next arm based on the sequence of rounds and obtained outcomes. The bandit regret of policy $\pi$ in this bandit environment $\theta$ over $N$ rounds is defined as the expected difference in cumulative outcomes over $N$ rounds between always selecting the optimal action and following a policy $\pi$ to choose actions. Bayesian bandit regret [Agrawal and Goyal, 2012, Lattimore and Szepesvári, 2020] is the expectation of this bandit regret, taken over the randomness in the distribution of bandit instances, and the action chosen by the policy:
$\begin{split}\textnormal{BanditRegret}(\theta):=&\mu^{*}(\mathcal{A}_{\operatorname*{full}},\theta)-\sum_{t=1}^{N}\mu_{A_{t}}(\theta),\\ \textnormal{BayesianBanditRegret}:=&\mathbb{E}\!\left[\mu^{*}(\mathcal{A}_{\operatorname*{full}},\theta)-\sum_{t=1}^{N}\mu_{A_{t}}(\theta)\right],\end{split}$
where $A_{t}$ is the action chosen by policy $\pi$ in round $t$. Since we consider settings where only a subset of the full action space is accessible, the action chosen by the policy can only from the subset $\mathcal{A}\subseteq\mathcal{A}_{\operatorname*{full}}$. This quantifies the expected difference in cumulative outcomes over $N$ rounds between always selecting the optimal action from the full action set $\mathcal{A}_{\operatorname*{full}}$ and following policy $\pi$ to choose actions from the restricted subset $\mathcal{A}$.
Next, we establish the connection between the regret defined in this paper and the Bayesian bandit regret.
A.1 Optimal policy
If $\pi$ is the optimal policy, it selects the best action in $\mathcal{A}$ in every round. Therefore, the regret of only having access to $\mathcal{A}$ instead of $\mathcal{A}_{\operatorname*{full}}$ for $N$ rounds, is determined by the optimal actions in $\mathcal{A}$ and $\mathcal{A}_{\operatorname*{full}}$:
[…]
A.2 Thompson sampling policy
If $\pi$ is the Thompson sampling policy, the Bayesian regret is bounded by
[…]
$=$ $\mathbb{E}\left[\sum_{t=1}^{N}\left(\mu^{*}(\mathcal{A})-\mu_{A_{t}}+\mu^{*}(\mathcal{A}_{\operatorname*{full}})-\mu^{*}(\mathcal{A})\right)\right]$
$\leq$ $C\sqrt{|\mathcal{A}|\cdot N\log N}+N\cdot\mathbb{E}_{\theta}[\textnormal{Regret}],$
where $|\mathcal{A}|$ denotes the cardinality, and $C>0$ is a constant. The inequality uses the Bayesian (bandit) regret of Thompson sampling (Theorem 36.1 in Lattimore and Szepesvári [2020]). Note that with access to the full action space, the Bayesian regret is bounded by $C\sqrt{|\mathcal{A}_{\operatorname*{full}}|\cdot N\log N}$. However, this can be worse than restricting to a subset $\mathcal{A}$ when $\mathbb{E}_{\theta}[\textnormal{Regret}]$ is sufficiently small and the time horizon $N$ is finite. Intuitively speaking, using a well-chosen subset of representative actions instead of the full action space can reduce Bayesian regret in multi-armed bandits.
Appendix B Proof of Lemma 4
Statement: Given a partition $\mathcal{R}$ of the full action space, and the importance measure $q$ assigning a value to each cluster $r\in\mathcal{R}$. Let $\mathcal{A}$ be the output of Algorithm 1 after $K$ samples. Then, with probability at least $1-\frac{1}{\epsilon}\exp(-K\epsilon)$, it holds that for any cluster $r\in\mathcal{R}$,
$r\cap\mathcal{A}\neq\emptyset\quad\text{whenever}\quad q(r)>\epsilon.$
Proof.
By definition, the algorithm draws $K$ i.i.d. samples from clusters in $\mathcal{R}$ according to the distribution $q$. Define a set of typical clusters $R_{\epsilon}:=\{r\in\mathcal{R}:q(r)>\epsilon\}$. Then,
$\Pr[r\cap\mathcal{A}\!=\!\emptyset]=(1-q(r))^{K}<(1-\epsilon)^{K}\leq\exp(-K\epsilon),\;\forall r\in R_{\epsilon},$ (7)
where the equality is the probability of missing the cluster $r$ for $K$ times, the first inequality uses the definition of $R_{\epsilon}$. The last inequality uses $\log(1-\epsilon)^{K}=K\log(1-\epsilon)$ and the inequality $\log(1-\epsilon)\leq-\epsilon$ for $\epsilon<1$ [TOPSØE1, 2007]. Therefore,
$\Pr[r\cap\mathcal{A}\neq\emptyset\;\text{whenever}\;q(r)>\epsilon]$
$=$ $1-\Pr[r\cap\mathcal{A}\neq\emptyset\;\exists r\in R_{\epsilon}]$
$\geq$ $1-\sum_{r\in R_{\epsilon}}\Pr[r\cap\mathcal{A}\neq\emptyset]$
$>$ $1-\sum_{r\in R_{\epsilon}}\exp(-K\epsilon)$
$=$ $1-|R_{\epsilon}|\exp(-K\epsilon)\geq 1-1/\epsilon\cdot\exp(-K\epsilon),$
where the first inequality uses union bound, the second inequality uses Equation (7). The last inequality uses the fact that there could be at most $\lfloor 1/\epsilon\rfloor$ clusters in $R_{\epsilon}$. ∎
Appendix C Proof of Theorem 7
Lemma 13 (Regret decomposition of reference subsets).
Consider a partition :=$\mathcal{R}:=\{r_{\ell}\}_{\ell\leq m}$ of the full action space and an arbitrary reference subset $\mathcal{A}$. For each $\ell\leq m$, define a Gaussian process $\left\{Z_{a}\right\}_{a\in r_{\ell}}$ where $Z_{a}:=\mu_{a}-\mu_{a_{\ell}}$. Define a non-negative random variable
$Y_{\ell}:=\sup_{a\in r_{\ell}}Z_{a}=\sup_{a\in r_{\ell}}\mu_{a}-\mu_{a_{\ell}}.$
Under Assumption 6, $\operatorname*{Regret}=Y_{\ell}$ where the cluster $r_{\ell}$ contains the optimal action. Hence,
$\mathbb{E}_{\theta}\left[\operatorname*{Regret}\right]=\sum_{\ell\leq m}q(r_{\ell})\cdot\mathbb{E}_{\theta}\left[Y_{\ell}\Big{|}a^{*}(\theta)\in r_{\ell}\right].$
Proof.
Let the cluster $r_{\ell}$ contains the optimal action, i.e., $a^{*}(\theta)\in r_{\ell}$, then
$\operatorname*{Regret}=\max_{a\in\mathcal{A}_{\operatorname*{full}}}\mu_{a}-\max_{a^{\prime}\in\mathcal{A}}\mu_{a^{\prime}}=\max_{a\in r_{\ell}}\mu_{a}-\mu_{a_{\ell}}=Y_{\ell},$
where the middle equality uses that when $a^{*}(\theta)\in r_{\ell}$, $\max_{a\in\mathcal{A}_{\operatorname*{full}}}\mu_{a}=\max_{a\in r_{\ell}}\mu_{a}$ and further with Assumption 6, $\max_{a^{\prime}\in\mathcal{A}}\mu_{a^{\prime}}\geq\mu_{a_{\ell}}$ when $a^{*}(\theta)\in r_{\ell}$. Then,
$\mathbb{E}\left[\operatorname*{Regret}\right]=\sum_{\ell\leq m}\Pr[a^{*}(\theta)\in r_{\ell}]\cdot\mathbb{E}\left[Y_{\ell}\Big{|}a^{*}(\theta)\in r_{\ell}\right]=\sum_{\ell\leq m}q(r_{\ell})\cdot\mathbb{E}\left[Y_{\ell}\Big{|}a^{*}(\theta)\in r_{\ell}\right],$
where the left equality uses tower rule, the right equality uses the definition of $q$ measure. ∎
To bound the expected regret of reference subsets, we need a few more lemmas:
Lemma 14 (Expectation integral identity).
Given a non-negative random variables $X$. If $\Pr[X\geq u]\leq c\exp{\left(-\frac{u^{2}}{\epsilon^{2}}\right)}$ for any $u>0$, then $\mathbb{E}X\leq C\epsilon\sqrt{\log c}$, where $\epsilon,c,C$ are positive constants.
Proof.
[…]
$=$ $u_{0}+\exp\left(-\frac{u_{0}^{2}}{\epsilon^{2}}\right)\cdot\frac{c\epsilon^{2}}{2u_{0}}$
$=$ $\epsilon\sqrt{\log c}+\frac{\epsilon}{2\sqrt{\log c}}\leq C\epsilon\sqrt{\log c},$
where the first equality uses integrated tail formula of expectation (cf. Lemma 1.6.1 of Vershynin [2018]). The last equality set $u_{0}:=\epsilon\sqrt{\log c}$. ∎
Lemma 15 (Borell-TIS inequality; Lemma 2.4.7 of [Talagrand, 2014]).
Given a set $\mathcal{S}$, and a zero-mean Gaussian process $\left(X_{a}\right)_{a\in\mathcal{S}}$. Let $\epsilon:=\sup_{a\in\mathcal{S}}\left(\mathbb{E}X_{a}^{2}\right)^{\frac{1}{2}}$. Then for $u>0$, we have
$\Pr\left[\left|\sup_{a\in\mathcal{S}}X_{a}-\mathbb{E}\sup_{a\in\mathcal{S}}X_{a}\right|\geq u\right]\leq 2\exp{\left(-\frac{u^{2}}{2\epsilon^{2}}\right)}.$
It means that the size of the fluctuations of $\mathbb{E}\sup_{a\in\mathcal{S}}X_{a}$ is governed by the size of the individual random variables $X_{a}$.
Statement of Theorem 7:
Consider a partition :=$\mathcal{R}:=\{r_{\ell}\}_{\ell\leq m}$ of the full action space, with $\epsilon:=\max_{r\in\mathcal{R}}\operatorname{diam}(r)$, and an arbitrary reference subset $\mathcal{A}$. Then, for some constant $C>0$,
$\mathbb{E}_{\theta}[\operatorname*{Regret}]\leq\max_{r\in\mathcal{R}}\mathbb{E}_{\theta}\left[\max_{a\in r}\mu_{a}\right]+C\epsilon\sqrt{\log|\mathcal{R}|}.$
If the partition $\mathcal{R}$ satisfies Assumption 6, then
$\mathbb{E}_{\theta}[\operatorname*{Regret}]\geq\min_{r\in\mathcal{R}}\mathbb{E}_{\theta}\left[\max_{a\in r}\mu_{a}\right]-C\epsilon\sqrt{\log|\mathcal{R}|}.$
Proof.
Fix $\ell$, define a Gaussian process $\left\{Z_{a}\right\}_{a\in r_{\ell}}$ where $Z_{a}:=\mu_{a}-\mu_{a_{\ell}}$. Define a non-negative random variable
$Y_{\ell}:=\sup_{a\in r_{\ell}}Z_{a}=\sup_{a\in r_{\ell}}\mu_{a}-\mu_{a_{\ell}}.$
Since $\mathbb{E}\mu_{a}=0$ for all $a\in\mathcal{A}_{\operatorname*{full}}$, it holds that $\mathbb{E}Y_{\ell}=\mathbb{E}\max_{a\in r_{\ell}}\mu_{a}$.
When the cluster $r_{\ell}$ contains the optimal action, i.e., $a^{*}(\theta)\in r_{\ell}$, we have
$\operatorname*{Regret}=\max_{a\in\mathcal{A}_{\operatorname*{full}}}\mu_{a}-\max_{a^{\prime}\in\mathcal{A}}\mu_{a^{\prime}}=\max_{a\in r_{\ell}}\mu_{a}-\max_{a^{\prime}\in\mathcal{A}}\mu_{a^{\prime}}\leq\max_{a\in r_{\ell}}\mu_{a}-\mu_{a_{\ell}}=Y_{\ell}\leq\max_{\ell^{\prime}\leq m}Y_{\ell^{\prime}},$
where the first equality follows from the definition of regret in Equation (1), and the second equality follows from the assumption that $a^{*}(\theta)\in r_{\ell}$. The inequality holds because $a_{\ell}\in\mathcal{A}$, and hence $\max_{a\in\mathcal{A}}\mu_{a}\geq\mu_{a_{\ell}}$. The final equality follows from the definition of $Y_{\ell}$. On the other hand, according to Lemma 13, under Assumption 6, $\operatorname*{Regret}=Y_{\ell}$, and thus $\operatorname*{Regret}\geq\min_{\ell^{\prime}\leq m}Y_{\ell^{\prime}}$, where the cluster $r_{\ell}$ contains the optimal action.
Therefore, we reach the important conclusion that
$\min_{\ell\leq m}Y_{\ell}\leq\operatorname*{Regret}\leq\max_{\ell\leq m}Y_{\ell},$ (8)
where the left inequality holds under Assumption 6. Further, by definition $r_{\ell}\subset B(a_{\ell},\epsilon)$, such that $\mathbb{E}Z_{a}^{2}=\mathbb{E}(\mu_{a}-\mu_{a_{\ell}})^{2}=\|a-a_{\ell}\|^{2}_{2}\leq\epsilon^{2}$. Using Lemma 15 on the process $\left\{Z_{a}\right\}_{a\in r_{\ell}}$, we have
$\Pr\left[\left|Y_{\ell}-\mathbb{E}Y_{\ell}\right|\geq u\right]\leq 2\exp{\left(-\frac{u^{2}}{2\epsilon^{2}}\right)}.$
By union bound, we have
$\Pr\left[\max_{\ell\leq m}\left|Y_{\ell}-\mathbb{E}Y_{\ell}\right|\geq u\right]\leq 2m\exp{\left(-\frac{u^{2}}{2\epsilon^{2}}\right)}.$
Using Lemma 14, we have for some absolute constant $C>0$
$\mathbb{E}\max_{\ell\leq m}\left|Y_{\ell}-\mathbb{E}Y_{\ell}\right|\leq C\epsilon\sqrt{\log m}.$ (9)
Upper bound:
$\mathbb{E}\operatorname*{Regret}\leq\mathbb{E}\max_{\ell\leq m}Y_{\ell}\leq\max_{\ell\leq m}\mathbb{E}Y_{\ell}+\mathbb{E}\max_{\ell\leq m}\left|Y_{\ell}-\mathbb{E}Y_{\ell}\right|\leq\max_{\ell\leq m}\mathbb{E}\max_{a\in r_{\ell}}\mu_{a}+C\epsilon\sqrt{\log m},$ (10)
where the first inequality uses Equation (8). The second inequality uses $\max_{\ell\leq m}Y_{\ell}\leq\max_{\ell\leq m}\mathbb{E}Y_{\ell}+\max_{\ell\leq m}\left|Y_{\ell}-\mathbb{E}Y_{\ell}\right|$, since $Y_{\ell}\leq\mathbb{E}Y_{\ell}+\left|Y_{\ell}-\mathbb{E}Y_{\ell}\right|$. The last inequality uses Equation (9) and the identity $\mathbb{E}[Y_{\ell}]=\mathbb{E}[\max_{a\in r_{\ell}}\mu_{a}]$.
Lower bound:
$\mathbb{E}\operatorname*{Regret}\geq\mathbb{E}\min_{\ell\leq m}Y_{\ell}\geq\min_{\ell\leq m}\mathbb{E}Y_{\ell}-\mathbb{E}\max_{\ell\leq m}\left|Y_{\ell}-\mathbb{E}Y_{\ell}\right|\geq\min_{\ell\leq m}\mathbb{E}\max_{a\in r_{\ell}}\mu_{a}-C\epsilon\sqrt{\log m},$ (11)
where the first inequality uses Equation (8) under Assumption 6. The second inequality uses $\min_{\ell\leq m}Y_{\ell}\geq\min_{\ell\leq m}\mathbb{E}Y_{\ell}-\max_{\ell\leq m}\left|Y_{\ell}-\mathbb{E}Y_{\ell}\right|$, since $Y_{\ell}\geq\mathbb{E}Y_{\ell}-\left|Y_{\ell}-\mathbb{E}Y_{\ell}\right|$. The last inequality uses Equation (9) and the identity $\mathbb{E}[Y_{\ell}]=\mathbb{E}[\max_{a\in r_{\ell}}\mu_{a}]$. ∎
Appendix D Proof of Theorem 8
Lemma 16 (Transformation invariance).
Given any vector $x\in\mathbb{R}^{n}$, let $S+x:=\{s+x:s\in\mathcal{S}\}$. If $\mathbb{E}[\theta]=0$, then $\mathbb{E}\left[\max_{a\in\mathcal{S}+x}\langle a,\theta\rangle\right]=\mathbb{E}\left[\max_{a^{\prime}\in\mathcal{S}}\langle a^{\prime},\theta\rangle\right]$.
Proof.
$\mathbb{E}\left[\max_{a\in\mathcal{S}+x}\langle a,\theta\rangle\right]=$ $\mathbb{E}\left[\max_{a^{\prime}\in\mathcal{S}}\langle a^{\prime}+x,\theta\rangle\right]$
$=$ $\mathbb{E}\left[\max_{a^{\prime}\in\mathcal{S}}\langle a^{\prime},\theta\rangle\right]+\mathbb{E}\left[\langle x,\theta\rangle\right]$
$=$ $\mathbb{E}\left[\max_{a^{\prime}\in\mathcal{S}}\langle a^{\prime},\theta\rangle\right],$
where the last equality uses $\mathbb{E}\left[\langle x,\theta\rangle\right]=\left\langle x,\mathbb{E}[\theta]\right\rangle=0$. ∎
Lemma 17.
Consider a partition $\mathcal{R}$ of the full action space. Let $\mathcal{A}$ be the output of Algorithm 1. Then, the event that the optimal action falls in a cluster $r$, i.e., $\{a^{*}(\theta)\in r\}$, is independent of whether the subset $\mathcal{A}$ intersects with the cluster. It holds that
$\Pr[a^{*}(\theta)\in r,\,r\cap\mathcal{A}\!=\!\emptyset]=$ $q(r)(1-q(r))^{K},$
$\Pr[a^{*}(\theta)\in r,r\cap\mathcal{A}\!\neq\!\emptyset]\leq$ $q(r).$
Proof.
Since $\mathcal{A}$ is the output of Algorithm 1, the event $\{a^{*}(\theta)\in r\}$ is independent from $\{r\cap\mathcal{A}\!\neq\!\emptyset\}$ or $\{r\cap\mathcal{A}\!=\!\emptyset\}$. We have
$\Pr[a^{*}(\theta)\in r,r\cap\mathcal{A}\!=\!\emptyset]=$ $\Pr[a^{*}(\theta)\in r]\Pr[r\cap\mathcal{A}\!=\!\emptyset]$
$=$ $q(r)\Pr[r\cap\mathcal{A}\!=\!\emptyset]$
$=$ $q(r)(1-q(r))^{K},$
where the first equality uses independence between $\{a^{*}(\theta)\in r\}$ and $\{r\cap\mathcal{A}\!\neq\!\emptyset\}$, the second equality uses the definition of measure $q$, the third equality use the probability of missing cluster $r$ in $K$ i.i.d. samplings. Similarly,
$\Pr[a^{*}(\theta)\in r,r\cap\mathcal{A}\!\neq\!\emptyset]=q(r)\cdot\Pr[r\cap\mathcal{A}\!\neq\!\emptyset].$
Using $\Pr[r\cap\mathcal{A}\!\neq\!\emptyset]\leq 1$, we complete the proof. ∎
Statement of Theorem 8:
Let $\mathcal{A}$ be output of Algorithm 1. Consider a partition :=$\mathcal{R}:=\{r_{\ell}\}_{\ell\leq m}$ of the full action space, with $\epsilon:=\max_{r\in\mathcal{R}}\operatorname{diam}(r)$. Then, for the same constant $C>0$ in Theorem 7,
$\begin{split}\mathbb{E}_{\theta,\mathcal{A}}[\textnormal{Regret}]\leq&\max_{r\in\mathcal{R}}\mathbb{E}_{\theta}\left[\max_{a\in r}\mu_{a}\right]+C\epsilon\sqrt{\log|\mathcal{R}|}\\ &+\left(\mathbb{E}_{q}\left[(1-q(r))^{2K}\right]\cdot\mathbb{E}_{\theta}\left[\max_{a\in\mathcal{A}_{\operatorname*{full}}}\mu^{2}_{a}\right]\right)^{1/2}.\end{split}$
Proof.
If $r_{\ell}\cap\mathcal{A}\neq\emptyset$, define $a_{\ell}\in r_{\ell}\cap\mathcal{A}$. If $r_{\ell}\cap\mathcal{A}\!=\!\emptyset$, choose an arbitrary point $a_{\ell}\in r_{\ell}$ as the representative. The set $\mathcal{A}^{\prime}:=\{a_{\ell}\}_{\ell\leq m}$ forms a reference subset of Definition 1. The cluster $r_{\ell}$ is contained in a closed Euclidean ball of radius $\epsilon$ centered at $a_{\ell}$, i.e., $r_{\ell}\subset B(a_{\ell},\epsilon)$. Define a Gaussian process $\left\{Z_{a}\right\}_{a\in r_{\ell}}$ where $Z_{a}:=\mu_{a}-\mu_{a_{\ell}}$. Define a random variable
$Y_{\ell}:=\sup_{a\in r_{\ell}}Z_{a}=\sup_{a\in r_{\ell}}\mu_{a}-\mu_{a_{\ell}}.$
Since $\mathbb{E}\mu_{a}=0$ for all $a\in\mathcal{A}_{\operatorname*{full}}$, we have $\mathbb{E}Y_{\ell}=\mathbb{E}\sup_{a\in r_{\ell}}\mu_{a}$.
Consider the case that $a^{*}(\theta)\in r_{\ell}$. We have
$\begin{split}&\mathbb{E}\left[\operatorname*{Regret}\Big{|}r_{\ell}\cap\mathcal{A}\!\neq\!\emptyset,a^{*}(\theta)\in r_{\ell}\right]\\ \leq&\mathbb{E}\left[Y_{\ell}\Big{|}r_{\ell}\cap\mathcal{A}\!\neq\!\emptyset,a^{*}(\theta)\in r_{\ell}\right]\\ =&\mathbb{E}\left[Y_{\ell}\Big{|}a^{*}(\theta)\in r_{\ell}\right],\\ \leq&\mathbb{E}\left[\max_{\ell\leq m}Y_{\ell}\Big{|}a^{*}(\theta)\in r_{\ell}\right],\end{split}$ (12)
where the first inequality uses $\operatorname*{Regret}\leq Y_{\ell}$ if $r_{\ell}\cap\mathcal{A}\!\neq\!\emptyset$, the equality uses that $Y_{\ell}$ is independent of $r_{\ell}\cap\mathcal{A}\!\neq\!\emptyset$. On the other hand, we assume that $0\in\operatorname*{Conv}(\mathcal{A})$ because even if it doesn’t hold, we can always find a vector $x\in\mathbb{R}^{n}$ such that $0\in\operatorname*{Conv}(\mathcal{A}+x)$, without changing the value of $\mathbb{E}[\max_{a\in\mathcal{A}}\mu_{a}]$ (c.f. Lemma 16). Therefore,
$\begin{split}&\mathbb{E}\left[\operatorname*{Regret}\Big{|}r_{\ell}\cap\mathcal{A}\!=\!\emptyset,a^{*}(\theta)\in r_{\ell}\right]\\ \leq&\mathbb{E}\left[\max_{a\in\mathcal{A}_{\operatorname*{full}}}\mu_{a}\Big{|}a^{*}(\theta)\in r_{\ell}\right],\end{split}$ (13)
where the inequality uses $\operatorname*{Regret}\leq\max_{a\in\mathcal{A}_{\operatorname*{full}}}\mu_{a}$, as a consequence of $0\in\operatorname*{Conv}(\mathcal{A})$, and that $\max_{a\in\mathcal{A}_{\operatorname*{full}}}\mu_{a}$ is independent of $r_{\ell}\cap\mathcal{A}\!=\!\emptyset$. Further,
$\mathbb{E}[\operatorname*{Regret}]=$ $\sum_{r\in\mathcal{R}}\Pr[r\cap\mathcal{A}\neq\emptyset,a^{*}(\theta)\in r]\cdot\mathbb{E}\left[\operatorname*{Regret}\Big{|}r\cap\mathcal{A}\neq\emptyset,a^{*}(\theta)\in r\right]$ (14)
$+\sum_{r\in\mathcal{R}}\Pr[r\cap\mathcal{A}=\emptyset,a^{*}(\theta)\in r]\cdot\mathbb{E}\left[\operatorname*{Regret}\Big{|}r\cap\mathcal{A}=\emptyset,a^{*}(\theta)\in r\right]$
$\leq$ $\sum_{\ell\leq m}q(r_{\ell})\cdot\left(\mathbb{E}\left[\max_{\ell\leq m}Y_{\ell}\Big{|}a^{*}(\theta)\in r_{\ell}\right]+(1-q(r_{\ell}))^{K}\cdot\mathbb{E}\left[\max_{a\in\mathcal{A}_{\operatorname*{full}}}\mu_{a}\Big{|}a^{*}(\theta)\in r_{\ell}\right]\right)$
$=$ $\mathbb{E}\left[\max_{\ell\leq m}Y_{\ell}\right]+\sum_{r\in\mathcal{R}}q(r)\cdot(1-q(r))^{K}\cdot\mathbb{E}\left[\max_{a\in\mathcal{A}_{\operatorname*{full}}}\mu_{a}\Big{|}a^{*}(\theta)\in r\right]$
$\leq$ $\mathbb{E}\left[\max_{\ell\leq m}Y_{\ell}\right]+\left(\mathbb{E}_{q}\left[(1-q(r))^{2K}\right]\cdot\mathbb{E}_{\theta}\left[\max_{a\in\mathcal{A}_{\operatorname*{full}}}\mu^{2}_{a}\right]\right)^{1/2},$
where the first equality uses tower rule. The first inequality uses Equations (12-13), and Lemma 17. The last equality uses tower rule again. The last inequality uses Cauchy-Schwarz inequality, which states that for any two random variables $X$ and $Y$, we have $|\mathbb{E}[XY]|\leq\sqrt{\mathbb{E}[X^{2}]\mathbb{E}[Y^{2}]}$.
Note that the Gaussian process assumption is not used in Equation (14); it is only needed to bound the term $\mathbb{E}\left[\max_{\ell\leq m}Y_{\ell}\right]$. By applying Equation (10) in the proof of Theorem 7, we bound this term and complete the proof. ∎
Appendix E Proof of Theorem 10
Lemma 18.
Let $q$ denote a discrete probability distribution over a finite support $\mathcal{R}$. Define
$\begin{split}M:=\max_{q}&\;\mathbb{E}_{q}\left[(1-q(r))^{K}\right]\\ \textnormal{s.t.}&\sum_{r\in\mathcal{R}}q(r)=1,\;q(r)\in[0,1)\;\;\forall r\in\mathcal{R}.\end{split}$ (15)
When $|\mathcal{R}|\geq K+1$, the maximum is $M=\left(1-\frac{1}{|\mathcal{R}|}\right)^{K}$, and it is attained when $q$ is uniform. When $|\mathcal{R}|<K+1$, the maximum is upper bounded by $M\leq\frac{|\mathcal{R}|}{K+1}\left(\frac{K}{K+1}\right)^{K}$, and there exists a feasible distribution $q^{\prime}$ such that $\mathbb{E}_{q^{\prime}}\left[(1-q^{\prime}(r))^{K}\right]\geq\frac{|\mathcal{R}|-1}{K+1}\left(\frac{K}{K+1}\right)^{K}$.
Proof.
Let =$\mathcal{R}=\{r_{1},\dots,r_{m}\}$ be the finite support of the measure $q$, where $m:=|\mathcal{R}|$. Let $q_{i}:=q(r_{i})$ denote the probability mass at each support point. Define the function $f(q_{\ell}):=q_{\ell}\cdot(1-q_{\ell})^{K}$.
Case I $m<K+1$: The first derivative of $f$ is
$f^{\prime}(q_{\ell})=(1-q_{\ell})^{K-1}\big{(}1-(K+1)q_{\ell}\big{)},$
which is positive on the interval $\left[0,\frac{1}{K+1}\right)$ and negative on the interval $\left(\frac{1}{K+1},1\right)$. Therefore, $f(q_{\ell})$ attains its maximum over [0,1)[0,1)[ 0 , 1 ) at
$q^{*}_{\ell}=\frac{1}{K+1},$
with the corresponding maximum value
$f(q^{*}_{\ell})\leq\frac{1}{K+1}\left(\frac{K}{K+1}\right)^{K}.$
Since there are $m$ support points, this yields the upper bound $M\leq\sum_{\ell=1}^{m}f(q^{*}_{\ell})\leq\frac{m}{K+1}\left(\frac{K}{K+1}\right)^{K}$. Also, since $\frac{1}{K+1}<\frac{1}{m}$, the solution $q_{1}=\dots=q_{m-1}=\frac{1}{K+1}$ and $q_{m}=\frac{K-m+2}{K+1}$ is feasible. Thus,
$M\geq$ $\sum_{\ell=1}^{m-1}f\left(\frac{1}{K+1}\right)+f\left(\frac{K-m+2}{K+1}\right)$
$\geq$ $\frac{m-1}{K+1}\left(\frac{K}{K+1}\right)^{K},$
where the right inequality uses that $f(q_{\ell})\geq 0$ for $q_{\ell}\in[0,1]$.
Case II $m\geq K+1$: Consider the relaxed maximization problem of Equation (15):
$\max_{q_{1},\dots,q_{m}\in[0,1)}\sum_{\ell=1}^{m}f(q_{\ell})\quad\text{subject to }\sum_{\ell=1}^{m}q_{\ell}\leq 1.$ (16)
Let $\lambda\geq 0$ be the Lagrange multiplier associated with the constraint. Define the Lagrangian:
$\mathcal{L}(q_{1},\dots,q_{m},\lambda)=\sum_{\ell=1}^{m}q_{\ell}(1-q_{\ell})^{K}-\lambda\left(\sum_{\ell=1}^{m}q_{\ell}-1\right).$
For each $\ell=1,\dots,m$, compute the partial derivative of $\mathcal{L}$ with respect to $q_{\ell}$:
$\frac{\partial}{\partial q_{\ell}}\left[q_{\ell}(1-q_{\ell})^{K}\right]=(1-q_{\ell})^{K}-Kq_{\ell}(1-q_{\ell})^{K-1}.$
Setting this derivative equal to zero yields the stationary condition:
$(1-q_{\ell})^{K-1}(1-(K+1)q_{\ell})=\lambda,\quad\text{with }\lambda\geq 0.$
To find critical points of Equation (16), we solve the system:
$(1-q_{\ell})^{K-1}(1-(K+1)q_{\ell})=\lambda\geq 0\quad\forall\ell\leq m,\quad\text{and}\quad\sum_{\ell=1}^{m}q_{\ell}\leq 1.$
Define the function $g(q_{\ell}):=(1-q_{\ell})^{K-1}(1-(K+1)q_{\ell})$. For $g(q_{\ell})\geq 0$, it must hold that $1-(K+1)q_{\ell}\geq 0$, i.e., $q_{\ell}\leq\frac{1}{K+1}$. Therefore, any feasible solution to this system must satisfy $q_{\ell}\in\left[0,\frac{1}{K+1}\right]$ for all $\ell\leq m$.
Then, over the interval $q_{\ell}\in\left[0,\frac{1}{K+1}\right]$, both factors $(1-q_{\ell})^{K-1}$ and $1-(K+1)q_{\ell}$ are positive and decreasing. Hence, $g(q_{\ell})$ is positive and strictly decreasing, so the equation $g(q_{\ell})=\lambda\geq 0$ has at most one solution. Therefore, all $q_{\ell}$’s must be equal at a critical point. Let $q_{\ell}=c$ for all $\ell\leq m$. Due to the assumption of $m\geq K+1$, we have $\frac{1}{m}\leq\frac{1}{K+1}$, so any choice of $c\in\left[0,\frac{1}{m}\right]$ satisfies the constraint $\sum_{\ell=1}^{m}q_{\ell}\leq 1$ and is feasible for the system.
Therefore, $q_{\ell}=c$ for all $\ell\leq m$ is a feasible critical point of Equation (16). The corresponding objective value is:
$\sum_{\ell=1}^{m}f(c)=m\cdot c\cdot(1-c)^{K},$
which is increasing in $c$ over the interval $\left[0,\frac{1}{m}\right]$. Hence, the maximum is attained at $c=\frac{1}{m}$, and the optimal value is:
$\sum_{\ell=1}^{m}f\left(\frac{1}{m}\right)=\left(1-\frac{1}{m}\right)^{K}=\left(1-\frac{1}{|\mathcal{R}|}\right)^{K},$
achieved when $q_{\ell}=\frac{1}{m}$ for all $\ell$.
To confirm that this critical point is indeed a maximum, observe that the Hessian of the objective function $f(q_{\ell})$ is diagonal (since the function is separable), and the diagonal entries are:
$\frac{\partial^{2}}{\partial q_{\ell}^{2}}\left[q_{\ell}(1-q_{\ell})^{K}\right]=-(1-q_{\ell})^{K-2}\left(2K-K(K+1)q_{\ell}\right),$
which is negative for $q_{\ell}\leq\frac{1}{K+1}$, because then $2-(K+1)q_{\ell}>0$. Hence, the Hessian is negative definite, and the critical point is a local (and thus global) maximum.
Finally, note that Equation (16) is a relaxation of Equation (15). While the optimum of the original problem is upper bounded by that of the relaxed problem, the optimal solution to the relaxed problem also lies within the feasible region of the original problem. Therefore, the maximum of Equation (15) is also attained when $q$ is uniform, with the maximum value being $\left(1-\frac{1}{|\mathcal{R}|}\right)^{K}$. ∎
Statement of Theorem 10
Under Assumption 3, there exists a point $a_{0}$ and a constant $M>0$ such that $\mathcal{A}_{\operatorname*{full}}\subset B(a_{0},M)$, a closed Euclidean ball of radius $M$ centered at $a_{0}$. Let the action space have dimension $n\in\mathbb{N}$, and fix a constant $0<\epsilon<M$. Let $\mathcal{A}$ be the output of Algorithm 1. For the same constant $C>0$ in Theorem 7, we have:
$\mathbb{E}_{\theta,\mathcal{A}}[\operatorname*{Regret}]\leq 2\epsilon\sqrt{n}+C\epsilon\sqrt{\log N(\mathcal{A}_{\operatorname*{full}},\epsilon)},\quad\text{where }K\geq\frac{1}{2}\left(\frac{M^{2}N(\mathcal{A}_{\operatorname*{full}},\epsilon)}{\epsilon^{2}e}-1\right).$
As $\epsilon\to 0^{+}$, we have $\mathbb{E}_{\theta,\mathcal{A}}[\operatorname*{Regret}]\to 0$ as $K\to\infty$.
Proof.
By Lemma 16, we can shift the action space to be centered at the origin. So, without loss of generality, we assume that $\mathcal{A}_{\operatorname*{full}}\subset M\cdot B_{2}^{n}$, the scaled unit Euclidean ball in $\mathbb{R}^{n}$. Then,
$\mathbb{E}_{\theta}\left[\max_{a\in\mathcal{A}_{\operatorname*{full}}}\mu_{a}^{2}\right]\leq M^{2}\cdot\mathbb{E}\|\theta\|_{2}^{2}=M^{2}n,$ (17)
where the inequality follows from $\mu_{a}=\langle\theta,a\rangle\leq M\|\theta\|_{2}$ for all $a\in MB_{2}^{n}$, and the equality uses that $\theta\sim\mathcal{N}(0,I)$, so each coordinate has variance 1 and $\|\theta\|_{2}^{2}=\sum_{i=1}^{n}\theta_{i}^{2}$ has expectation $n$.
Let $\{a_{1},\dots,a_{m}\}\subseteq\mathcal{A}_{\operatorname*{full}}$ be a minimal geometric $\epsilon$-net under the Euclidean norm, so that $m=N(\mathcal{A}_{\operatorname*{full}},\epsilon)$. Define $\pi(a)$ as the closest point in the $\epsilon$-net to $a$, and let the partition =$\mathcal{R}=\{r_{1},\dots,r_{m}\}$ be given by
$r_{\ell}=\left\{a\in\mathcal{A}_{\operatorname*{full}}:\pi(a)=a_{\ell}\right\}.$
Then,
$\max_{\ell\leq m}\mathbb{E}_{\theta}\left[\max_{a\in r_{\ell}}\mu_{a}\right]\leq\mathbb{E}_{\theta}\left[\max_{a\in B(a_{\ell},\epsilon)}\mu_{a}\right]=\mathbb{E}_{\theta}\left[\max_{a\in\epsilon B_{2}^{n}}\mu_{a}\right]\leq\epsilon\sqrt{n},$ (18)
where the first inequality follows from $r_{\ell}\subset B(a_{\ell},\epsilon)$ by construction; the equality uses Lemma 16 to shift; and the final bound uses Claim 3 from Supplementary K. Then,
$\mathbb{E}_{\theta,\mathcal{A}}[\textnormal{Regret}]\leq$ $\max_{r\in\mathcal{R}}\mathbb{E}_{\theta}\left[\max_{a\in r}\mu_{a}\right]+C\epsilon\sqrt{\log m}+\left(\mathbb{E}_{q}\left[(1-q(r))^{2K}\right]\cdot\mathbb{E}_{\theta}\left[\max_{a\in\mathcal{A}_{\operatorname*{full}}}\mu^{2}_{a}\right]\right)^{1/2},$
$\leq$ $\epsilon\sqrt{n}+C\epsilon\sqrt{\log m}+\sqrt{\frac{m}{2K+1}}\left(\frac{2K}{2K+1}\right)^{K}\cdot M\sqrt{n}$
$\leq$ $\epsilon\sqrt{n}+C\epsilon\sqrt{\log m}+\epsilon\sqrt{n},$
where the first inequality follows from the algorithm-dependent upper bound in Theorem 8. The second inequality follows from Equations (17) and (18), and the term $\mathbb{E}_{q}\left[(1-q(r))^{2K}\right]$ is bounded by $\frac{m}{2K+1}\left(\frac{2K}{2K+1}\right)^{2K}$ (see Lemma 18). The last inequality comes from
$\log\left(\sqrt{\frac{m}{2K+1}}\left(\frac{2K}{2K+1}\right)^{K}M\right)=$ $\tfrac{1}{2}\log\left(\tfrac{m}{2K+1}\right)+K\log\left(\tfrac{2K}{2K+1}\right)+\log M$
$\leq$ $\tfrac{1}{2}\log m-\tfrac{1}{2}\log(2K+1)-\tfrac{1}{2}+\log M$
$\leq$ $\tfrac{1}{2}\log m-\tfrac{1}{2}\log\left(\frac{mM^{2}}{\epsilon^{2}e}\right)-\tfrac{1}{2}+\log M\leq\log\epsilon,$
where the first inequality uses $\tfrac{1}{2}\log\left(\tfrac{m}{2K+1}\right)=\tfrac{1}{2}(\log m-\log(2K+1))$. Also, $\log\left(\tfrac{2K}{2K+1}\right)=\log\left(1-\tfrac{1}{2K+1}\right)\leq-\tfrac{1}{2K+1}$ due to the inequality $\log(1-X)\leq-X$ for $X<1$, and for a large $K$, we approximate $-\tfrac{K}{2K+1}\approx-\frac{1}{2}$. The last inequality uses the assumption $K\geq\frac{1}{2}\left(\frac{mM^{2}}{\epsilon^{2}e}-1\right)$.
Finally, since $\mathcal{A}_{\operatorname*{full}}\subset MB_{2}^{n}$, the covering number satisfies $N(\mathcal{A}_{\operatorname*{full}},\epsilon)\leq C^{\prime}(M/\epsilon)^{n}$ for some constant $C^{\prime}>0$; see Proposition 4.2.12 of Vershynin [2018]. Since $\epsilon\sqrt{\log(M/\epsilon)}\to 0$ as $\epsilon\to 0^{+}$, it follows that $\epsilon\sqrt{\log N(\mathcal{A}_{\operatorname*{full}},\epsilon)}\to 0$ as $\epsilon\to 0^{+}$. ∎
Appendix F Proof of Theorem 11
Consider a partition :=$\mathcal{R}:=\{r_{\ell}\}_{\ell\leq m}$ of the full action space, where each cluster contains more than one action and satisfies Assumption 6. Let $\epsilon:=\max_{r\in\mathcal{R}}\operatorname{diam}(r)$. Let $\mathcal{A}$ be the output of Algorithm 1. Then, for the same constant $C$ in Theorem 7 and another constant $c>0$, for any reference subset $\mathcal{A}^{\prime}$:
$\begin{split}\mathbb{E}_{\theta,\mathcal{A}}[\operatorname*{Regret}]\geq&c\cdot\left(\mathbb{E}_{q}(1-q(r))^{2K}\right)^{1/2}\cdot\mathbb{E}_{\theta}\left[\operatorname*{Regret}\textnormal{of }\mathcal{A}^{\prime}\right]\\ \geq&c\cdot\left(\mathbb{E}_{q}(1-q(r))^{2K}\right)^{1/2}\cdot\left(\min_{r\in\mathcal{R}}\mathbb{E}_{\theta}\left[\max_{a\in r}\mu_{a}\right]-C\epsilon\sqrt{\log|\mathcal{R}|}\right).\end{split}$
Proof.
Choose an arbitrary point $a_{\ell}\in r_{\ell}$ as the representative, such that the subset $\mathcal{A}^{\prime}=\{a_{\ell}\}_{\ell\leq m}$ forms a reference subset of Definition 1. Fix $\ell$. As in Lemma 13, define a Gaussian process $(Z_{a})_{a\in r_{\ell}}$ by setting $Z_{a}:=\mu_{a}-\mu_{a_{\ell}}$, and define random variable $Y_{\ell}:=\sup_{a\in r_{\ell}}Z_{a}$. The key idea is that $\operatorname*{Regret}\geq Y_{\ell}$ whenever $r_{\ell}\cap\mathcal{A}=\emptyset$ and $a^{*}(\theta)\in r_{\ell}$, such that
$\begin{split}\mathbb{E}\left[\operatorname*{Regret}\Big{|}r_{\ell}\cap\mathcal{A}\!=\!\emptyset,a^{*}(\theta)\in r_{\ell}\right]=&\mathbb{E}\left[\max_{a\in r_{\ell}}\mu_{a}-\max_{a^{\prime}\in\mathcal{A}}\mu_{a^{\prime}}\Big{|}r_{\ell}\cap\mathcal{A}\!=\!\emptyset,a^{*}(\theta)\in r_{\ell}\right]\\ \geq&\mathbb{E}\left[\max_{a\in r_{\ell}}\mu_{a}-\mu_{a_{\ell}}\Big{|}r_{\ell}\cap\mathcal{A}\!=\!\emptyset,a^{*}(\theta)\in r_{\ell}\right]\\ =&\mathbb{E}\left[Y_{\ell}\Big{|}a^{*}(\theta)\in r_{\ell}\right],\end{split}$ (19)
where the inequality follows from Assumption 6. The last equality follows from $Y_{\ell}$ is independent from $r_{\ell}\cap\mathcal{A}\!=\!\emptyset$. Further,
$\begin{split}\mathbb{E}\left[\operatorname*{Regret}\right]=&\sum_{r\in\mathcal{R}}\Pr[r\cap\mathcal{A}\neq\emptyset,a^{*}(\theta)\in r]\cdot\mathbb{E}\left[\operatorname*{Regret}\Big{|}r\cap\mathcal{A}\neq\emptyset,a^{*}(\theta)\in r\right]\\ &+\sum_{r\in\mathcal{R}}\Pr[r\cap\mathcal{A}=\emptyset,a^{*}(\theta)\in r]\cdot\mathbb{E}\left[\operatorname*{Regret}\Big{|}r\cap\mathcal{A}=\emptyset,a^{*}(\theta)\in r\right]\\ \geq&\sum_{r\in\mathcal{R}}q(r)\cdot(1-q(r))^{K}\cdot\mathbb{E}\left[\operatorname*{Regret}\Big{|}r\cap\mathcal{A}=\emptyset,a^{*}(\theta)\in r\right],\\ \geq&\sum_{\ell\leq m}q(r_{\ell})\cdot(1-q(r_{\ell}))^{K}\cdot\mathbb{E}\left[Y_{\ell}\Big{|}a^{*}(\theta)\in r_{\ell}\right]\\ \geq&c\cdot\left(\mathbb{E}_{q}(1-q(r))^{2K}\right)^{1/2}\cdot\left(\sum_{\ell\leq m}q(r_{\ell})\cdot\mathbb{E}^{2}\left[Y_{\ell}\Big{|}a^{*}(\theta)\in r_{\ell}\right]\right)^{1/2}\\ \geq&c\cdot\left(\mathbb{E}_{q}(1-q(r))^{2K}\right)^{1/2}\cdot\mathbb{E}_{\theta}\left[\operatorname*{Regret}\textnormal{of }\mathcal{A}^{\prime}\right]\\ \geq&c\cdot\left(\mathbb{E}_{q}(1-q(r))^{2K}\right)^{1/2}\cdot\left(\min_{r\in\mathcal{R}}\mathbb{E}_{\theta}\left[\max_{a\in r}\mu_{a}\right]-C\epsilon\sqrt{\log|\mathcal{R}|}\right)\end{split}$
where the equality uses the tower rule. The first inequality holds because the regret is lower bounded by zero when $r\cap\mathcal{A}\neq\emptyset$ and $a^{*}(\theta)\in r$, and Lemma 17. The second inequality follows from Equation (19). The third inequality uses the Pólya-Szegő inequality [Dragomir, 2004], corresponding to a constant $c>0$:
$c:=\frac{1}{2}\left(\sqrt{\frac{M_{1}M_{2}}{m_{1}m_{2}}}+\sqrt{\frac{m_{1}m_{2}}{M_{1}M_{2}}}\right),$
where $m_{1},m_{2},M_{1},M_{2}$ are constants such that
[…]
Note that $m_{2}$ is positive due to the assumption that each cluster contains more than one action. The fourth inequality follows from Jensen’s inequality that for a random variable $X$, $\mathbb{E}[X^{2}]\geq\mathbb{E}[X]^{2}$, and Lemma 13. Since the reference subset $\mathcal{A}^{\prime}$ is arbitrary, the lower bound $c\cdot\left(\mathbb{E}_{q}(1-q(r))^{2K}\right)^{1/2}\cdot\mathbb{E}_{\theta}\left[\operatorname*{Regret}\textnormal{of }\mathcal{A}^{\prime}\right]$ holds for any reference subset. The last inequality uses Theorem 7. ∎
Appendix G Proof of Theorem 12
Definition 2.
The random process $\{\mu_{a}\}_{a\in\mathcal{S}}$ is a mean-zero sub-Gaussian process if the process $\mathbb{E}\mu_{a}=0$ and has the increment condition:
$\forall u>0,\Pr[|\mu_{a}-\mu_{a^{\prime}}|\geq u]\leq 2\exp\left(-\frac{u^{2}}{2\|a-a^{\prime}\|^{2}_{2}}\right).$
Let $\gamma_{2}(\mathcal{S})$ be the Talagrand’s chaining functional [Talagrand, 2014] for a set $\mathcal{S}\in\mathbb{R}^{n}$, $n\in\mathbb{N}\cup\{+\infty\}$ and Euclidean norm. It is well-known that under the assumption that $\{\mu_{a}\}_{a\in\mathcal{S}}$ is a Gaussian process, this gives the tightest bound, for a universal constant $L$:
$\frac{1}{L}\gamma_{2}(\mathcal{S})\leq\mathbb{E}\left[\max_{a\in\mathcal{S}}\mu_{a}\right]\leq L\gamma_{2}(\mathcal{S}).$
Within the proof of the upper bound for general sub-Gaussian process, it also derives the deviation bound for the term $\max_{a\in\mathcal{S}}\mu_{a}$:
Lemma 19 (Theorem 2.2.22 of Talagrand [2014]).
Let $\{\mu_{a}\}_{a\in\mathcal{S}}$ be a mean-zero sub-Gaussian process, then there exists a constant $c>0$:
$\Pr\left[\sup_{a,a^{\prime}\in\mathcal{S}}|\mu_{a}-\mu_{a^{\prime}}|\geq u\right]\leq 2\exp\left(-\frac{cu^{2}}{\gamma^{2}_{2}(\mathcal{S})}\right).$
Lemma 20.
Let $\{\mu_{a}\}_{a\in\mathcal{A}_{full}}$ be a mean-zero sub-Gaussian process. Consider a partition :=$\mathcal{R}:=\{r_{\ell}\}_{\ell\leq m}$ of the full action space, and an arbitrary reference subset $\mathcal{A}:=\{a_{\ell}\}_{\ell\leq m}$.
Fix $\ell$, define a non-negative random variable
$Y_{\ell}:=\sup_{a\in r_{\ell}}\mu_{a}-\mu_{a_{\ell}}.$
Then, for some constant $C>0$,
$\mathbb{E}_{\theta}\left[\max_{\ell\leq m}Y_{\ell}\right]\leq C\sqrt{\log m}\cdot\max_{\ell\leq m}\gamma_{2}(r_{\ell}).$
Proof.
Let $\epsilon:=\max_{\ell\leq m}\gamma_{2}(r_{\ell})$. For any $\ell\leq m$, we have
$\Pr\left[|Y_{\ell}|\geq u\right]=$ $\Pr\left[\left|\sup_{a\in r_{\ell}}\mu_{a}-\mu_{a_{\ell}}\right|\geq u\right]$
$\leq$ $\Pr\left[\sup_{a,a^{\prime}\in r_{\ell}}|\mu_{a}-\mu_{a^{\prime}}|\geq u\right]$
$\leq$ $2\exp\left(-\frac{cu^{2}}{\gamma^{2}_{2}(r_{\ell})}\right)$
$\leq$ $2\exp\left(-\frac{cu^{2}}{\epsilon^{2}}\right),$
where the equality uses definition of $Y_{\ell}$. The first inequality uses
$\left|\sup_{a\in r_{\ell}}\mu_{a}-\mu_{a_{\ell}}\right|\leq\sup_{a\in r_{\ell}}|\mu_{a}-\mu_{a_{\ell}}|\leq\sup_{a,a^{\prime}\in r_{\ell}}|\mu_{a}-\mu_{a^{\prime}}|.$
The second inequality uses the fact that $\{\mu_{a}\}_{a\in r_{\ell}}$ is a mean-zero sub-Gaussian process and applies Lemma 19. The last inequality uses the definition of $\epsilon$, and $c>0$ is a constant.
By union bound, we have
$\Pr\left[\max_{\ell\leq m}|Y_{\ell}|\geq u\right]\leq 2m\exp\left(-\frac{cu^{2}}{\epsilon^{2}}\right).$
Further, using Lemma 14, we have for some absolute constant $C>0$:
$\mathbb{E}\left[\max_{\ell\leq m}|Y_{\ell}|\right]\leq C\epsilon\sqrt{\log m}.$
Statement of Theorem 12:
Let $\mathcal{A}$ be output of Algorithm 1. Let $\{\mu_{a}\}_{a\in\mathcal{A}_{\operatorname*{full}}}$ be a mean-zero sub-Gaussian process. Consider a partition :=$\mathcal{R}:=\{r_{\ell}\}_{\ell\leq m}$ of the full action space. Then, for the same constant $C>0$,
$\begin{split}\mathbb{E}_{\theta,\mathcal{A}}[\textnormal{Regret}]\leq&C\sqrt{\log m}\cdot\max_{\ell\leq m}\gamma_{2}(r_{\ell})\\ &+\left(\mathbb{E}_{q}\left[(1-q(r))^{2K}\right]\cdot\mathbb{E}_{\theta}\left[\max_{a\in\mathcal{A}_{\operatorname*{full}}}\mu^{2}_{a}\right]\right)^{1/2}.\end{split}$
Proof.
If $r_{\ell}\cap\mathcal{A}\neq\emptyset$, define $a_{\ell}\in r_{\ell}\cap\mathcal{A}$. If $r_{\ell}\cap\mathcal{A}\!=\!\emptyset$, choose an arbitrary point $a_{\ell}\in r_{\ell}$ as the representative. The set $\mathcal{A}^{\prime}:=\{a_{\ell}\}_{\ell\leq m}$ forms a reference subset of Definition 1.
For each $\ell\leq m$, define a random variable
$Y_{\ell}:=\sup_{a\in r_{\ell}}\mu_{a}-\mu_{a_{\ell}}.$
Following the same reasoning used in the proof of Theorem 8:
$\mathbb{E}[\operatorname*{Regret}]\leq$ $\mathbb{E}\left[\max_{\ell\leq m}Y_{\ell}\right]+\left(\mathbb{E}_{q}\left[(1-q(r))^{2K}\right]\cdot\mathbb{E}_{\theta}\left[\max_{a\in\mathcal{A}_{\operatorname*{full}}}\mu^{2}_{a}\right]\right)^{1/2}$
$\leq$ $C\sqrt{\log m}\cdot\max_{\ell\leq m}\gamma_{2}(r_{\ell})+\left(\mathbb{E}_{q}\left[(1-q(r))^{2K}\right]\cdot\mathbb{E}_{\theta}\left[\max_{a\in\mathcal{A}_{\operatorname*{full}}}\mu^{2}_{a}\right]\right)^{1/2},$
The first inequality follows from Equation (14) in the proof of Theorem 8. The last inequality uses Lemma 20. ∎
Appendix H The effect of clustering structure on regret
In Figure 2, we study the effect of the cluster diameters (controlled by a spread parameter) on regret. We structure the clustered action space: Five center points are fixed on the unit sphere in $\mathbb{R}^{3}$, and around each center, 200 points are sampled to form five clusters. Each point is obtained by adding Gaussian noise (mean zero, standard deviation equal to the spread parameter) to the center direction, followed by projection back onto the unit sphere. Bandits are sampled from a 3-dimensional standard Gaussian distribution, i.e., $\theta\sim\mathcal{N}(0,I)$. The left subplot shows the expected regret of Algorithm 1 with $K=10$, computed using $10^{4}$ additional bandits, as the spread varies from 0.01 to 0.5. The curves and error shade represent the mean $\pm$ one standard deviation of expected regret over 30 repetitions. The middle and right subplots display example action spaces for spread values of 0.01 and 0.5, with representative actions (purple stars) selected by Algorithm 1 with $K=10$.
Figure 2: Illustration of clustered action spaces on unit sphere in $\mathbb{R}^{3}$ and the effect of cluster diameters on regret. Five clusters are formed by generating 5 fixed center points, with 200 points sampled around each using Gaussian noise (spread controls the variance). Bandits are drawn from $\mathcal{N}(0,I)$. The left subplot shows the mean $\pm$ standard deviation of the expected regret (over 30 trials) as the spread varies from 0.01 to 0.5, using $10^{4}$ additional bandits. The middle and right subplots show example action spaces (blue dots) for spread values 0.01 and 0.5, with representative actions (purple stars) selected by Algorithm 1 with $K=10$.
Appendix I Varying-dependence actions with RBF/Gibbs kernels
We study the effect of varying action dependence using a Gaussian process with a kernel. To control the degree of dependence, we use stationary RBF kernel and non-stationary Gibbs kernel [Williams and Rasmussen, 2006].
[…]
where $l$ is a length-scale parameter and $l(a):=0.1+0.9\cdot\exp(-\|a\|^{2})$ is a location-dependent length-scale function. Both of them control the dependence between actions. But, unlike the stationary RBF kernels, the Gibbs kernel allows the correlation to depend not only on the distance between actions, but also on their locations. When $l(a)=l$ is a constant, the Gibbs kernel reduces to the RBF kernel.
Sampling Outcome Functions from a Kernel: We first construct the kernel matrix $\mathbf{K}$, where each entry is given by $\mathbf{K}_{a,a^{\prime}}=k(a,a^{\prime})$, for $a,a^{\prime}\in\mathcal{A}_{\operatorname*{full}}$, depending on the choice of kernel. We then sample a Gaussian vector (a Gaussian process function evaluated at finite input) $f\sim\mathcal{N}(0,\mathbf{K})$. Under either kernels defined in Equation (20), the variance of the function value $f(a)$ is one for all $a\in\mathcal{A}_{\operatorname*{full}}$. In this way, we sample functions from a RKHS function class; See [Kanagawa et al., 2018, Theorem 4.12].
Figure 3: Experiments with outcome functions sampled from RBF/Gibbs kernels in Equation (20). Sampled outcome functions from Gibbs kernel over fixed 1000 grid points in [0,2][0,2][ 0 , 2 ] (blue curves, right y-axis). The histogram (purple bars, left y-axis) shows action selection frequencies by Algorithm 1 with $K=5000$, favoring regions with rougher functions and edge points.
To study the effect of varying dependence on the output of Algorithm 1, we consider a fixed action space consisting of 1000 grid points in the interval [0,2][0,2][ 0 , 2 ], using Gibbs kernel in Equation (20) to sample outcome functions. To simplify computations, we marginalize a Gaussian process defined by the kernel over the grid. Figure 3 provides examples of sampled outcome functions (blue curves, with the y-axis on the right-hand side), which become smoother as the actions approach the left end of the interval—indicating stronger correlations among function values in that region. We run Algorithm 1 on this action space with $K=5000$ to select actions and record the frequency of each action being selected. The resulting histogram (purple bars, with the y-axis on the left) reflects the importance measure $q$, highlighting that Algorithm 1 tends to select more actions from regions where the outcome functions are rougher—i.e., where action outcomes are less correlated and their features $\Phi(a)$ are are farther apart. Another interesting aspect of this subplot is the two high bars at the edges. Recall that the actual action space consists of feature vectors $\Phi(a)$ for $a\in\mathcal{A}_{ful}$. For actions indexed closer to 0, their feature vectors become more densely packed compared to those indexed closer to 2, resulting in more correlated outcomes. The two actions at the edges, indexed by 0 and 2, correspond to the two farthest points in the actual feature space.
Appendix J Independent and identically distributed actions
As an extreme case, suppose that $\{\mu_{a}\}_{a\in\mathcal{A}_{\operatorname*{full}}}$ is a set of i.i.d. random variables. This corresponds to the canonical process in which the action space is given by the orthonormal basis of $\mathbb{R}^{n}$, where $n=|\mathcal{A}_{\operatorname*{full}}|$. To see this, we associate each action $a$ with a unit vector $e_{a}$, which has a value of 111 at the $a$th coordinate and 0 elsewhere, and define the expected outcome as $\mu_{a}(\theta):=\langle e_{a},\theta\rangle$. With this construction, the collection $\{\mu_{a}\}_{a\in\mathcal{A}_{\operatorname*{full}}}$ consists of mutually independent random variables.
Let the set $\mathcal{A}_{\operatorname*{full}}=\{e_{i}:i=1,\dots,n\}$ denote the $n$ unit vectors aligned with the coordinate axes in $\mathbb{R}^{n}$. Hence, $\operatorname{diam}(\mathcal{A}_{\operatorname*{full}})=\sqrt{2}$. In this case, $\max_{a\in\mathcal{A}_{\operatorname*{full}}}\langle a,\theta\rangle$ is equivalent to the maximum among the $n$ entries of $\theta$, where each entry is i.i.d. from the standard normal distribution $\mathcal{N}(0,1)$. By symmetry, each coordinate has the same probability of attaining the maximum value. If each cluster contains only one unit vector, then the importance measure $q$ over clusters is the uniform distribution.
Expected maximum in $\mathcal{A}_{\operatorname*{full}}$: Let $X_{i}$, $i=1,\dots,n$ be i.i.d. samples from $\mathcal{N}(0,1)$, then
$\mathbb{E}\max_{a\in\mathcal{A}_{\operatorname*{full}}}\mu_{a}=\mathbb{E}_{\theta}\max_{i=1,\dots,n}\theta_{i}=\mathbb{E}\max_{i=1,\dots,n}X_{i},$
where the equivalence comes from that each entry $\theta_{i}$ is i.i.d. sample from $\mathcal{N}(0,1)$. Then, according to the bounds of expected maximum of Gaussian in Supplementary L, we have
$\frac{\sqrt{\log n}}{\sqrt{\pi\log 2}}\leq\;\mathbb{E}\max_{a\in\mathcal{A}_{\operatorname*{full}}}\mu_{a}\leq\sqrt{2\log n},\quad\frac{\sqrt{\log|\mathcal{A}|}}{\sqrt{\pi\log 2}}\leq\;\mathbb{E}\max_{a\in\mathcal{A}}\mu_{a}\leq\sqrt{2\log|\mathcal{A}|}.$
Bounds of expected regret of arbitrary $\mathcal{A}$: By definition of regret in Equation (1),
$\frac{\sqrt{\log n}}{\sqrt{\pi\log 2}}-\sqrt{2\log|\mathcal{A}|}\leq\mathbb{E}_{\theta}[\textnormal{Regret}]\leq\sqrt{2\log n}-\frac{\sqrt{\log|\mathcal{A}|}}{\sqrt{\pi\log 2}}.$
As a result, any algorithm, including Algorithm 1, would perform poorly unless the subset size $|\mathcal{A}|$ is sufficiently large.
Appendix K Properties of Gaussian width
Given a set $\mathcal{S}\in\mathbb{R}^{n}$, the term $\mathbb{E}[\max_{a\in\mathcal{S}}\mu_{a}]$ where $\theta\sim\mathcal{N}(0,I)$ is called Gaussian (mean) width.
Claim 1: $\mathbb{E}\left[\max_{a\in\mathcal{S}}\langle a,\theta\rangle\right]=\mathbb{E}\left[\max_{a^{\prime}\in-\mathcal{S}}\langle a^{\prime},\theta\rangle\right]$.
$\mathbb{E}\left[\max_{a\in\mathcal{S}}\langle a,\theta\rangle\right]=\mathbb{E}\left[\max_{a\in\mathcal{S}}\langle a,-\theta\rangle\right]=\mathbb{E}\left[\max_{a\in\mathcal{S}}\langle-a,\theta\rangle\right]=\mathbb{E}\left[\max_{a^{\prime}\in-\mathcal{S}}\langle a^{\prime},\theta\rangle\right],$
where the first equality uses $-\theta$ and $\theta$ are identically distributed. The third equality uses for any $a\in\mathcal{S}$, it holds that $-a\in-\mathcal{S}$.
Claim 2: $\mathbb{E}\left[\max_{a\in\mathcal{S}}\langle a,\theta\rangle\right]\leq\frac{1}{2}\mathbb{E}\left[\max_{a,a^{\prime}\in\mathcal{S}}\langle a-a^{\prime},\theta\rangle\right]$.
Let $a^{*}(-\mathcal{S},\theta)$ denote the optimal action in $\mathcal{S}$ for bandit instance $\theta$.
$2\cdot\mathbb{E}\left[\max_{a\in\mathcal{S}}\langle a,\theta\rangle\right]=$ $\mathbb{E}\left[\max_{a\in\mathcal{S}}\langle a,\theta\rangle\right]+\mathbb{E}\left[\max_{a^{\prime}\in-\mathcal{S}}\langle a^{\prime},\theta\rangle\right]$
$=$ $\mathbb{E}\left[\langle a^{*}(\mathcal{S},\theta),\;\theta\rangle\right]+\mathbb{E}\left[\langle a^{*}(-\mathcal{S},\theta),\;\theta\rangle\right]$
$=$ $\mathbb{E}\left[\langle a^{*}(\mathcal{S},\theta)+a^{*}(-\mathcal{S},\theta),\;\theta\rangle\right]$
$\leq$ $\mathbb{E}\left[\max_{a,a^{\prime}\in\mathcal{S}}\langle a-a^{\prime},\theta\rangle\right],$
where the first equality uses Claim 1, the second equality uses the definition of $a^{*}(-\mathcal{S},\theta)$. The third equality uses linearity of expectation. The inequality uses that the vector $a^{*}(\mathcal{S},\theta)+a^{*}(-\mathcal{S},\theta)$ belongs to the set of vectors $\{a-a^{\prime}:a,a^{\prime}\in\mathcal{S}\}$. In fact, one can prove the equality that $\mathbb{E}\left[\max_{a\in\mathcal{S}}\langle a,\theta\rangle\right]=\frac{1}{2}\mathbb{E}\left[\max_{a,a^{\prime}\in\mathcal{S}}\langle a-a^{\prime},\theta\rangle\right]$, but we only need this inequality to prove the next claim.
Claim 3: $\mathbb{E}\left[\max_{a\in\mathcal{S}}\mu_{a}\right]\leq\frac{\operatorname{diam}(\mathcal{S})}{2}\cdot\sqrt{n}$.
$\mathbb{E}\left[\max_{a\in\mathcal{S}}\mu_{a}\right]=$ $\mathbb{E}\left[\max_{a\in\mathcal{S}}\langle a,\theta\rangle\right]$
$\leq$ $\frac{1}{2}\mathbb{E}\left[\max_{a,a^{\prime}\in\mathcal{S}}\langle a-a^{\prime},\theta\rangle\right]$
$\leq$ $\frac{1}{2}\mathbb{E}\max_{a,a^{\prime}\in\mathcal{S}}\|\theta\|_{2}\|a-a^{\prime}\|_{2}$
$\leq$ $\frac{1}{2}\mathbb{E}\operatorname{diam}(\mathcal{S})\|\theta\|_{2}\leq\frac{\operatorname{diam}(\mathcal{S})}{2}\cdot\sqrt{n},$
where the first equality uses the definition of $\mu_{a}$. The first inequality uses Claim 2. The second inequality uses Cauchy-Schwarz inequality. The third inequality uses the definition of $\operatorname{diam}(\cdot)$. The last inequality uses $\mathbb{E}\|\theta\|_{2}\leq\sqrt{n}$.
Appendix L Bounds of expected maximum of Gaussian
Let $X_{1},\dots,X_{N}$ be $N$ random Gaussian variables (no necessarily independent) with zero mean and variance of marginals smaller than $\sigma^{2}$, then
$\mathbb{E}\left[\max_{i=1,\dots,N}X_{i}\right]\leq\sigma\sqrt{2\log N}.$
Proof.
for any $\delta>0$,
$\mathbb{E}\left[\max_{i=1,\dots,N}X_{i}\right]=$ $\frac{1}{\delta}\mathbb{E}\left[\log\exp(\delta\max_{i=1,\dots,N}X_{i})\right]\leq\frac{1}{\delta}\log\mathbb{E}\left[\exp(\delta\max_{i=1,\dots,N}X_{i})\right]$
$=$ $\frac{1}{\delta}\log\mathbb{E}\left[\max_{i=1,\dots,N}\exp(\delta X_{i})\right]\leq\frac{1}{\delta}\log\sum_{i=1}^{N}\mathbb{E}\left[\exp(\delta X_{i})\right]$
$\leq$ $\frac{1}{\delta}\log\sum_{i=1}^{N}\exp(\sigma^{2}\delta^{2}/2)=\frac{\log N}{\delta}+\frac{\sigma^{2}\delta}{2},$
where the first inequality uses Jensen’s inequality. Taking $\delta:=\sqrt{2(\log N)/\sigma^{2}}$ yields the results. ∎
Let $X_{1},\dots,X_{N}$ be i.i.d. $\mathcal{N}(0,\sigma^{2})$ random variables, then according to [Kamath, 2015]:
$\mathbb{E}_{\theta}\left[\max_{i=1,\dots,N}X_{i}\right]\geq\frac{\sigma\sqrt{\log N}}{\sqrt{\pi\log 2}}.$